### v1

# On the rank index of projective curves of almost minimal degree

Abstract.
In this article, we investigate the rank index of projective curves $\mathcal{C}\subset\mathbb{P}^{r}$ of degree $r+1$ when $\mathcal{C}=\pi_{p}(\tilde{\mathcal{C}})$ for the standard rational normal curve $\tilde{\mathcal{C}}\subset\mathbb{P}^{r+1}$ and a point $p\in\mathbb{P}^{r+1}\setminus\tilde{\mathcal{C}}^{3}$. Here, the rank index of a closed subscheme $X\subset\mathbb{P}^{r}$ is defined to be the least integer $k$ such that its homogeneous ideal can be generated by quadratic polynomials of rank $\leq k$. Our results show that the rank index of $\mathcal{C}$ is at most $4$, and it is exactly equal to $3$ when the projection center $p$ is a coordinate point of $\mathbb{P}^{r+1}$. We also investigate the case where $p\in\tilde{\mathcal{C}}^{3}\setminus\tilde{\mathcal{C}}^{2}$.
Key words and phrases:
apolarity, linear projections, low rank quadrics, projective curves with parametrization, property $\operatorname{QR}(k)$
2020 Mathematics Subject Classification:
14A25, 14H45, 14N05, 15A63, 16E45
1. Introduction
Throughout this paper we work over an algebraically closed field $\mathbb{K}$ with $\operatorname{char}(\mathbb{K})=0$. Due to [HLMP21], a non-degenerate projective variety $X\subset\mathbb{P}^{r}$ is said to satisfy property $\operatorname{QR}(k)$ if its homogeneous ideal is generated by quadratic polynomials of rank at most $k$. When $X$ is cut out ideal-theoretically by quadrics, we define the rank index of $X$, denoted by $\operatorname{rank-index}(X)$, as the least integer $k$ such that $X$ satisfies property $\operatorname{QR}(k)$. Thus
$\operatorname{rank-index}(X)\geq 3$
since $X$ is irreducible and nondegenerate in $\operatorname{\mathbb{P}}^{r}$. Many classical constructions in projective geometry such as rational normal scrolls and Segre-Veronese varieties satisfy property $\operatorname{QR}(4)$ since their homogeneous ideals are generated by $2$-minors of one or several matrices of linear forms (cf. [Ber08], [H0̀2], [Puc98], [SS11], [Sul08], etc). Recently, it has been found that there are many projective varieties whose rank index is equal to $3$, which is possibly the minimal value. The $d$-th Veronese embedding of $\operatorname{\mathbb{P}}^{n}$ satisfies property $\operatorname{QR}(3)$ for all $n\geq 1$ and $d\geq 2$ if $\operatorname{char}(\mathbb{K})\neq 3$ (cf. [HLMP21]). Every linearly normal curve of arithmetic genus $g$ and degree $\geq 4g+4$ satisfies property $\operatorname{QR}(3)$ (cf. [Par22]). Also, the authors in [MP23] found the rank index of various projective varieties, including rational normal scrolls and del Pezzo varieties. For example, if $X$ is a smooth rational normal scroll of dimension $\geq 2$, then $\operatorname{rank-index}(X)=4$.
The main purpose of this paper is to investigate the rank index of projective curves $\mathcal{C}\subset\operatorname{\mathbb{P}}^{r}$ of degree $r+1$. To put things in perspective, we recall some of the previously known results regarding this problem. To give precise statements, we require some notation and definitions. Let $\mathcal{C}\subset\operatorname{\mathbb{P}}^{r}$ be a nondegenerate projective curve of degree $d$. Then $d\geq r$ and equality is attained if and only if it is a rational normal curve, which is obviously very well-understood. In particular, it holds that $\operatorname{rank-index}(\mathcal{C})=3$ (cf. [HLMP21, Corollary 2.4]). The next case, i.e., $d=r+1$, is exactly one of the following two cases:
  1. Case $1$ $\mathcal{C}$ is a linearly normal curve of arithmetic genus one;
  2. Case $2$ $\mathcal{C}$ is an isomorphic projection of a rational normal curve $\widetilde{\mathcal{C}}\subset\operatorname{\mathbb{P}}^{r+1}$.
In the former case, it is already known that $\operatorname{rank-index}(\mathcal{C})=3$ (cf. [Par22, Theorem 1.1]). In the latter case, let
$\tilde{\mathcal{C}}=\mathcal{C}_{r+1}=\left\{[s^{r+1}:s^{r}t:\cdots:t^{r+1}]~{}|~{}[s:t]\in\mathbb{P}^{1}\right\}\subset\mathbb{P}^{r+1},d\geq 4$
be the standard rational normal curve of degree $r+1$, and let $\tilde{\mathcal{C}}^{k}$ denote the $k$-th join of $\tilde{\mathcal{C}}$ with itself. That is, $\tilde{\mathcal{C}}^{k}$ is the closure of the set of points lying in $(k-1)$-dimensional linear subspaces spanned by general collections of $k$ points in $\tilde{\mathcal{C}}$. For joins of projective varieties, we refer to [Zak04]. Then
$\mathcal{C}=\pi_{p}(\tilde{\mathcal{C}})$
for some $p\in\operatorname{\mathbb{P}}^{r+1}\setminus\tilde{\mathcal{C}}^{2}$ where $\pi_{p}:\operatorname{\mathbb{P}}^{r+1}\dashrightarrow\operatorname{\mathbb{P}}^{r}$ is the linear projection from $p$. For a point $p\in\operatorname{\mathbb{P}}^{r+1}$, the integer
$\mbox{rank}_{\tilde{\mathcal{C}}}(p)=\min\{~{}k~{}\mid~{}p\in\tilde{\mathcal{C}}^{k}~{}\},$
is called the rank of $p$ with respect to $\tilde{\mathcal{C}}$, or simply $\tilde{\mathcal{C}}$-rank of $p$. Note that if $p_{c}$ is the $c$-th coordinate point of $\mathbb{P}^{r+1}$ for some $0\leq c\leq\lfloor(r+1)/2\rfloor$, then $\mbox{rank}_{\tilde{\mathcal{C}}}(p)=c+1$ (cf. [Par07, Section 5]). It is well known that various properties of $\mathcal{C}$ are determined by the rank of $p$ with respect to $\tilde{\mathcal{C}}$. For example, if $p\in\operatorname{\mathbb{P}}^{r+1}\setminus\tilde{C}$, then $\mathcal{C}=\pi_{p}(\tilde{\mathcal{C}})$ is an isomorphic projection of $\widetilde{\mathcal{C}}$ if and only if $\mbox{rank}_{\tilde{\mathcal{C}}}(p)\geq 3$. Moreover, the graded Betti numbers of $\mathcal{C}$ are uniquely determined by $\mbox{rank}_{\tilde{\mathcal{C}}}(p)$ (cf. [Par07] and [LP16]). In particular, the homogeneous ideal of $\mathcal{C}$ is generated by quadratic polynomials if and only if $\mbox{rank}_{\tilde{\mathcal{C}}}(p)\geq 4$. Also there is a surprising connection between some invariants of $\mathcal{C}$ in real algebraic geometry and the rank of $p$ with respect to $\tilde{\mathcal{C}}$ (cf. [BCJ24]).
Now, let us state our first main result in this paper.
Theorem 1.1.
For the standard rational normal curve $\widetilde{\mathcal{C}}\subset\operatorname{\mathbb{P}}^{r+1}$ and a point $p\in\operatorname{\mathbb{P}}^{r+1}$ with $\mbox{rank}_{\tilde{\mathcal{C}}}(p)\geq 4$, let
$\mathcal{C}=\pi_{p}(\tilde{\mathcal{C}})\subset\operatorname{\mathbb{P}}^{r}$
be a smooth rational curve of degree $r+1$. Then
  1. (1)
    $\operatorname{rank-index}(\mathcal{C})\leq 4$.
  2. (2)
    Suppose that $p=p_{c}$ is the $c$-th coordinate point of $\mathbb{P}^{r+1}$ (and hence $23\leq c\leq r-$ 2). Then $\operatorname{rank-index}(\mathcal{C})=3$.
The proofs of Theorem 1.1. (1) and (2) are respectively in Section 3 and Section 4.
One of the reasons why the problem of determining the rank index of the curve $\mathcal{C}$ in Theorem 1.1 is interesting is that, unlike previously known varieties satisfying property $\operatorname{QR}(3)$, which are all linearly normal, there is the difficulty that $\mathcal{C}$ is not linearly normal and hence its properties change continuously as the projection center $p$ varies.
In the proof of Theorem 1.1.(1), we use the fact that there is a unique rational normal surface scroll $\mathcal{S}$ that contains $\mathcal{C}$. Indeed, the homogeneous ideal of $\mathcal{C}$ is generated by that of $\mathcal{S}$ and the quadratic polynomials obtained from the equation of $\mathcal{C}$ as a divisor of $\mathcal{S}$. Note that the existence of such a surface $\mathcal{S}$ can be found in [EEK88, Proposition 1.2] and [BS07, Theorem 7.3] and the uniqueness of $\mathcal{S}$ is recently proven in [KMP24, Theorem 5.2]. We provide a new constructive proof of the existence of $\mathcal{S}$ by using the apolar ideal of the binary form corresponding to the projection center $p$ in Section 3.
The proof of Theorem 1.1. (2) is obtained by combining the framework of induction on the degree of $\mathcal{C}$ with the investigation of rank $3$ quadratic equations of $\tilde{\mathcal{C}}$ that are also contained in the homogeneous ideal of $\mathcal{C}$.
Regarding Theorem 1.1, one can naturally ask whether there exists a point $p\in\operatorname{\mathbb{P}}^{r+1}\setminus\tilde{\mathcal{C}}^{3}$ such that the rank index of $\mathcal{C}=\pi_{p}(\tilde{\mathcal{C}})$ is equal to $4$. We do not have such an example. On the contrary, Theorem 1.1.(2) leads us to formulate the following conjecture.
Conjecture 1.2.
Let $\mathcal{C}\subset\operatorname{\mathbb{P}}^{r}$ be as in Theorem 1.1. Then $\operatorname{rank-index}(\mathcal{C})=3$.
Next, we consider the case where $\mbox{rank}_{\tilde{\mathcal{C}}}(p)=3$ and so the homogeneous ideal of $\mathcal{C}$ fails to be generated by quadratic polynomials. Then, $\mathcal{C}$ admits a unique trisecant line, say $\mathbb{L}$. Moreover, the homogeneous ideal of the reducible curve
$X:=\mathcal{C}\cup\mathbb{L}$
is generated by the quadratic polynomials in the homogeneous ideal of $\mathcal{C}$. Note that $\mathcal{C}$ is $3$-regular and hence $\mbox{length}(\mathcal{C}\cap\mathbb{L})=3$ (cf. [HSV91]). Regarding the rank index of $X$, we obtain the following interesting partial result.
Theorem 1.3.
Suppose that $r\geq 4$. Let $\mathcal{C}$, $\mathbb{L}$, and $X=\mathcal{C}\cup\mathbb{L}$ be as above. Then
  1. (1)
    $\operatorname{rank-index}(X)\leq 4$.
  2. (2)
    If $\mathcal{C}\cap\mathbb{L}$ has at most two points, or equivalently, if $\mathcal{C}\cap\mathbb{L}$ contains a point with multiplicity at least two, then $\operatorname{rank-index}(X)=4$.
The proofs of Theorem 1.3 is in Section 5.
By Theorem 1.3, it remains to determine the rank index of $X$ when $\mathcal{C}\cap\mathbb{L}$ consists of three simple points. There are many example such that $X$ can satisfy property $\operatorname{QR}(3)$ and hence $\operatorname{rank-index}(X)=3$. This leads us to formulate the following conjecture.
Conjecture 1.4.
Let $\mathcal{C}$, $\mathbb{L}$ and $X=\mathcal{C}\cup\mathbb{L}$ be as in Theorem 1.3. Then $\operatorname{rank-index}(X)=3$ if and only if $\mathcal{C}\cap\mathbb{L}$ consists of three simple points.
Obviously, Theorem 1.3 proves one direction of 1.4.
Organization of the paper.
Section 2 outlines key definitions and foundational concepts including (smooth) curves of almost minimal degree, apolarity of binary forms, and the rank index of varieties. Section 3 presents our primary results on the rank index for projected curves. In Section 4, we focus on the quadratic rank index of projected curves obtained by linear projection from a standard basis in projective spaces. Section 5 includes explicit calculations for certain projective schemes not covered in the previous sections that provide additional supports for conjectures concerning the rank index.
Acknowledgement. The first named author was supported by the Institute for Basic Science (IBS-R032-D1-2023-a00). The second named author was supported by a KIAS Individual Grant(MG083101) at Korea Institute for Advanced Study. The third named author was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. 2022R1A2C1002784).
2. Preliminaries
2.1. Apolarity on binary forms
Given a polynomial ring $\mathbb{K}[s,t]$ with the standard grading i.e., $\deg(s)=\deg(t)=1$. A monomial is denoted by $s^{a}t^{b}$ for nonnegative integers $a$ and $b$. We define the degree of a monomial $s^{a}t^{b}$ as $a+b$.
Definition 2.1.
Let $R=\mathbb{K}[s,t]$ be the graded ring over $\mathbb{K}$ with the standard grading on the variables. That is, $R=\bigoplus_{i\in\mathbb{Z}_{\geq 0}}R_{i}$ where $R_{i}$ is the space of homogeneous polynomials of degree $i$. Let $\mathcal{R}=\mathbb{K}[S,T]$ be a graded ring whose graded components $\mathcal{R}_{i}$ are the $\mathbb{K}$-vector spaces of homogeneous polynomials of degree $i$.
We consider a bilinear map of vector spaces
$\mathcal{R}_{j}\times R_{i}\rightarrow R_{i-j}$
defined by
$(F,f)\mapsto F\circ f=\sum_{0\leq k\leq j}c_{k}\frac{\partial^{j}}{\partial s^{j-k}\partial t^{k}}f$
where $F=\sum_{0\leq k\leq j}c_{k}S^{j-k}T^{k}$. The apolar ideal of a homogeneous polynomial $f\in R$ is defined as the set of homogeneous polynomials $F\in\mathcal{R}$ such that $F\circ f=0$, denoted by $(f)^{\perp}$. In other words,
$(f)^{\perp}:=\{F\in\mathcal{R}:F\circ f=0\}.$
Proposition 2.2.
(cf. [IK99, Theorem 1.44.]) Suppose $f\in\mathbb{K}[s,t]$ is a binary form of degree $d$. Then, $(f)^{\perp}$ is a complete intersection ideal generated by two binary forms in $\mathcal{R}$. This means that if we let $F_{l}$ be the lower-degree generator and $F_{h}$ the higher-degree generator of $(f)^{\perp}\subset\mathbb{K}[S,T]$, then
$(f)^{\perp}=(F_{l},F_{h}),$
where $V(F_{l},F_{h})=\emptyset$ and $\deg F_{l}+\deg F_{h}=d+2$ with $\deg F_{l}\leq\deg F_{h}$. If $\deg F_{l}<\deg F_{h}$, then $F_{l}$ is uniquely determined by $f$, while $F_{h}$ is unique modulo the principal ideal $\langle F_{l}\rangle$.
For example, if $f=\alpha s+\beta t\in\mathbb{K}[s,t]$ is a linear binary form, then the lower-degree generator of $(f)^{\perp}$ is $F_{l}=\beta S-\alpha T\in\mathcal{R}$ and $F_{h}$ is any quadratic forms in $\mathcal{R}$.
The apolarity of forms is a useful tool for studying the decomposition of forms into sums of powers of linear forms. This decomposition is known as the Waring decomposition or the (generalized) additive decomposition of forms. For more details on the generalized additive decomposition and apolarity, see [IK99].
We remark that given a vector space $R_{d}$ for a fixed integer $d$, the pairing $R_{d}\times\mathcal{R}_{d}\rightarrow R_{0}=\mathbb{K}$ forms an inner product on these vector spaces. Therefore, we can identify $\mathcal{R}_{d}$ with the dual vector space of $R_{d}$ via the differential operator. In particular, for $f\in R_{d}$ and $g\in\mathcal{R}_{d}$, the operation $g\circ f=f\circ g$ is defined by
$g\circ f=\sum_{i=0}^{d}c_{i}\left(\frac{\partial}{\partial s}\right)^{d-i}\left(\frac{\partial}{\partial t}\right)^{i}f(s,t),$
where $g(S,T)=\sum_{i=0}^{d}c_{i}S^{d-i}T^{i}$.
Let $\{S^{d},S^{d-1}T,\cdots,T^{d}\}$ be the monomial basis of $\mathcal{R}_{d}$. Consider the dual basis of this monomial basis for $R_{d}$ with above inner product, which is given by $\{s^{[d]},\cdots,s^{[d-i]}t^{[i]},\cdots,t^{[d]}\}:=\{\frac{1}{d!}s^{d},\cdots,\frac{1}{(d-i)!i!}s^{d-i}t^{i},\cdots,\frac{1}{d!}t^{d}\}$. Indeed,
(2.1) $S^{d-i}T^{i}\circ s^{[d-j]}t^{[j]}=\delta_{ij}$
where $\delta_{ij}$ is the Kronecker delta. Therefore, any form $f(s,t)$ can be expressed as
(2.2) $f(s,t)=\sum_{i=0}^{d}\left(S^{d-i}T^{i}\circ f(s,t)\right)s^{[d-i]}t^{[i]}.$
We have the following identity:
(2.3) $(\alpha s+\beta t)^{d}=\sum_{i=0}^{d}\binom{d}{i}\alpha^{d-i}\beta^{i}s^{d-i}t^{i}=d!\sum_{i=0}^{d}\alpha^{d-i}\beta^{i}s^{[d-i]}t^{[i]}$
Lemma 2.3.
Let $d$ be a positive integer. For binary forms $F(S,T)\in\mathbb{K}[S,T]_{d}$ and $(\alpha s+\beta t)^{d}\in\mathbb{K}[s,t]_{d}$,
$F(S,T)\circ(\alpha s+\beta t)^{d}=d!F(\alpha,\beta)$
Proof.
Let $F(S,T)=\sum_{j=0}^{d}c_{j}S^{d-j}T^{j}\in\mathbb{K}[S,T]_{d}$. By (2.1) and (2.3),
$F(S,T)\circ(\alpha s+\beta t)^{d}$ $=d!\sum_{i=0}^{d}\sum_{j=0}^{d}\alpha^{d-i}\beta^{i}c_{j}\left(S^{d-i}T^{i}\circ s^{[d-j]}t^{[j]}\right)$
$=d!\sum_{i=0}^{d}c_{i}\alpha^{d-i}\beta^{i}=d!F(\alpha,\beta)$
Thus, the inner product between $F(S,T)$ and $(\alpha s+\beta t)^{d}$ can be interpreted as the evaluation of $F(S,T)$ at $(\alpha,\beta)$.
2.2. Projected curves
We now discuss a correspondence between binary forms of degree $d$ and points in a projective space $\mathbb{P}^{d}$. Let $\nu_{d}:\mathbb{P}^{1}\to\mathbb{P}^{d}$ be the $d$-uple embedding (or the $d$th Veronese map). Let $\mathcal{C}_{d}:=\nu_{d}(\mathbb{P}^{1})\subseteq\mathbb{P}^{d}$ denote the image of the Veronese map, which is known as the standard rational normal curve.
Given a point $p\in\mathbb{P}^{d}$, consider the vector space of linear forms on $\mathbb{P}^{d}$ that vanish at $p$. Pulling this space back to $\mathbb{P}^{1}$ via $\nu_{d}$ yields a $d$-dimensional vector space of degree $d$ binary forms, which corresponds to a hyperplane in the space of binary forms of degree $d$. We denote by $f_{p}$ the binary form of degree $d$ (unique up to scalar multiplication) that is orthogonal to this hyperplane with respect to the inner product. In particular, one can see that given a binary form $f_{p}$ in the vector space $\mathbb{K}[s,t]_{d}$, the orthogonal complement of $f_{p}$ with respect to the inner product defined by the differential operator is $((f_{p})^{\perp})_{d}$.
An alternative way to associate a point $p$ with the binary form $f_{p}$ is as follows. Under the $d$-uple embedding $\nu_{d}$, a point $\nu_{d}([\alpha:\beta])$ on the rational normal curve corresponds to the $d$th power $(\alpha s+\beta t)^{d}\in\mathbb{K}[s,t]_{d}$ (using the identification between points of $\mathbb{P}^{1}$ and binary linear forms). Since distinct points on the rational normal curve are in linearly general position, this extends additively to give a correspondence between all points in $\mathbb{P}^{d}$ and binary forms of degree $d$.
Explicitly, for $p\in\mathbb{P}^{d}$, we may express $p$ as a linear combination of $r\leq d+1$ points on $\mathcal{C}_{d}$, say $p=\sum_{i=1}^{r}c_{i}p_{i}$. Setting $p_{i}=\nu_{d}([\alpha_{i}:\beta_{i}])$, we have
$f_{p}=\sum_{i=1}^{r}c_{i}(\alpha_{i}s+\beta_{i}t)^{d}\in\mathbb{K}[s,t]_{d}.$
In this correspondence, the torus-fixed points $e_{i}\in\mathbb{P}^{d}$ correspond to the monomials $s^{[d-i]}t^{[i]}\in\mathbb{K}[s,t]_{d}$.
Example 2.4.
Let $\mathcal{C}_{6}$ be the standard rational normal curve of degree $6$ in $\mathbb{P}^{6}$, and let $p=e_{2}+e_{4}$ be a point where $e_{2}$ and $e_{4}$ are the second and fourth torus-fixed point in $\mathbb{P}^{6}$, respectively. Then
$(f_{p})^{\perp}=(S^{3}T-ST^{3},S^{4}-2S^{2}T^{2}+T^{4})\subset\mathbb{K}[S,T].$
Note that there is a unique degree $6$ form $f_{p}$, up to constant multiples, determined by $(f_{p})^{\perp}$:
$f_{p}=s^{[4]}t^{[2]}+s^{[2]}t^{[4]}=\frac{1}{4!2!}s^{4}t^{2}+\frac{1}{2!4!}s^{2}t^{4}=\frac{1}{6!}\cdot\left(-t^{6}-s^{6}+\frac{1}{2}(s+t)^{6}+\frac{1}{2}(s-t)^{6}\right).$
The powers of linear forms $\{t^{6},s^{6},(s+t)^{6},(s-t)^{6}\}$ correspond to the four points $\{\nu_{6}([0:1]),\nu_{6}([1:0]),\nu_{6}([1:1]),\nu_{6}([1:-1])\}$ on the standard rational normal curve.
Let $\pi_{p}:\mathbb{P}^{d}\dashrightarrow\mathbb{P}^{d-1}$ be the linear projection map away from a point $p\in\mathbb{P}^{d}$. The projected curve $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})$ can be classified in terms of the $\mathcal{C}_{d}$-rank of the point $p$. For example,
  • If $p$ lies on $\mathcal{C}_{d}$, then $\mathcal{C}$ is the curve of minimal degree.
  • If $p$ lies in $\mathcal{C}_{d}^{2}\setminus\mathcal{C}_{d}$, then $\mathcal{C}$ is an arithmetically Cohen-Macaulay(aCM) curve of almost minimal degree.
  • If $p$ lies in $\mathbb{P}^{d}\setminus\mathcal{C}_{d}^{2}$, then $\mathcal{C}$ is a smooth non-aCM curve of almost minimal degree.
In fact, the graded Betti numbers of $\mathcal{C}$ are determined by the $\mathcal{C}_{d}$-rank of $p$ for $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}^{2}$, see [Par07].
Through the identification between points in $\mathbb{P}^{d}$ and binary forms of degree $d$, we can regard the projection map $\pi_{p}$ as a $\mathbb{K}$-linear map $\pi_{p}:\mathbb{K}[s,t]_{d}\rightarrow((f_{p})^{\perp})_{d}$ on vector spaces. Indeed, for $f_{p}\in\mathbb{K}[s,t]_{d}$ associated with the point $p\in\mathbb{P}^{d}$, $((f_{p})^{\perp})_{d}$ forms a hyperplane orthogonal to $f_{p}$ with respect to the inner product $\circ$. Indeed, given $p$, let $\{p_{1},\dots,p_{d}\}$ be a basis for $p^{\perp}=\{p^{\prime}\in\mathbb{P}^{d}:p\circ p^{\prime}=0\}$ for any inner product $\circ$ on the space $\mathbb{P}^{d}$. Since the linear projection $\pi_{p}:\mathbb{P}^{d}\dashrightarrow\mathbb{P}^{d-1}$ is the map sending $q=\alpha_{0}p+\sum_{i=1}^{d}\alpha_{i}p_{i}$ to $\sum_{i=1}^{d}\alpha_{i}p_{i}$, the linear projection $\pi_{p}:\mathbb{K}[s,t]_{d}\rightarrow\mathbb{K}[s,t]_{d}$ sends the binary form $f_{q}=\alpha_{0}f_{p}+\sum_{i=1}^{d}\alpha_{i}f_{i}\in\mathbb{K}[s,t]_{d}$ to $\sum_{i=1}^{d}\alpha_{i}f_{i}$, where $\{f_{1},f_{2},\dots,f_{d}\}$ is the basis for $((f_{p})^{\perp})_{d}$.
Furthermore, by associating a point $p\in\mathbb{P}^{d}$ with a binary form $f_{p}\in\mathbb{K}[s,t]_{d}$ and applying apolarity to this form, we can obtain deeper insights into the projected curve $\mathcal{C}$. In detail, for a given point $p\in\mathbb{P}^{d}$, the apolar ideal is given by $(f_{p})^{\perp}=(g_{1},g_{2})\subset\mathbb{K}[s,t]$ with $\deg g_{1}\leq\deg g_{2}$. The $\mathcal{C}_{d}$-rank of the point $p$ equals to $\deg(g_{1})$ which also specifies the type of rational normal surface containing the curve $\pi_{p}(\mathcal{C}_{d})$.
2.3. Quadratic rank index of projective schemes
For a nondegenerate projective scheme $X\subset\operatorname{\mathbb{P}}^{r}$ defined by quadratic polynomials, we denote by $I(X)$ the homogeneous ideal in the polynomial ring $R=\mathbb{K}[z_{0},z_{1},\ldots,z_{r}]$ with standard grading on variables and let $\{Q_{0},\ldots,Q_{t}\}$ be a basis of $I(X)_{2}$. We say $X$ satisfies property $\operatorname{QR}(k)$ for an integer $k$ if $I(X)$ is generated by a collection of quadratic forms of rank at most $k$. We define the quadratic rank-index of a variety $X$ (defined by quadrics) as the minimal number $k$ such that $X$ satisfies property $\operatorname{QR}(k)$. In this section we discuss how to compute the rank index of $X$ given the basis $\{Q_{0},\ldots,Q_{t}\}$.
A general member $Q$ of $I(X)_{2}$ is of the form
$Q=x_{0}Q_{0}+\cdots+x_{t}Q_{t}\quad\mbox{for some}\quad x_{0},\ldots,x_{t}\in\mathbb{K}.$
We identify $\operatorname{\mathbb{P}}(I(X)_{2})$ with $\operatorname{\mathbb{P}}^{t}$ by sending $[Q]$ to $[x_{0}:\cdots:x_{t}]$. Now, let $M_{Q}$ be the $(r+1)\times(r+1)$ symmetric matrix associated with $Q$. Note that its entries are linear forms in $x_{0},\ldots,x_{t}$. For each $1\leq k\leq r$, we define the locus
$\Phi_{k}(X):=\{[Q]~{}|~{}Q\in I(X)_{2}-\{0\},~{}\mbox{rank}(Q)\leq k\}\subset\operatorname{\mathbb{P}}(I(X)_{2})=\operatorname{\mathbb{P}}^{t}$
of all quadratic polynomials of rank at most $k$ in the projective space $\operatorname{\mathbb{P}}^{t}$. Note that $\Phi_{k}(X)$ is a projective algebraic set in $\operatorname{\mathbb{P}}^{t}$ since it is the zero set of the ideal $I(k+1,M)$ of $\mathbb{K}[x_{0},\ldots,x_{t}]$ generated by all $(k+1)\times(k+1)$ minors of $M$. There is a descending filtration of $\operatorname{\mathbb{P}}^{t}$ associated to $X$:
$\emptyset=\Phi_{1}(X)=\Phi_{2}(X)\subset\Phi_{3}(X)\subset\Phi_{4}(X)\subset\cdots\subset\Phi_{r}(X)\subset\operatorname{\mathbb{P}}^{t}$
Note that any change of coordinates $T:\operatorname{\mathbb{P}}^{r}\rightarrow\operatorname{\mathbb{P}}^{r}$ induces a projective equivalence between $\Phi_{k}(X)$ and $\Phi_{k}(T(X))$.
Proposition 2.5.
[MP23, Proposition 2.1.] Keep the previous notations. Then
$\operatorname{rank-index}(X)=\min\{~{}s\mid\sqrt{I(s+1,M_{Q})}\quad\mbox{has no nonzero linear form.}\}$. }
2.4. Rank $3$ generators
We review the $Q$-map on a given decomposition of the vector bundle over varieties, which produces rank 3 quadratic generators, introduced in [HLMP21]. Let $L=\mathcal{O}_{X}(1)$ and suppose that $X\subset\operatorname{\mathbb{P}}^{r}$ is the linearly normal embedding of $X$ by $L$. Thus there is a natural isomorphism
$f:H^{0}(X,L)\rightarrow R_{1}$
of $\mathbb{K}$-vector spaces. Now, assume that $L$ is decomposed as $L=A^{2}\otimes B$ for some line bundles $A$ and $B$ on $X$ such that $h^{0}(X,A)\geq 2$ and $h^{0}(X,B)\geq 1$. Define the map
$Q_{A,B}:H^{0}(X,A)\times H^{0}(X,A)\times H^{0}(X,B)\rightarrow I(X)_{2}$
by
$Q_{A,B}(s,t,h)=f(s\otimes s\otimes h)f(t\otimes t\otimes h)-f(s\otimes t\otimes h)^{2}.$
This map is well-defined and any rank $3$ quadratic generators of $I(X)$ with the line bundle $L=A^{2}\otimes B$ on $X$ lies on the image of the $Q_{A,B}$ map.
3. Generators of projected curves
Let $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})$ for $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}$ be the projected curve of degree $d$ in $\mathbb{P}^{d-1}$. In this section, we demonstrate the existence of a parameterized curve that is projectively equivalent to $\mathcal{C}$, and we show that $\mathcal{C}$ satisfies property $\operatorname{QR}(4)$. We note that, since the projected curve is not linearly normal, determining whether the projected curves satisfy property $\operatorname{QR}(k)$ is not straightforward.
Theorem 3.1.
Let $f_{p}\in\mathbb{K}[s,t]_{d}$ be the binary form corresponding to $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}$. Let $g_{1}$ and $g_{2}$ be binary forms that generate the ideal $(f_{p})^{\perp}$, and let $d_{i}=\deg g_{i}$ for $i=1,2$. Then, the curve $\mathcal{C}$ is projectively equivalent to a curve with the following parametrization:
$\{[g_{2}(\alpha,\beta)\alpha^{d_{1}-2}:\dots:g_{2}(\alpha,\beta)\beta^{d_{1}-2}:g_{1}(\alpha,\beta)\alpha^{d_{2}-2}:\dots:g_{1}(\alpha,\beta)\beta^{d_{2}-2}]\mid[\alpha:\beta]\in\mathbb{P}^{1}\}.$
Proof.
Given $p\in\mathbb{P}^{d}$, we identify $p$ with the binary form $f_{p}\in\mathbb{K}[s,t]_{d}$ and let $\tilde{p}\in\mathbb{K}^{d+1}$ be an affine representative of $p$. We view $\tilde{p}$ as a vector whose initial point is the origin and endpoint is $\tilde{p}$. Note that the orthogonal complement of the vector $\tilde{p}$ corresponds to the vector space $((f_{p})^{\perp})_{d}\subset\mathbb{K}[S,T]_{d}$, and
$((f_{p})^{\perp})_{d}$ $=g_{1}\mathbb{K}[S,T]_{d-d_{1}}+g_{2}\mathbb{K}[S,T]_{d-d_{2}}$
$=\langle g_{1}S^{d-d_{1}},\dots,g_{1}T^{d-d_{1}},g_{2}S^{d-d_{2}},\dots,g_{2}T^{d-d_{2}}\rangle,$
where $d_{i}=\deg g_{i}$ for $i=1,2$ with $2\leq d_{1}\leq d_{2}$.
Let $\langle\tilde{p}\rangle^{\perp}=W=\{w\mid w\circ\tilde{p}=0\}$. Then the set $\{\tilde{p},w_{1},\dots,w_{d}\}$ forms a basis of $\mathbb{K}^{d+1}$ for any independent set $\{w_{1},\dots,w_{d}\}$ in $W$. Let $\{v_{0},v_{1},\dots,v_{d}\}$ be the dual basis to $\{\tilde{p},w_{1},\dots,w_{d}\}$ in $(\mathbb{K}^{d+1})^{*}$. Note that $v_{i}\circ\tilde{p}=0$ for $1\leq i\leq d$ and $v_{i}\circ w_{j}=\delta_{ij}$ for $0\leq i\leq d$ and $1\leq j\leq d$.
We choose $\{v_{1},v_{2},\dots,v_{d}\}\subset(\mathbb{K}^{d+1})^{*}$ such that each vector in $\{v_{1},v_{2},\dots,v_{d}\}$ corresponds to the binary forms $\{g_{2}S^{d-d_{2}},\dots,g_{2}T^{d-d_{2}},g_{1}S^{d-d_{1}},\dots,g_{1}T^{d-d_{1}}\}\subset\mathbb{K}[S,T]_{d}$, and take $\{w_{1},w_{2},\dots,w_{d}\}\subset\mathbb{K}^{d+1}$ such that $\{w_{1},w_{2},\dots,w_{d}\}$ is dual to $\{v_{1},v_{2},\dots,v_{d}\}$.
Since $\{\tilde{p},w_{1},\dots,w_{d}\}$ forms a basis of $\mathbb{K}^{d+1}$, for any $\tilde{q}\in\mathbb{K}^{d+1}$, we have $\tilde{q}=a_{0}\tilde{p}+\sum_{i=1}^{d}(v_{i}\circ\tilde{q})w_{i}$. Therefore, for $q\in\mathbb{P}^{d}$ whose affine representative is $\tilde{q}\in\mathbb{K}^{d+1}$,
$f_{q}=a_{0}f_{p}+(g_{2}S^{d-d_{2}}\circ f_{q})f_{w_{1}}+\cdots+(g_{1}T^{d-d_{1}}\circ f_{q})f_{w_{d}}.$
Then $\pi_{p}:\mathbb{P}^{d}\dashrightarrow\mathbb{P}^{d-1}$ sends $q$ to $[g_{2}S^{d-d_{2}}\circ f_{q}:\cdots:g_{1}T^{d-d_{1}}\circ f_{q}]$.
For $q\in\mathcal{C}_{d}=\{[\alpha^{d}:\alpha^{d-1}\beta:\cdots:\beta^{d}]\mid[\alpha:\beta]\in\mathbb{P}^{1}\}$, $q$ is identified with
$f_{q}=(\alpha s+\beta t)^{d}=\sum_{i=0}^{d}\binom{d}{i}\alpha^{d-i}\beta^{i}s^{d-i}t^{i}.$
Then the inner products between the basis elements of $((f_{p})^{\perp})_{d}$ and $f_{q}$ are the evaluations of these basis elements at $(\alpha,\beta)$ by Lemma 2.3. Thus, $\pi_{p}(\mathcal{C})$ is projectively equivalent to the curve
$\{[g_{2}(\alpha,\beta)\alpha^{d-d_{2}}:\dots:g_{2}(\alpha,\beta)\beta^{d-d_{2}}:g_{1}(\alpha,\beta)\alpha^{d-d_{1}}:\dots:g_{1}(\alpha,\beta)\beta^{d-d_{1}}]\mid[\alpha:\beta]\in\mathbb{P}^{1}\}.$
Due to Theorem 3.1, we can immediately see that $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})$ is contained in a rational surface scroll $S=S(d-d_{2},d-d_{1})=S(d_{1}-2,d_{2}-2)\subset\mathbb{P}^{d-1}$ as a divisor. This provides a constructive proof of the existence of a surface scroll containing the projected curve. Moreover, we can prove that the ideals of the projected curves can be generated by quadratic forms of rank at most four.
Corollary 3.2.
For $d\geq 4$, any projected curve $\pi_{p}(\mathcal{C}_{d})\subset\mathbb{P}^{d-1}$ for $p\in\mathbb{P}^{d}$ satisfies property $\operatorname{QR}(4)$.
Proof.
For points $p\in\mathcal{C}_{d}^{2}$, the projected curve $\pi_{p}(\mathcal{C}_{d})$ satisfies property $\operatorname{QR}(3)$, as shown in [Par22]. Therefore, we assume that $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}^{2}$. Let $f_{p}\in\mathbb{K}[s,t]_{d}$ be the binary form corresponding to $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}^{2}$ with $d\geq 4$. By Theorem 3.1, the projected curve $\pi_{p}(\mathcal{C}_{d})$ is projectively equivalent to the curve
$\mathcal{C}=\left\{[g_{2}(\alpha,\beta)\alpha^{d_{1}-2}:\dots:g_{2}(\alpha,\beta)\beta^{d_{1}-2}:g_{1}(\alpha,\beta)\alpha^{d_{2}-2}:\dots:g_{1}(\alpha,\beta)\beta^{d_{2}-2}]\mid[\alpha:\beta]\in\mathbb{P}^{1}\right\}$
where $(f_{p})^{\perp}=(g_{1},g_{2})$ with $d_{1}+d_{2}=d+2$.
Therefore, the curve $\mathcal{C}$ is contained in
$\mathcal{S}$ =$=\left\{[\alpha^{d_{1}-2}\gamma:\dots:\beta^{d_{1}-2}\gamma:\alpha^{d_{2}-2}\eta:\dots:\beta^{d_{2}-2}\eta]\mid[\alpha:\beta],[\gamma:\eta]\in\mathbb{P}^{1}\right\}$
$=\mathcal{S}(d_{1}-2,d_{2}-2)$
and $\mathcal{C}$ is obtained by the section defined by $g_{1}x-g_{2}y=0$, where $x$ and $y$ are parameters of $\mathcal{S}$. Note that $\mathcal{C}$ is a divisor in $\mathcal{S}$ as the class $H+2F$ where $H$ and $F$ are respectively a hyperplane section and a ruling of $\mathcal{S}$.
Therefore, the ideal $I(\pi_{p}(\mathcal{C}_{d}))_{2}$ is generated by the generators of $I(\mathcal{S})$ and quadratic forms corresponding to a basis of $(H-2F)(g_{1}x-g_{2}y)$. Note that the surface scroll $\mathcal{S}$ satisfies property $\operatorname{QR}(4)$ since it is obtained by the $2\times 2$ minors of a $2\times(d-2)$ matrix consisting of linear forms in the coordinate ring of $\mathcal{S}$.
Since $H-2F$ is spanned by $\{S^{d_{1}-j-4}T^{j}x\mid 0\leq j\leq d_{1}-4\}\cup\{S^{d_{2}-j-4}T^{j}y\mid 0\leq j\leq d_{2}-4\}$, $I(\pi_{p}(\mathcal{C}_{d}))_{2}/I(\mathcal{S})_{2}$ is generated by quadratic forms on $\mathcal{S}$ corresponding to
$\{S^{d_{1}-j-4}T^{j}g_{1}(S,T)x^{2}-S^{d_{1}-j-4}T^{j}g_{2}(S,T)xy\mid 0\leq j\leq d_{1}-4\}$
[…]
There exist two quadratic forms $q_{i}(S,T),i=1,2$ such that $q_{i}\mid g_{i}$ for $i=1,2$, since $g_{1}$ and $g_{2}$ are binary forms. Let $L_{1,j}$ be the linear form on $\mathcal{S}$ corresponding to $S^{d_{1}-j-4}T^{j}q_{1}x$, $L_{2}$ be the linear form on $\mathcal{S}$ corresponding to $(g_{1}/q_{1})x$, $L_{3,j}$ be the linear form on $\mathcal{S}$ corresponding to $S^{d_{1}-j-4}T^{j}q_{2}x$, and $L_{4}$ be the linear form on $\mathcal{S}$ corresponding to $(g_{2}/q_{2})y$. Then the quadratic generators corresponding to $\{S^{d_{1}-j-4}T^{j}g_{1}x^{2}-S^{d_{1}-j-4}T^{j}g_{2}xy\mid 0\leq j\leq d_{1}-4\}$ are
$\{L_{1,j}L_{2}-L_{3,j}L_{4}\mid 0\leq j\leq d_{1}-4\},$
which is a set of quadratic forms whose ranks are at most four.
Similarly, let $L_{5,j}$ be the linear form on $\mathcal{S}$ corresponding to $S^{d_{2}-j-4}T^{j}q_{1}y$ and $L_{6,j}$ be the linear form on $\mathcal{S}$ corresponding to $S^{d_{2}-j-4}T^{j}q_{2}y$. Then the quadratic generators corresponding to $\{S^{d_{2}-j-4}T^{j}g_{1}xy-S^{d_{2}-j-4}T^{j}g_{2}y^{2}\mid 0\leq j\leq d_{2}-4\}$ are
$\{L_{5,j}L_{2}-L_{6,j}L_{4}\mid 0\leq j\leq d_{2}-4\},$
which is a set of quadratic forms whose ranks are at most four. Thus, any quadratic generators of the ideal of $\mathcal{C}$ are generated by quadratic forms of rank at most four, implying that the curve $\pi_{p}(\mathcal{C}_{d})$ satisfies property $\operatorname{QR}(4)$. ∎
Example 3.3 (Monomial projections).
Let $e_{c}$ be the $c$-th standard basis for $\mathbb{P}^{d}$ with $c\geq 1$ and $d\geq 4$ Then the corresponding binary form is $f_{e_{c}}=s^{d-c}t^{c}\in\mathbb{K}[s,t]_{d}$ and $f^{\perp}=(S^{d-c+1},T^{c+1})\subset\mathbb{K}[S,T]$. Therefore, by Theorem 3.1, the parametrization of the curve $\mathcal{C}=\pi_{e_{c}}(\mathcal{C}_{d})$ is
$\{[\alpha^{d}:\alpha^{d-1}\beta:\cdots:\alpha^{d-c+1}\beta^{c-1}:\alpha^{d-c-1}\beta^{c+1}:\cdots:\beta^{d}]\mid[\alpha:\beta]\in\mathbb{P}^{1}\}.$
Thus, $\mathcal{C}$ is a divisor of
$\mathcal{S}$ =$=\{[\alpha^{c-1}\gamma:\alpha^{c-2}\beta\gamma:\cdots:\beta^{c-1}\gamma:\alpha^{d-c-1}\eta:\alpha^{d-c-2}\beta\eta:\cdots:\beta^{d-c-1}\eta]|[\alpha:\beta],[\gamma:\eta]\in\mathbb{P}^{1}\}$
$(=\mathcal{S}(c-1,d-c-1))$
and $\mathcal{C}$ is a section defined by $xT^{c+1}-yS^{d-c+1}$ in $\mathcal{S}$. Since $H-2F$ is spanned by $\{S^{c-j-3}T^{j}x\mid 0\leq j\leq c-3\}\cup\{S^{d-c-j-3}T^{j}y\mid 0\leq j\leq d-c-3\}$, $I(\mathcal{C})_{2}/I(\mathcal{S})_{2}$ is generated by quadratic forms on $\mathcal{S}$ corresponding to
$\{S^{c-j-3}T^{c+j+1}x^{2}-S^{d-j-2}T^{j}xy\mid 0\leq j\leq c-3\}$
[…]
Therefore,
$I(\mathcal{S})=\left(2\times 2\text{ minors of }\begin{bmatrix}z_{0}&z_{1}&\cdots&z_{c-2}&z_{c}&z_{c+1}&\cdots&z_{d-2}\\ z_{1}&z_{2}&\cdots&z_{c-1}&z_{c+1}&z_{c+2}&\cdots&z_{d-1}\end{bmatrix}\right)$
and $I(\mathcal{C})$ is generated by the generators of $I(\mathcal{S})$ and
$\{z_{j}z_{c-1}-z_{j}z_{c}\mid 0\leq j\leq c-3\}\cup\{z_{c-1}z_{c+j+2}-z_{c}z_{c+j}\mid 0\leq j\leq d-c-3\}.$
Since
$I(\pi_{e_{c}}(\mathcal{C}_{d}))=I(\mathcal{C}_{d})\cap\mathbb{K}[z_{0},\dots,z_{c-1},z_{c+1},\dots,z_{d}],$
this gives a constructive proof of the result from [ES96] which states that the elimination of a variable from the ideal of the toric variety $\mathcal{C}_{d}$ results in an ideal generated by binomials.
4. Monomial projections
In Corollary 3.2, we have seen that the projected curves satisfy property $\operatorname{QR}(4)$. However, we will demonstrate that the projected curves also satisfy property $\operatorname{QR}(3)$ in certain cases.
Let $\mathcal{C}_{d,c}=\pi_{e_{c}}(\mathcal{C}_{d})\subset\mathbb{P}^{d-1}$ be the projected curve where $e_{c}$ for $0\leq c\leq d$ is the standard basis of $\mathbb{P}^{d}$. For some $c$, we can immediately see that property $\operatorname{QR}(3)$ is satisfied:
  1. (1)
    If $c=0\text{ or }d$, then $\mathcal{C}_{d,c}$ is a rational normal curve. Therefore, it satisfies property $\operatorname{QR}(3)$ (see [HLMP21]).
  2. (2)
    If $c=1\text{ or }d-1$, then $\mathcal{C}_{d,c}$ is a del Pezzo curve. Therefore, it satisfies property $\operatorname{QR}(3)$ (see [Par22]).
Thus, we assume $22\leq c\leq d-$ 2 and $d\geq 6$.
We distinguish between the cases $c=2\text{ or }d-2$ and $33\leq c\leq d-$ 3 because, in the case $c=2\text{ or }d-2$, $\mathcal{C}_{d,c}$ is not generated by quadrics, whereas for $33\leq c\leq d-$ 3, the curve $\mathcal{C}_{d,c}$ is generated by quadrics. For $\mathcal{C}_{d,2}$ (or $\mathcal{C}_{d,d-2}$), we instead consider the scheme $X_{d}=\mathcal{C}_{d,2}\cup\mathbb{L}$ where $\mathbb{L}$ is the trisecant line to $\mathcal{C}_{d,2}$. Note that $X_{d}$ is the scheme generated by quadrics on $\mathcal{C}_{d,2}$.
As we aim to determine whether the curves $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})\subset\mathbb{P}^{d-1}$ satisfy property $\operatorname{QR}(3)$ and to provide additional numerical information beyond simply determining if the property holds, we introduce the following notation:
Notation and Remark 4.1.
There is a combinatorial way to find a generating set of $I(\mathcal{C}_{d,c})$. Suppose
$I(\mathcal{C}_{d})=\left(2\times 2\text{ minors of }\begin{bmatrix}x_{0}&x_{1}&\cdots&x_{d-1}\\ x_{1}&x_{2}&\cdots&x_{d}\end{bmatrix}\right)\subset\mathbb{K}[x_{0},x_{1},\dots,x_{d}],$
and define the index of a quadratic monomial $x_{i}x_{j}$ by $i+j$. Observe that
  1. (1)
    differences of any pairs $(x_{i}x_{j},x_{k}x_{l})$ of quadratic monomials with the same index are generators of $I(\mathcal{C}_{d})$ and
  2. (2)
    the generators of $I(\mathcal{C}_{d})_{2}$ that do not involve $x_{c}$ can be generators of $I(\mathcal{C}_{d,c})_{2}$.
Therefore, the differences between any pairs of quadratic monomials $(x_{i}x_{j},x_{k}x_{l})$ with same index $i+j=k+l$ for $i,j,k,l\neq c$ are generators of $I(\mathcal{C}_{d,c})_{2}$. Moreover, for each integer $0\leq t\leq 2d$, we can obtain #{quadratic monomials whose index is $\#\{\text{quadratic monomials whose index is }t\}-1$ linearly independent quadratic forms for each index.
Among the list of monomials, there are $d+1$ quadratic monomials whose support contains $x_{c}$. For $11\leq c\leq d-$ 1, we lose exactly $d+1$ quadratic generators when eliminating $x_{c}$ from $I(\mathcal{C}_{d})$. Therefore, we have $\binom{d}{2}-(d+1)=\binom{d-1}{2}-2$ linearly independent quadratic generators in $I(\mathcal{C}_{d,c})_{2}$. These generators span $I(\mathcal{C}_{d,c})_{2}$ because $\dim_{\mathbb{K}}I(\mathcal{C}_{d,c})_{2}=\binom{d-2}{2}+d-4=\binom{d-1}{2}-2$ where $I(\mathcal{C}_{d,c})_{2}=I(\mathcal{S})_{2}+(I(\mathcal{C}_{d,c})_{2}/I(\mathcal{S})_{2})$.
Let $X$ be a projective scheme defined by quadratic forms and $I$ be the saturated ideal defining $X$.
  1. (1)
    For quadratic forms $q$ and $q^{\prime}$ in $I$, we write
    $q\sim_{I}q^{\prime}$
    if $q-q^{\prime}$ can be expressed as a sum of rank 3 generators of $I$. In particular,
    1. (i)
      $\sim_{I}$ is an equivalence relation and
    2. (ii)
      if $q\sim_{I}0$, then $q$ is spanned by rank 3 quadratic forms in $I$.
  2. (2)
    For $0\leq c\leq d$, let $I_{d,c}$ denote the ideal of $\mathcal{C}_{d,c}=\pi_{e_{c}}(\mathcal{C}_{d})$. We write
    $q\sim_{(d,c)}q^{\prime}$
    for quadratic forms $q$ and $q^{\prime}$ in $I_{d,c}$ if $q-q^{\prime}$ can be expressed as a sum of rank 3 generators in $(I_{d,c})_{2}$.
  3. (3)
    $\delta(X,t)$ denotes the number of independent quadratic generators of $I$ whose rank is at most $t$.
We first claim that $\mathcal{C}_{d,c}$ for $d\geq 6$ and $33\leq c\leq d-$ 3 satisfies property $\operatorname{QR}(3)$. To show this claim, we begin by proving a series of computational lemmas.
Lemma 4.2.
Let $i$, $j$, $k$, and $l$ be integers such that $0\leq i<j\leq d$, $i<j-k$, $i+j+k\equiv 0\pmod{2}$, and $c\neq i,i+k,j-k,j,\frac{i+j-k}{2},\frac{i+j+k}{2}$. Then,
$x_{i}x_{j}+x_{i+k}x_{j-k}-2x_{\frac{i+j-k}{2}}x_{\frac{i+j+k}{2}}\sim_{(d,c)}0.$
Proof.
We fix the decomposition of the line bundle $\mathcal{O}_{\mathbb{P}^{1}}(d)=A^{\otimes 2}\times B$ where
$A=\mathcal{O}_{\mathbb{P}^{1}}\left(\frac{j-i-k}{2}\right)\text{ and }B=\mathcal{O}_{\mathbb{P}^{1}}(d-j+i+k).$
Then
$Q_{A,B}(S^{\frac{j-i-k}{2}},T^{\frac{j-i-k}{2}},S^{d-j}T^{i}(S^{k}+T^{k}))$
$=$ $(x_{i}+x_{i+k})(x_{j-k}+x_{j})-(x_{\frac{i+j-k}{2}}+x_{\frac{i+j+k}{2}})^{2}$
$=$ $(x_{i}x_{j-k}-x_{\frac{i+j-k}{2}}^{2})+(x_{i+k}x_{j}-x_{i+j+k}^{2})+x_{i}x_{j}+x_{i+k}x_{j-k}-2x_{\frac{i+j-k}{2}}x_{\frac{i+j+k}{2}}.$
Since $(x_{i}x_{j-k}-x_{\frac{i+j-k}{2}}^{2})$ and $(x_{i+k}x_{j}-x_{i+j+k}^{2})$ are rank $3$ generators in $I_{d,c}$, we have $x_{i}x_{j}+x_{i+k}x_{j-k}-2x_{\frac{i+j-k}{2}}x_{\frac{i+j+k}{2}}\sim_{(d,c)}0$. ∎
Lemma 4.3.
For each $60\leq i\leq d-$ 6, the two quadratic forms $x_{i}x_{i+5}-x_{i+1}x_{i+4}$ and $x_{i+1}x_{i+6}-x_{i+2}x_{i+5}$ satisfy
$(x_{i}x_{i+5}-x_{i+1}x_{i+4})\sim_{(d,i+3)}0\text{ and }(x_{i+1}x_{i+6}-x_{i+2}x_{i+5})\sim_{(d,i+3)}0.$
Proof.
For a decomposition $\mathcal{O}_{\mathbb{P}^{1}}(d)=\mathcal{O}_{\mathbb{P}^{1}}(2)^{\otimes 2}\otimes\mathcal{O}_{\mathbb{P}^{1}}(d-4)$ of the line bundle,
$\;Q_{\mathcal{O}_{\mathbb{P}^{1}}(2),\mathcal{O}_{\mathbb{P}^{1}}(d-4)}(S^{2},\;ST+aT^{2},\;S^{d-i-4}T^{i}-2aS^{d-i-5}T^{i+1}+2a^{2}S^{d-i-6}T^{i+2})$
$=$ $\;(x_{i}x_{i+2}-x_{i+1}^{2})+a^{2}(-x_{i}x_{i+4}+x_{i+2}^{2})+2a^{3}(x_{i}x_{i+5}-x_{i+1}x_{i+4})$
$+2a^{4}(x_{i}x_{i+6}-2x_{i+1}x_{i+5}+x_{i+2}x_{i+4})+4a^{5}(-x_{i+1}x_{i+6}+x_{i+2}x_{i+5})$
$+4a^{6}(x_{i+2}x_{i+6}-x_{i+4}^{2})$
for $a\in\mathbb{K}$. Since $x_{i}x_{i+6}+x_{i+2}x_{i+4}-2x_{i+1}x_{i+5}\sim_{(d,i+3)}(x_{i}+x_{i+4})(x_{i+2}+x_{i+6})-(x_{i+1}+x_{i+5})^{2}$ and the forms $x_{i}x_{i+2}-x_{i+1}^{2}$, $x_{i}x_{i+4}-x_{i+2}^{2}$, $x_{i+2}x_{i+6}-x_{i+4}^{2}$, and $(x_{i}+x_{i+4})(x_{i+2}+x_{i+6})-(x_{i+1}+x_{i+5})^{2}$ are rank 3 quadratic generators in $I_{d,i+3}$, it follows that
$02a^{3}(x_{i}x_{i+5}-x_{i+1}x_{i+4})-4a^{5}(x_{i+1}x_{i+6}-x_{i+2}x_{i+5})\sim_{(d,i+3)}$ 0
for any $a\in\mathbb{K}$.
Thus $x_{i}x_{i+5}-x_{i+1}x_{i+4}\sim_{(d,i+3)}0$ and $x_{i+1}x_{i+6}-x_{i+2}x_{i+5}\sim_{(d,i+3)}0$. ∎
Lemma 4.4.
For each $50\leq i\leq d-$ 5, the two quadratic forms $x_{i}x_{i+5}-x_{i+2}x_{i+3}$ and $x_{i+2}x_{i+5}-x_{i+3}x_{i+4}$ satisfy
$(x_{i}x_{i+5}-x_{i+2}x_{i+3})\sim_{(d,i+1)}0\text{ and }(x_{i+2}x_{i+5}-x_{i+3}x_{i+4})\sim_{(d,i+1)}0.$
Proof.
For a decomposition $\mathcal{O}_{\mathbb{P}^{1}}(d)=\mathcal{O}_{\mathbb{P}^{1}}(2)^{\otimes 2}\otimes\mathcal{O}_{\mathbb{P}^{1}}(d-4)$ of the line bundle,
$Q_{\mathcal{O}_{\mathbb{P}^{1}}(2),\mathcal{O}_{\mathbb{P}^{1}}(d-4)}(S^{2}+aST,\;T^{2},\;S^{d-i-4}T^{i}-2aS^{d-i-5}T^{i+1})$
=$\quad$= $(x_{i}-3a^{2}x_{i+2}-2a^{3}x_{i+3})(x_{i+4}-2ax_{i+5})-(x_{i+2}-ax_{i+3}-2a^{2}x_{i+4})^{2}$
=$\quad$= $(x_{i}x_{i+4}-x_{i+2}^{2})+2a(-x_{i}x_{i+5}+x_{i+2}x_{i+3})+a^{2}(x_{i+2}x_{i+4}-x_{i+3}^{2})$
$+6a^{3}(x_{i+2}x_{i+5}-x_{i+3}x_{i+4})+4a^{4}(x_{i+3}x_{i+5}-x_{i+4}^{2})$
Since $x_{i}x_{i+4}-x_{i+2}^{2}$, $x_{i+2}x_{i+4}-x_{i+3}^{2}$, and $x_{i+3}x_{i+5}-x_{i+4}^{2}$ are rank 3 quadratic generators in $I_{d,i+3}$,
$02a(-x_{i}x_{i+5}+x_{i+2}x_{i+3})+6a^{3}(x_{i+2}x_{i+5}-x_{i+3}x_{i+4})\sim_{(d,i+1)}$ 0
for any $a\in\mathbb{K}$.
Hence $x_{i}x_{i+5}-x_{i+2}x_{i+3}\sim_{(d,i+1)}0$ and $x_{i+2}x_{i+5}-x_{i+3}x_{i+4}\sim_{(d,i+1)}0$. ∎
Theorem 4.5.
For $d\geq 6$ and $33\leq c\leq d-$ 3, the ideal of $\mathcal{C}_{d,c}$ is generated by quadratic forms and satisfies property $\operatorname{QR}(3)$.
Proof.
Since $I_{d,c}\cong I_{d,d-c}$ for any $0\leq c\leq d$, we assume that $3\leq c\leq\lfloor\frac{d}{2}\rfloor$. We recall that the quadratic generators of $I_{d,c}$ are given by binomials of the form $x_{i}x_{j}-x_{k}x_{l}$ where $i+j=k+l$ and $i,j,k,l\neq c$. There are four initial cases for using induction:
$(d,c)=(6,3),(7,3),(8,3),\text{ and }(8,4).$
Initial Case 1 : $d=6,c=3$
index monomials
$2$ $x_{0}x_{2},x_{1}^{2}$
$3$ $x_{1}x_{2}$
$4$ $x_{0}x_{4},x_{2}^{2}$
$5$ $x_{0}x_{5},x_{1}x_{4}$
$6$ $x_{0}x_{6},x_{1}x_{5},x_{2}x_{4}$
$7$ $x_{1}x_{6},x_{2}x_{5}$
$8$ $x_{2}x_{6},x_{4}^{2}$
$9$ $x_{4}x_{5}$
$10$ $x_{4}x_{6},x_{5}^{2}$
Table 1. List of monomials given indices.
The index $2,4,8,10$ generators are $x_{0}x_{2}-x_{1}^{2},x_{0}x_{4}-x_{2}^{2},x_{2}x_{6}-x_{4}^{2},x_{4}x_{6}-x_{5}^{2}$ and they are already rank $3$ quadratic generators. The remaining generators are index $5,6,7$ generators. By Lemma 4.3 with $i=0$, index $5$ generator $x_{0}x_{5}-x_{1}x_{4}\sim_{(6,3)}0$ and index $7$ generator $x_{1}x_{6}-x_{2}x_{5}\sim_{(6,3)}0$. By Lemma 4.2, $x_{0}x_{6}+x_{2}x_{4}-2x_{2}x_{4}=x_{0}x_{6}-x_{2}x_{4}\sim_{(6,3)}0$ and $x_{0}x_{6}+x_{2}x_{4}-2x_{1}x_{5}=(x_{0}x_{6}-x_{2}x_{4})+2(x_{2}x_{4}-x_{1}x_{5})\sim_{(6,3)}0$. Hence $I_{6,3}$ satisfies $\operatorname{QR}(3)$ property.
Initial Case 2 : $d=7,c=3$.
Since $I_{7,3}$ contains the same generators as $I_{6,3}$, we only need to examine the new generators that include $x_{7}$ as supports. The generators of index $\geq 10$ are given by the $2\times 2$ minors of the matrix
$\begin{pmatrix}x_{4}&x_{5}&x_{6}\\ x_{5}&x_{6}&x_{7}\end{pmatrix}$
These minors correspond to the generators of a rational normal curve and are spanned by rank $3$ quadratic forms.
The remaining monomials are $x_{0}x_{7}$, $x_{1}x_{7}$, and $x_{2}x_{7}$. We need to find quadratic generators containing these monomials that can be expressed as a sum of rank $3$ quadrics without using $x_{3}$. By Lemma 4.2, we have $x_{0}x_{7}+x_{2}x_{5}-2x_{1}x_{6}\sim_{(7,3)}0$. The generator $x_{1}x_{7}-x_{4}^{2}$ is already a rank $3$ quadratic form. In addition, we obtain $x_{2}x_{7}-x_{4}x_{5}\sim_{(7,3)}0$ by applying Lemma 4.4 with $i=2$. Thus, we conclude that $I_{7,3}$ satisfies property $\operatorname{QR}(3)$.
Initial Case 3 : $d=8,c=3$
As in the previous case, since $I_{8,3}$ contains the same generators as $I_{7,3}$, we only need to check the new generators that involve $x_{8}$. The generators of index $\geq 11$ are given by the $2\times 2$ minors of the matrix
$\begin{pmatrix}x_{4}&x_{5}&x_{6}&x_{7}\\ x_{5}&x_{6}&x_{7}&x_{8}\end{pmatrix}$
These minors are generators of a rational normal curve, which already satisfies property $\operatorname{QR}(3)$.
The remaining monomials are $x_{0}x_{8}$, $x_{1}x_{8}$, and $x_{2}x_{8}$. The binomials $x_{0}x_{8}-x_{4}^{2}$ and $x_{2}x_{8}-x_{5}^{2}$ are already rank $3$ generators. By Lemma 4.2, we have $x_{1}x_{8}+x_{2}x_{7}-2x_{4}x_{5}\sim_{(8,3)}0$. Thus, $I_{8,3}$ satisfies property $\operatorname{QR}(3)$.
Initial Case 4 : $d=8,c=4$
Since $I_{7,4}\cong I_{7,3}$, we only need to check the new generators that involve $x_{8}$. The generators of index $\geq 12$ are given by the $2\times 2$ minors of the matrix
$\begin{pmatrix}x_{5}&x_{6}&x_{7}\\ x_{6}&x_{7}&x_{8}\end{pmatrix}$
These minors correspond to generators of a rational normal curve, which already satisfies property $\operatorname{QR}(3)$.
The remaining monomials are $x_{0}x_{8}$, $x_{1}x_{8}$, $x_{2}x_{8}$, and $x_{3}x_{8}$. By Lemma 4.2, we have $x_{0}x_{8}+x_{2}x_{6}-2x_{3}x_{5}\sim_{(8,4)}0$ and $x_{1}x_{8}+x_{3}x_{6}-2x_{2}x_{7}\sim_{(8,4)}0$. In addition, the binomial $x_{2}x_{8}-x_{5}^{2}$ is already a rank $3$ quadratic form. Moreover, we obtain $x_{3}x_{8}-x_{5}x_{6}\sim_{(8,4)}0$ by Lemma 4.4 with $i=3$. Hence, $I_{8,4}$ satisfies property $\operatorname{QR}(3)$.
For the inductive hypothesis, assume that the ideal $I_{d-1,c}$ satisfies property $\operatorname{QR}(3)$ for $d\geq 9$ and $c$ such that $3\leq c\leq\lfloor\frac{d}{2}\rfloor$. Since $I_{d,c}$ contains the generators of $I_{d-1,c}$, we only need to find the rank $3$ quadratic forms that involve $x_{d}$. The generators of index $\geq d+c$ are those of a rational normal curve, so the remaining task is to check the quadratic generators with indices between $d$ and $d+c-1$. We will consider the cases based on the parity of $d$ and $c$.
Case 1-1 : $d$ is odd and $3\leq c<\frac{d-1}{2}$.
index quadratic forms involving $x_{d}$
$d$ $x_{0}x_{d}+x_{1}x_{d-1}-2x_{\frac{d-1}{2}}x_{\frac{d+1}{2}}$
$d+1$ $x_{1}x_{d}-x_{\frac{d+1}{2}}^{2}$
$d+2$ $x_{2}x_{d}+x_{3}x_{d-1}-2x_{\frac{d+1}{2}}x_{\frac{d+3}{2}}$
$d+2i$ $x_{2i}x_{d}+x_{2i+1}x_{d-1}-2x_{i+\frac{d-1}{2}}x_{i+\frac{d+1}{2}}$
$d+2i+1$ $x_{2i+1}x_{d}-x_{i+\frac{d+1}{2}}^{2}$
$d+c-2$$\left\{\begin{array}[]{lr}x_{c-2}x_{d}+x_{c-1}x_{d-1}-2x_{\frac{d+c-3}{2}}x_{\frac{d+c-1}{2}},&\text{for even }c\\ x_{c-2}x_{d}-x_{\frac{d+c-2}{2}}^{2},&\text{for odd }c\\ \end{array}\right.$
Table 2. Quadratic generators in $I_{d,c}$ for odd $d$ and $3\leq c<\frac{d-1}{2}$.
Since $\frac{d-1}{2}>c$, all of binomals in Table 2 can be expressed as sums of rank 3 quadrics by Lemma 4.2.
If $d+c-1$ is an even number, the quadratic form $x_{c-1}x_{d}-x_{\frac{d+c-1}{2}}^{2}$ is already a rank 3 quadratic form. If $d+c-1$ is odd, we use the following binomial:
$x_{c-1}x_{d}+x_{c+2}x_{d-3}-2x_{\frac{d+c-4}{2}}x_{\frac{d+c+2}{2}}\sim_{(d,c)}0.$
The relation holds because $c<d-4$ implied by $c\leq\frac{d-1}{2}$ and $d\geq 9$.
Case 1-2 : $d$ is odd and $c=\frac{d-1}{2}$. The only issue arises at index $d$ because $x_{\frac{d-1}{2}}=x_{c}$. To resolve this, we use the following binomial:
$x_{0}x_{d}+x_{3}x_{d-3}-2x_{\frac{d-3}{2}}x_{\frac{d+3}{2}}\sim_{(d,c)}0.$
Case 2-1 : $d$ is even and $3\leq c<\frac{d}{2}$.
index quadrics containing $x_{d}$
$d$ $x_{0}x_{d}-x_{\frac{d}{2}}^{2}$
$d+1$ $x_{1}x_{d}+x_{2}x_{d-1}-2x_{\frac{d}{2}}x_{\frac{d+2}{2}}$
$d+2i$ $x_{2i}x_{d}-x_{i+\frac{d}{2}}^{2}$
$d+2i+1$ $x_{2i+1}x_{d}+x_{2i+2}x_{d-1}-2x_{i+\frac{d}{2}}x_{i+\frac{d+2}{2}}$
$d+c-2$$\left\{\begin{array}[]{lr}x_{c-2}x_{d}-x_{\frac{d+c-2}{2}}^{2},&\text{for even}c\\ x_{c-2}x_{d}+x_{c-1}x_{d-1}-2x_{\frac{d+c-3}{2}}x_{\frac{d+c-1}{2}},&\text{forodd }c\\ \end{array}\right.$
Table 3. Quadratic generators in $I_{d,c}$ for even $d$ and $3\leq c<\frac{d}{2}$.
Since $\frac{d}{2}>c$, all binomials listed in Table 3 are spanned by rank 3 quadratic forms by Lemma 4.2. For the index $d+c-1$, if $c$ is odd, the form $x_{c-1}x_{d}-x_{\frac{d+c-1}{2}}^{2}$ is already a rank $3$ quadratic generator. However, if $c$ is even, then we can use
$x_{c-1}x_{d}+x_{c+2}x_{d-3}-2x_{\frac{d+c-4}{2}}x_{\frac{d+c+2}{2}}\sim_{(d,c)}0.$
as a generator because $c<d-4$.
Case 2-2 : $d$ is even and $c=\frac{d}{2}$.
The issue arises at indices $d$ and $d+1$ because $x_{\frac{d}{2}}=x_{c}$. To resolve this, we replace these forms with the following binomials:
$x_{0}x_{d}+x_{2}x_{d-2}-2x_{\frac{d-2}{2}}x_{\frac{d+2}{2}}\sim_{(d,c)}0\text{and }x_{1}x_{d}+x_{4}x_{d-3}-2x_{\frac{d-2}{2}}x_{\frac{d+4}{2}}\sim_{(d,c)}0.$
Proof of Theorem 1.1.
Combine Corollary 3.2 and Theorem 4.5. ∎
Now, we prove that $X_{d}=\mathcal{C}_{d,2}\cup\mathbb{L}$ does not satisfy property $\operatorname{QR}(3)$ for $d\geq 5$.
Proposition 4.6.
Let $X_{d}$ be the projective scheme defined by the quadrics on $\mathcal{C}_{d,2}$ for $d\geq 5$. Then, we have the following:
  1. (1)
    $X_{d}$ does not satisfy property $\operatorname{QR}(3)$; Moreover,
  2. (2)
    $\delta(X_{d},4)-\delta(X_{d},3)=2$.
Proof.
We first prove that $\delta(X_{d},4)-\delta(X_{d},3)\geq 2$ for $d\geq 5$. Let $Q=\sum_{i=1}^{\frac{1}{2}(d^{2}-3d-2)}a_{i}Q_{i}$ where $a_{i}\in\mathbb{K}$ for $1\leq i\leq\frac{1}{2}(d^{2}-3d-2)$ and let $M_{Q}$ be the corresponding $\frac{1}{2}(d^{2}-3d-2)\times\frac{1}{2}(d^{2}-3d-2)$ symmetric matrix. We claim that there exist at least two independent linear forms in $\sqrt{I(4,M_{Q})}$.
We fix $Q_{1}=x_{0}x_{4}-x_{1}x_{3}$ and $Q_{2}=x_{0}x_{5}-x_{1}x_{4}$. Consider the $4\times 4$ submatrix of $M_{Q}$ obtained by selecting the rows and columns of $M_{Q}$ which correspond to $x_{0},x_{1},x_{3},x_{4}$, respectively:
$\begin{array}[]{c c}&\begin{array}[]{c c c c}x_{0}&x_{1}&x_{3}&x_{4}\\ \end{array}\\ \begin{array}[]{c c c c}x_{0}\\ x_{1}\\ x_{3}\\ x_{4}\end{array}&\begin{bmatrix}0&0&0&a_{1}\\ 0&0&-a_{1}&*\\ 0&-a_{1}&*&*\\ a_{1}&*&*&*\end{bmatrix}\end{array}.$
The determinant of this submatrix is $a_{1}^{4}$ which is contained in $I(4,M_{Q})$. Hence, $a_{1}\in\sqrt{I(4,M_{Q})}$. Now, consider the $4\times 4$ submatrix of $M_{Q}$ obtained by selecting the rows and columns of $M_{Q}$ which correspond to $x_{0},x_{1},x_{4},x_{5}$, respectively:
$\begin{array}[]{c c}&\begin{array}[]{c c c c}x_{0}&x_{1}&x_{4}&x_{5}\\ \end{array}\\ \begin{array}[]{c c c c}x_{0}\\ x_{1}\\ x_{4}\\ x_{5}\end{array}&\begin{bmatrix}0&0&a_{1}&a_{2}\\ 0&0&-a_{2}&*\\ a_{1}&-a_{2}&*&*\\ a_{2}&*&*&*\end{bmatrix}\end{array}.$
The determinant of this submatrix is $g\cdot a_{1}+a_{2}^{4}\in I(4,M_{Q})$ for $g\in\mathbb{K}[a_{1},\dots,a_{4}]_{3}$. Since $a_{1}\in\sqrt{I(4,M_{Q})}$, it follows that $a_{2}^{4}\in\sqrt{I(4,M_{Q})}$ $a_{2}\in\sqrt{I(4,M_{Q})}$. Therefore, $\delta(X_{5},4)-\delta(X_{5},3)\geq 2$
Now, we show that $\delta(X_{d},4)-\delta(X_{d},3)\leq 2$ by induction on $d\geq 5$. For the initial cases, we consider $d=5$ and $d=6$. When $d=5$, we have
$(I_{5,2})_{2}=\langle x_{4}^{2}-x_{3}x_{5},x_{1}x_{4}-x_{0}x_{5},x_{3}^{2}-x_{1}x_{5},x_{1}x_{3}-x_{0}x_{4}\rangle\subset\mathbb{K}[x_{0},x_{1},x_{3},x_{4},x_{5}].$
Thus, $\delta(X_{5},4)-\delta(X_{5},3)\leq 2$.
Suppose $d=6$. There are $8$ quadratic generators of $I(X_{6})$ as follows:
$Q_{1}=x_{0}x_{4}-x_{1}x_{3},Q_{2}=x_{0}x_{5}-x_{1}x_{4},Q_{3}=x_{0}x_{6}-x_{3}^{2},Q_{4}=x_{1}x_{5}-x_{3}^{2},$
$Q_{5}=x_{1}x_{6}-x_{3}x_{4},Q_{6}=x_{3}x_{5}-x_{4}^{2},Q_{7}=x_{3}x_{6}-x_{4}x_{5},Q_{8}=x_{4}x_{6}-x_{5}^{2}.$
Hence $\delta(\pi_{e_{2}}(C^{6}),4)=8$.
We claim that there are $6$ generators can be expressed as a sum of rank $3$ generators without $x_{2}$. We can check that
  1. (1)
    $Q_{3}=x_{0}x_{6}-x_{3}^{2},Q_{4}=x_{1}x_{5}-x_{3}^{2},Q_{6}=x_{3}x_{5}-x_{4}^{2},x_{8}=x_{4}x_{6}-x_{5}^{2}$ are rank $3$ itself.
  2. (2)
    $Q_{5}=x_{1}x_{6}-x_{3}x_{4}\sim_{(6,2)}0$ by Lemma 4.4 with $i=1$.
  3. (3)
    $Q_{6}=x_{3}x_{6}-x_{4}x_{5}=x_{3}x_{6}+x_{4}x_{5}-2x_{4}x_{5}\sim_{(6,2)}0$ by Lemma 4.2.
Thus, $\delta(\pi_{e_{2}}(C^{6}),4)-\delta(\pi_{e_{2}}(C^{6}),3)=2$
For the inductive hypothesis, let $d\geq 7$ and suppose that $\delta(X_{d-1},4)-\delta(X_{d-1},3)=2$. It is enough to show that the generators in $I_{2,d}\setminus I_{2,d-1}$ can be expressed as sums of rank $3$ quadratic generators.
  1. Case 1 :
    $d$ is odd
For the indices $d$ and $d+1$, the following hold:
$x_{0}x_{d}+x_{1}x_{d-1}-2x_{\frac{d-1}{2}}x_{\frac{d+1}{2}}\sim_{(d,2)}0\text{and }x_{1}x_{d}-x_{\frac{d+1}{2}}^{2}\sim_{(d,2)}0.$
For indices greater than or equal to $d+2$, all generators of $I_{2,d}$ can be obtained from the $2\times 2$ minors of the matrix
$\begin{pmatrix}x_{3}&x_{4}&\cdots&x_{d-1}\\ x_{4}&x_{5}&\cdots&x_{d}\end{pmatrix}.$
These generators can be expressed as sums of rank $3$ quadratic forms as the rational normal curves satisfy property $\operatorname{QR}(3)$.
  1. Case 2 :
    $d$ is even
For the indices $d$ and $d+1$, we have
$x_{0}x_{d}-x_{\frac{d}{2}}^{2}\sim_{(d,2)}0\text{ and }x_{1}x_{d}+x_{4}x_{d-3}-2x_{\frac{d-2}{2}}x_{\frac{d+4}{2}}\sim_{(d,2)}0.$
For indices greater than or equal to $d+2$, the generators are spanned by rank $3$ quadratic forms by the same reasoning above. ∎
5. $\operatorname{QR}(k)$ properties of $\pi_{p}(\mathcal{C}_{d})$ for rank $3$ point $p$.
In this section, we investigate whether the space of quadrics on the projected curve $\pi_{p}(\mathcal{C}_{d})$ can be spanned by rank 3 quadratic forms depending on points of rank $3$. We recall that $\pi_{p}(\mathcal{C}_{d})$ satisfies property $\operatorname{QR}(3)$ for points of rank at most two.
Suppose $p$ is a point in $\mathcal{C}_{d}^{3}\setminus\mathcal{C}_{d}^{2}$ for $d\geq 5$, and let $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})\subset\mathbb{P}^{d-1}$ be the image of $\mathcal{C}_{d}$ under the linear projection. In this case, $\mathcal{C}$ is not generated by quadrics due to the existence of a trisecant line $\mathbb{L}$ to $\mathcal{C}$, but $\mathcal{C}\cup\mathbb{L}$ is generated by quadratic polynomials. Let $X$ be the scheme defined by the quadrics in $\mathcal{C}$, i.e., $X=\mathcal{C}\cup\mathbb{L}$.
There are three cases depending on how $\mathcal{C}$ and $\mathbb{L}$ intersect, and each case can be described in terms of the multiplicity type of the lower-degree generator $F_{l}$ of the apolar ideal $(f_{p})^{\perp}=(F_{l},F_{h})$:
  1. (1)
    $F_{l}=\ell^{3}$ for a linear binary form $\ell$ $\iff$ $\mathcal{C}$ intersects $\mathbb{L}$ at a triple point.
  2. (2)
    $F_{l}=\ell_{1}^{2}\ell_{2}$ for linearly independent linear binary forms $\ell_{1},\ell_{2}$
    $\iff$ $\mathcal{C}$ intersects $\mathbb{L}$ at a double point and a simple point.
  3. (3)
    $F_{l}=\ell_{1}\ell_{2}\ell_{3}$ for linearly independent linear binary forms $\ell_{1},\ell_{2},\ell_{3}$
    $\iff$ $\mathcal{C}$ intersects $\mathbb{L}$ at three distinct simple points.
By [IK99, Lemma 1.31 and Proposition 1.36], given the degree of the lower-degree generator of $f_{p}^{\perp}$, we have a unique general additive decomposition of $f_{p}$. Moreover, each term in this general additive decomposition corresponds to a point on $C_{d}$ (with multiplicities). For example, in the case of (2), from the additive decomposition of $f_{p}$, the point $p$ is spanned by a double point and a simple point on $C_{d}$ (corresponding to $\ell_{1}^{2}$ and $\ell_{2}$, respectively). Therefore, the trisecant line $\mathbb{L}$ must intersect $\pi_{p}(C_{d})$ at the image of the double point and the simple point.
Example 5.1.
Let $e_{i}$ for $i=0,\dots,d$ be the torus-fixed points in $\mathbb{P}^{d}$ for $d\geq 5$. Let $\mathbb{L}$ be the trisecant line to $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})$.
  1. (1)
    If $p=e_{2}$, i.e., $(f_{p})^{\perp}=(S^{3},T^{d-1})$, then $\mathcal{C}\cap\mathbb{L}$ consists of a triple point.
  2. (2)
    If $p=e_{1}+e_{d}$, i.e., $(f_{p})^{\perp}=(S^{2}T,dS^{d-1}-T^{d-1})$, then $\mathcal{C}\cap\mathbb{L}$ consists of a double point and a simple point.
  3. (3)
    If $p=\sum_{i=0}^{d}(-1)^{i}e_{i}+(-1)^{d}\sum_{i=0}^{d}e_{i}+(-1)^{d}2e_{d}$, i.e., $(f_{p})^{\perp}=(S^{3}-ST^{2},T^{d-1})$, then $\mathcal{C}\cap\mathbb{L}$ consists of three distinct simple points.
We claim that the scheme $X$ does not satisfy $\operatorname{QR}(3)$ if $\mathcal{C}\cap\mathbb{L}$ contains a non-simple point (i.e., the first two cases). However, in the third case where $\mathcal{C}\cap\mathbb{L}$ consists of three simple points, the scheme $X$ possibly satisfies property $\operatorname{QR}(3)$.
5.1. A triple point.
Let $p$ be a point in $\mathcal{C}_{d}^{3}\setminus\mathcal{C}_{d}^{2}$ for $d\geq 5$ such that $\pi_{p}(\mathcal{C}_{d})\cap\mathbb{L}$ consists of a point with multiplicity $3$. This implies that $(f_{p})^{\perp}=(\ell^{3},g)\subset\mathbb{K}[S,T]$ with $\deg\ell=1$ and $\deg g=d-1$. We claim that the scheme $X:=\pi_{p}(\mathcal{C}_{d})\cup L$ for $d\geq 5$ does not satisfy property $\operatorname{QR}(3)$.
Proposition 5.2.
For $d\geq 5$, let $p$ be a point in $\mathbb{P}^{d}$ such that $(f_{p})^{\perp}=(\ell^{3},g)$ with $\deg\ell=1$ and $\deg g=d-1$.
  1. (1)
    The scheme $X$ does not satisfy property $\operatorname{QR}(3)$. Moreover,
  2. (2)
    $\delta(X,4)-\delta(X,3)\geq 2.$
Proof.
We first note that it suffices to consider the point $p\in\mathbb{P}^{d}$ corresponding to $f_{p}\in\mathbb{K}[s,t]_{d}$ such that $(f_{p})^{\perp}=(T^{3},g(S,T))\subset\mathbb{K}[S,T]$ by using an automorphism of $\mathbb{P}^{1}$ that sends the linear form $\ell$ to $T$ in $\mathbb{K}[S,T]_{1}$. Moreover, we can take any higher degree generator of $(f_{p})^{\perp}$ among any forms in $\{g(S,T)-T^{3}h(S,T)\mid h(S,T)\in\mathbb{K}[S,T]_{d-3}\}$. Therefore, we can assume
$(f_{p})^{\perp}=(T^{3},\alpha_{0}S^{d-1}+\alpha_{1}S^{d-2}T+\alpha_{2}S^{d-3}T^{2})$
for some $\alpha_{0},\alpha_{1},\alpha_{2}\in\mathbb{K}$.
We claim that $\sqrt{I(4,M_{Q})}$ contains at least one linear form. Following the proof of Corollary 3.2, $I(X)$ is generated by the quadratic generators of
$I(\mathcal{S}(1,d-2))=\left(2\times 2\text{ minors of }\begin{bmatrix}z_{0}&z_{2}&z_{3}&\cdots&z_{d-2}\\ z_{1}&z_{3}&z_{4}&\cdots&z_{d-1}\end{bmatrix}\right)$
and
$\left\{z_{1}z_{j+4}-\alpha_{0}z_{2}z_{j+2}-\alpha_{1}z_{2}z_{j+3}-\alpha_{2}z_{2}z_{j+4}\mid 0\leq j\leq d-5\right\}.$
Let $Q=\sum_{i=0}^{N}a_{i}Q_{i}$ be the general element in $\mathbb{K}[a_{0},\dots,a_{N}]$ where $Q_{i}$ are the quadratic generators of $I(X)$ for $30\leq i\leq N=\binom{d-1}{2}-$ 3. Without loss of generality, suppose that $Q_{0}=-z_{1}z_{2}+z_{0}z_{3}$ and $Q_{1}=-z_{1}z_{3}+z_{0}z_{4}$.
Let $M_{1}$ be the $4\times 4$ submatrix of $M_{Q}$ obtained by selecting the 0th, 1st, 2nd, and 3rd rows and columns of $M_{Q}$:
$M_{1}=\begin{array}[]{c c}&\begin{array}[]{c c c c}z_{0}&z_{1}&z_{2}&z_{3}\\ \end{array}\\ \begin{array}[]{c c c c}z_{0}\\ z_{1}\\ z_{2}\\ z_{3}\end{array}&\begin{bmatrix}0&0&0&a_{0}\\ 0&0&-a_{0}&*\\ 0&-a_{0}&*&*\\ a_{0}&*&*&*\end{bmatrix}\end{array}.$
where * denotes some linear forms in $\mathbb{K}[a_{0},\dots,a_{N}]$. The determinant of this submatrix is $a_{0}^{4}$. Thus $a_{0}\in\sqrt{I(4,M_{Q})}$.
Next, let $M_{2}$ be the $4\times 4$ submatrix of $M_{Q}$ obtained by selecting the 0th, 1st, 3rd, and 4nd rows and columns of $M_{Q}$:
$M_{2}=\begin{array}[]{c c}&\begin{array}[]{c c c c}z_{0}&z_{1}&z_{3}&z_{4}\\ \end{array}\\ \begin{array}[]{c c c c}z_{0}\\ z_{1}\\ z_{3}\\ z_{4}\end{array}&\begin{bmatrix}0&0&a_{0}&a_{1}\\ 0&0&-a_{1}&*\\ a_{0}&-a_{1}&*&*\\ a_{1}&*&*&*\end{bmatrix}\end{array}.$
where * are some linear forms in $\mathbb{K}[a_{0},\dots,a_{N}]$. The determinant of this submatrix is $a_{0}f+a_{1}^{4}\in I(4,M_{Q})$ where $f$ is a cubic form. Since $a_{0}\in\sqrt{I(4,M_{Q})}$, it follows that $a_{1}\in\sqrt{I(4,M_{Q})}$.
Hence, $X$ does not satisfy property $\operatorname{QR}(3)$ and $\delta(X,4)-\delta(X,3)\geq 2$. ∎
Considering the analysis of the case $p=e_{2}$ (see Proposition 4.6), we conjecture that
$\delta(X,4)-\delta(X,3)=2.$
For $d\leq 6$, we have observed that $\delta(X,4)-\delta(X,3)=2$ for randomly generated values of $\alpha_{0},\alpha_{1},\alpha_{2}$ provided by the computer algebra system [GS].
5.2. A double point and a simple point.
Let $p$ be a point in $\mathcal{C}_{d}^{3}\setminus\mathcal{C}_{d}^{2}$ for $d\geq 5$ such that $\pi_{p}(\mathcal{C}_{d})\cap\mathbb{L}$ consists of a double point and a simple point. This implies that $(f_{p})^{\perp}=(\ell_{1}^{2}\ell_{2},g)\subset\mathbb{K}[S,T]$ with $\deg\ell_{1}=\deg\ell_{2}=1$ and $\deg g=d-1$. We claim that the scheme $X=\pi_{p}(\mathcal{C}_{d})\cup\mathbb{L}$ for $d\geq 5$ does not satisfy property $\operatorname{QR}(3)$.
Proposition 5.3.
Let $p$ be a point in $\mathbb{P}^{d}$ such that $(f_{p})^{\perp}=(\ell_{1}^{2}\ell_{2},g)\subset\mathbb{K}[S,T]$ with $\deg l_{1}=\deg l_{2}=1$ and $\deg g=d-1$. The scheme $X$ does not satisfy property $\operatorname{QR}(3)$ for $d\geq 5$.
Proof.
We first remark that it suffices to consider the point $p\in\mathbb{P}^{d}$ corresponding to $f_{p}\in\mathbb{K}[s,t]_{d}$ such that $(f_{p})^{\perp}=(T^{2}(S+\alpha T),g(S,T))$ for some $\alpha\in\mathbb{K}$, by applying an automorphism on $\mathbb{P}^{1}$ that sends the linear form $\ell_{1}$ to $T$ in $\mathbb{K}[S,T]_{1}$.
Moreover, we can take any higher-degree generator of $(f_{p})^{\perp}$ among any forms in $\{g(S,T)-T^{2}(S+\alpha T)h(S,T)\mid h(S,T)\in\mathbb{K}[S,T]_{d-3}\}$. Therefore, we can assume
$(f_{p})^{\perp}=(T^{2}(S+\alpha T),\beta_{0}S^{d-1}+\beta_{1}S^{d-2}T+\beta_{2}T^{d-1})$
for some $\alpha,\beta_{0},\beta_{1},\beta_{2}\in\mathbb{K}$.
We claim that $\sqrt{I(4,M_{Q})}$ contains at least one linear form. Following the proof of Corollary 3.2, $I(\mathcal{C}_{d})_{2}$ is spanned by the generators of $I(S(1,d-2))$ and
$\left\{z_{1}z_{j+3}+\alpha z_{1}z_{j+4}-\beta_{0}z_{2}z_{j+2}-\beta_{1}z_{2}z_{j+3}-\beta_{2}z_{j+2}z_{d-1}\mid 0\leq j\leq d-5\right\}.$
Then, $I(4,M_{Q})$ contains $a^{4}$ where $a$ is the coefficient of the quadratic form $z_{0}z_{3}-z_{1}z_{2}$ in $Q$ since $M_{Q}$ contains a $4\times 4$ lower triangular submatrix whose diagonal entries are $\{a,-a,-a,a\}$. Thus, $\sqrt{I(4,M_{Q})}$ contains a linear form $a$, and so $\mathcal{C}_{d}$ fails to satisfy property $\operatorname{QR}(3)$. ∎
For $d=5$, we observe that $\delta(X,4)-\delta(X,3)=1$ for randomly generated values of $\alpha,\beta_{0},\beta_{1},\beta_{2}$ provided by the computer algebra system [GS]. In addition, for a point $p=e_{1}+e_{6}\in\mathbb{P}^{6}$ (where $f_{p}^{\perp}=(S^{2}T,6S^{5}-T^{5})$), we can check that $\delta(X,4)-\delta(X,3)=1$ by [GS]. These examples support 1.4.
5.3. Three simple points.
Suppose $X=\pi_{p}(\mathcal{C}_{d})\cup\mathbb{L}$ where the intersection $\pi_{p}(\mathcal{C}_{d})\cap\mathbb{L}$ consists of three simple points. We note that it suffices to consider the point $p\in\mathbb{P}^{d}$ such that
$(f_{p})^{\perp}=(S(S+\alpha_{1}T)(S+\alpha_{2}T),\beta_{0}S^{d-1}+\beta_{1}S^{d-2}T+\beta_{2}T^{d-1}),$
for $\alpha_{1},\alpha_{2},\beta_{0},\beta_{1},\beta_{2}\in\mathbb{K}$.
For $d=5$, we check that $X$ satisfies property $\operatorname{QR}(3)$ for randomly generated values of $\alpha_{1},\alpha_{2},\beta_{0},\beta_{1},\beta_{2}$ using the computer algebra system [GS].
We claim that $X$ satisfies property $\operatorname{QR}(3)$ if
$p=\sum_{i=0}^{d}(-1)^{i}e_{i}+(-1)^{d}\sum_{i=0}^{d}e_{i}+(-1)^{d}2e_{d}.$
We first verify this claim for low-degree cases, specifically for $d=6$ and $d=7$.
Lemma 5.4.
  1. (1)
    For a point $p\in\mathbb{P}^{6}$ such that $(f_{p})^{\perp}=(S^{3}-ST^{2},T^{5})$, $\pi_{p}(C_{6})\cup\mathbb{L}$ satisfies property $\operatorname{QR}(3)$.
  2. (2)
    For a point $p\in\mathbb{P}^{7}$ such that $(f_{p})^{\perp}=(S^{3}-ST^{2},T^{6})$, $\pi_{p}(C_{7})\cup\mathbb{L}$ satisfies property $\operatorname{QR}(3)$.
Proof of (1).
Let $LT:\mathbb{P}^{6}\to\mathbb{P}^{6}$ be a linear transformation that maps a point $p$ to $e_{0}$, and let $T^{*}$ denote the $k$-linear homomorphism induced by $LT$. Let $I=(g_{1},\dots,g_{8})$ be the saturated ideal of the rational normal curve $C_{6}$, and let $J=(h_{1},\dots,h_{8})$ where each $h_{i}$ for $i=1,\dots,8$ is the quadratic form obtained by applying $T^{*}$ to the variables in $g_{i}$.
Then,
$I(\pi_{p}(C_{6}))=J\cap\mathbb{K}[y_{1},\dots,y_{6}],$
where $\mathbb{K}[y_{0},y_{1},\dots,y_{d}]$ is the coordinate ring of the image under $LT$. The ideal of $X=\pi_{p}(\mathcal{C}_{d})\cup\mathbb{L}$ is given by $\left(J\cap\mathbb{K}[y_{1},\dots,y_{d}]\right)_{2}$. Considering the isomorphism $f:H^{0}(C_{6},\mathcal{O}_{\mathbb{P}^{1}}(6))\to R_{1}$, the elimination of a variable is equivalent to requiring that the coefficient of the term mapping to $y_{0}$ via the $Q_{A,B}$-map is zero given a decomposition of the line bundle $\mathcal{O}_{\mathbb{P}^{1}}(6)=A^{2}\otimes B$ on $\mathcal{C}_{6}$.
We fix the decomposition $\mathcal{O}_{\mathbb{P}^{1}}(6)=\mathcal{O}_{\mathbb{P}^{1}}(2)^{2}\otimes\mathcal{O}_{\mathbb{P}^{1}}(2)$ and let $l,m,\text{ and }n$ be general members of $\mathcal{O}_{\mathbb{P}^{1}}(2)\simeq\mathbb{K}[S,T]_{2}$. The vanishing condition of the coefficient can be decomposed into following three cases:
  1. (i)
    $l=S^{2}-T^{2},~{}m=\gamma_{0}ST+\gamma_{1}T^{2},~{}n=2\gamma_{0}\gamma_{1}S^{2}-(\gamma_{0}^{2}+\gamma_{1}^{2})ST$;
  2. (ii)
    $l=S^{2}-ST,~{}m=\gamma_{0}ST+\gamma_{1}T^{2},~{}n=(-2\gamma_{0}\gamma_{1}-\gamma_{1}^{2})S^{2}+\gamma_{1}^{2}ST+(\gamma_{0}+\gamma_{1})^{2}T^{2}$;
  3. (iii)
    $l=S^{2}+ST,~{}m=\gamma_{0}ST,~{}n=(-2\gamma_{0}\gamma_{1}-\gamma_{1}^{2})S^{2}-\gamma_{1}^{2}ST+(\gamma_{0}+\gamma_{1})^{2}T^{2}$
for $\gamma_{0},\gamma_{1}\in\mathbb{K}\setminus\{0\}$.
Since we obtain degree $6$ binary forms in $\gamma_{0}$ and $\gamma_{1}$ from $Q_{\mathcal{O}_{\mathbb{P}^{1}}(2),\mathcal{O}_{\mathbb{P}^{1}}(2)}$, there are $7$ generators and there are $5$ independent quadratic forms in $I(\pi_{p}(\mathcal{C}_{6}))_{2}$ among them. Therefore, we obtain $5$ independent rank $3$ quadratic generators of $I(\pi_{p}(\mathcal{C}_{6}))$ for each cases and obtained $8$ independent generators in total, as verified by the computer software system [GS].
Proof of (2). As in (1), the ideal of $X$ is given by $\left(J\cap\mathbb{K}[y_{1},\dots,y_{7}]\right)_{2}$ where $J$ is generated by the quadratic forms obtained by applying the $\mathbb{K}$-linear map $T^{*}:\mathbb{K}[x_{0},\dots,x_{7}]\to\mathbb{K}[y_{0},\dots,y_{7}]$ to the variables in the quadratic generators of the defining ideal of $C_{7}\subset\mathbb{P}^{7}$.
We fix the decomposition $\mathcal{O}_{\mathbb{P}^{1}}(7)=\mathcal{O}_{\mathbb{P}^{1}}(2)^{2}\otimes\mathcal{O}_{\mathbb{P}^{1}}(3)$. Let $l$ and $m$ be general elements of $\mathcal{O}_{\mathbb{P}^{1}}(2)\simeq\mathbb{K}[S,T]_{2}$, and $n$ be a general element of $\mathcal{O}_{\mathbb{P}^{1}}(3)\simeq\mathbb{K}[S,T]_{3}$. The cases where $l$, $m$, and $n$ satisfy the vanishing condition of the coefficient are distinguished as the following three cases:
  1. (i)
    $\begin{cases}l=S^{2}-T^{2},\;m=\gamma_{0}ST+\gamma_{1}T^{2},\\ n=\gamma_{2}(S^{3}-ST^{2})-2\gamma_{0}\gamma_{1}S^{2}T+(\gamma_{0}^{2}+\gamma_{1}^{2})ST^{2}\end{cases}$
  2. (ii)
    $\begin{cases}l=S^{2}-ST,\;m=\gamma_{1}St+\gamma_{0}T^{2},\\ n=\gamma_{2}(S^{3}-ST^{2})+((\gamma_{1}+\gamma_{0})^{2}-\gamma_{0}^{2})(S^{2}T+ST^{2})-(\gamma_{1}+\gamma_{0})^{2}(T^{3}+ST^{2})\end{cases}$
  3. (iii)
    $\begin{cases}l=S^{2}+ST,\;m=-\gamma_{1}ST+\gamma_{0}T^{2},\\ n=\gamma_{2}(S^{3}-ST^{2})+((\gamma_{1}+\gamma_{0})^{2}-\gamma_{0}^{2})(S^{2}T-ST^{2})-(\gamma_{1}+\gamma_{0})^{2}(T^{3}-ST^{2})\end{cases}$
for $\gamma_{0},\gamma_{1},\gamma_{2}\in\mathbb{K}\setminus\{0\}$.
We obtain $9$ independent quadratic forms in $I(\pi_{p}(\mathcal{C}_{7}))_{2}$ in each case and the 27 quadratic forms span the space $I(\pi_{p}(\mathcal{C}_{d}))_{2}$ which is verified using the computer algebra system [GS].
We now prove the claim that $X=\pi_{p}(\mathcal{C}_{d})\cup\mathbb{L}$ satisfies property $\operatorname{QR}(3)$ for the point $p=\sum_{i=0}^{d}(-1)^{i}e_{i}+(-1)^{d}\sum_{i=0}^{d}e_{i}-2e_{d}$.
Proposition 5.5.
Let $p$ be a point in $\operatorname{\mathbb{P}}^{d}$ (for $d\geq 6$) corresponding a binary form $f_{p}$ such that $(f_{p})^{\perp}=(S^{3}-ST^{2},T^{d-1})$. Then, the scheme $X=\pi_{p}(\mathcal{C}_{d})\cup\mathbb{L}$ satisfies $\operatorname{QR}(3)$ property.
Proof.
Let $LT:\mathbb{P}^{d}\to\mathbb{P}^{d}$ be a linear map on $\mathbb{P}^{d}$ that sends a point $p$ to $e_{0}$. Let $I$ be the ideal generated by the $2\times 2$ minors of the matrix
$\begin{bmatrix}x_{0}&x_{1}&\cdots&x_{d-1}\\ x_{1}&x_{2}&\cdots&x_{d}\end{bmatrix}$
and let $J$ be the ideal generated by the quadratic forms obtained by applying the $\mathbb{K}$-linear map $T^{*}:\mathbb{K}[x_{0},\dots,x_{d}]\to\mathbb{K}[y_{0},\dots,y_{d}]$ to the variables in the generators of $I$. Then
$I(\pi_{p}(\mathcal{C}_{d}))=J\cap\mathbb{K}[y_{1},\dots,y_{d}]$
where $J$ is $2\times 2$ minors of
$\begin{bmatrix}y_{0}&y_{1}&y_{2}+y_{0}&y_{3}+y_{1}&\cdots&y_{d-1}+\cdots+y_{1}\\ y_{1}&y_{2}+y_{0}&y_{3}+y_{1}&y_{4}+y_{2}+y_{0}&\cdots&y_{d}\end{bmatrix}.$
Therefore, we obtain the ideal of $X=\pi_{p}(\mathcal{C}_{d})\cup\mathbb{L}$ as follows:
$I(X)=$ (2×2 minors of [$(2\times 2\text{ minors of }\begin{bmatrix}y_{2}&y_{3}&\cdots&y_{d-2}&y_{d-1}+\cdots+y_{1}\\ y_{3}&y_{4}&\cdots&y_{d-1}&y_{d}\end{bmatrix})$
$+(y_{1}y_{i+2}-y_{1}y_{i}-y_{3}y_{i}:2\leq i\leq d-3)$
First, the cases $d=6,7$ are covered by Lemma 5.4.(1) and Lemma 5.4.(2).
Suppose $d\geq 8$ and $X=\mathbb{L}\cup\pi_{p}(\mathcal{C}_{d^{\prime}})$ satisfies $\operatorname{QR}(3)$ for all integers $d^{\prime}<d$. For brevity, we omit the line bundles $A$ and $B$ from the decomposition $\mathcal{O}_{\mathbb{P}^{1}}(d)=A^{\otimes 2}\otimes B$ in the $Q_{A,B}$-map and simply write $Q$ instead.
Case 1: $d$ is even.
Among the generarors of $I=I(X)$, $2\times 2$ minors of $\begin{bmatrix}y_{2}&y_{3}&\cdots&y_{d-2}\\ y_{3}&y_{4}&\cdots&y_{d-1}\end{bmatrix}$ can be written as a sum of rank $3$ quadratic forms. Therefore, we claim that
$y_{i}y_{d}-y_{i+1}(y_{d-1}+\dots+y_{1})\sim_{I}0\text{ for }2\leq i\leq d-2$
and
$y_{1}y_{i+2}-y_{1}y_{i}-y_{3}y_{i}\sim_{I}0\text{ for }2\leq i\leq d-3.$
By the induction hypothesis, we have
$z_{i}z_{d-1}-z_{i+1}(z_{d-2}+\cdots+z_{0})=\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j})$
for some $Q(l_{j},m_{j},h_{j})\in\mathbb{K}[z_{0},z_{2},z_{3},\dots,z_{d-1}]$. Therefore, we have that
$\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j}T)=y_{i+1}y_{d}-y_{i+2}(y_{d-1}+\cdots+y_{1})$
for $22\leq i\leq d-$ 2. Note that we didn’t check $y_{2}y_{d}-y_{3}(y_{d-1}+\cdots+y_{1})\sim_{I}0$ yet.
By induction hypothesis,
$w_{1}w_{i+2}-w_{i}(w_{1}+w_{3})=\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j})\in k[w_{0},w_{1},w_{3},\dots,w_{d-2}]$
for $52\leq i\leq d-$ 5. Then, we have
$\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j}S^{2})=y_{1}(y_{i+2})-(y_{3}+y_{1})y_{i}$
for $52\leq i\leq d-$ 5.
Moreover, since $\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j}T^{2})=y_{3}y_{i+4}-(y_{5}+y_{3}+y_{1})y_{i+2}$ for $52\leq i\leq d-$ 5 and $y_{3}y_{i+4}-y_{5}y_{i+2}=(y_{3}y_{i+4}-y_{4}y_{i+3})+(y_{4}y_{i+3}-y_{5}y_{i+2})\sim_{I}0$, we obtain
$y_{1}y_{i+4}-(y_{3}+y_{1})y_{i+2}\sim_{I}0$
for $52\leq i\leq d-$ 5. This implies that
$y_{1}y_{i+2}-(y_{3}+y_{1})y_{i}\sim_{I}0$
for $32\leq i\leq d-$ 3.
Now, we claim that
$y_{2}y_{d}-y_{3}(y_{d-1}+\cdots+y_{1})\sim_{I}0.$
Suppose $k$ is an integer such that $d=2k$. Note that
(5.1) $Q(S^{k}-S^{k-2}T^{2},T^{k},1)=(-y_{2}+y_{4})y_{d}-y_{k+2}^{2},$
(5.2) $y_{4}y_{d}-y_{5}(y_{d-1}+\cdots+y_{1}),$
(5.3) $y_{1}y_{5}-(y_{3}+y_{1})y_{3},\text{ and }$
(5.4) $y_{5}y_{d-1}-y_{k+2}^{2}=\sum_{i=0}^{k-4}(y_{i+5}y_{d-1-i}-y_{i+6}y_{d-2-i})$
are spanned by rank $3$ quadratic forms. Therefore,
[…]
$=y_{2}y_{d}-(y_{5}y_{d-3}+y_{5}y_{d-5}+\cdots+y_{5}y_{3})-y_{3}(y_{3}+y_{1})$
$\sim_{I}y_{2}y_{d}-y_{3}(y_{d-1}+y_{d-3}+\cdots+y_{3}+y_{1})$
since $y_{5}y_{d-1-2t}\sim_{I}y_{3}y_{d+1-2t}$ for $21\leq t\leq k-$ 2.
Case 2: $d$ is odd.
By the inductive hypothesis, we have $z_{i}z_{d-1}-z_{i+1}(z_{d-2}+\cdots+z_{1})=\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j})\in\mathbb{K}[z_{1},z_{2},\dots,z_{d-1}].$ Therefore,
$\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j}T)=y_{i+1}y_{d}-y_{i+2}(y_{d-1}+\cdots+y_{0})\in\mathbb{K}[z_{1},z_{2},\dots,z_{d-1}]$
for $22\leq i\leq d-$ 2.
On the other hand, by the inductive hypothesis, $w_{0}w_{i+2}-(w_{2}+w_{0})w_{i}=\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j})\in\mathbb{K}[w_{0},\dots,w_{d-2}]$ for $52\leq i\leq d-$ 5. Then,
$\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j}S^{2})=y_{0}y_{i+2}-(y_{2}+y_{0})y_{i}$
for $52\leq i\leq d-$ 5 and
$\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j}T^{2})=(y_{2}+y_{0})y_{i+4}-(y_{4}+y_{2}+y_{0})y_{i+2}$
for $52\leq i\leq d-$ 5. Moreover,
$y_{0}y_{i+4}-y_{4}y_{i+2}=(y_{2}y_{i+4}-y_{3}y_{i+3})+(y_{3}y_{i+3}-y_{4}y_{i+2})\sim_{I}0$
for $52\leq i\leq d-$ 5. Therefore,
$y_{0}y_{i+2}-(y_{2}+y_{0})y_{i}\sim_{I}0$
for $32\leq i\leq d-$ 3.
Now, we claim that
(5.5) $y_{2}y_{d}-y_{3}(y_{d-1}+\dots+y_{0})\sim_{I}0.$
Let $k$ be the integer such that $d=2k+1$. We have
(5.6) $(-y_{2}+y_{4})y_{d}+(-y_{3}+y_{5})(y_{d-1}+\cdots+y_{0})-2(y_{k+2})(y_{k+3})\sim_{I}0$
since (5.6) equals to $Q(S^{k}-S^{k-2}T^{2},\,T^{k},\,S+T)-Q(S^{k}-S^{k-2}T^{2},\,T^{k},\,S)-Q(S^{k}-S^{k-2}T^{2},\,T^{k},\,T)$. Moreover,
(5.7) $y_{4}y_{d}-y_{5}(y_{d-1}+\cdots+y_{0})\sim_{I}0\text{ and }$
(5.8) $y_{0}y_{5}-(y_{2}+y_{0})y_{3}\sim_{I}0.$
Then, we have
[…]
=\;= $-2y_{3}(y_{d-1}+\cdots+y_{4})+2y_{5}(y_{d-1}+\cdots+y_{2})-2y_{k+2}y_{k+3}$
=\;= $-2\left((y_{3}y_{d-1}-y_{5}y_{d-3})+(y_{3}y_{d-3}-y_{5}y_{d-5})+\cdots+(y_{3}y_{4}-y_{5}y_{2})\right)$
$+2(y_{5}y_{d-1}-y_{k+2}y_{k+3})$
[…]