### v2

# On the rank index of projective curves of almost minimal degree

Abstract.
[…]
Key words and phrases:
apolarity, linear projections, low rank quadrics, projective curves with parametrization, property $\operatorname{QR}(k)$
2020 Mathematics Subject Classification:
14A25, 14H45, 14N05, 15A63, 16E45
Corresponding author: Jaewoo Jung (jaewoojung@dgist.ac.kr)
1. Introduction
Throughout this paper we work over an algebraically closed field $\mathbb{K}$ with $\operatorname{char}(\mathbb{K})=0$. According to [8], a non-degenerate projective variety $X\subset\mathbb{P}^{r}$ is said to satisfy property $\operatorname{QR}(k)$ if its homogeneous ideal is generated by quadratic polynomials of rank at most $k$. When $X$ is cut out ideal-theoretically by quadrics, we define the rank index of $X$, denoted by $\operatorname{rank-index}(X)$, as the least integer $k$ such that $X$ satisfies property $\operatorname{QR}(k)$. Thus
$\operatorname{rank-index}(X)\geq 3$
since $X$ is irreducible and nondegenerate in $\operatorname{\mathbb{P}}^{r}$. Many classical constructions in projective geometry such as rational normal scrolls and Segre-Veronese varieties satisfy property $\operatorname{QR}(4)$ since their homogeneous ideals are generated by $2$-minors of one or several matrices of linear forms (cf. [1], [7], [17], [18], [19], etc). Recently, it has been found that there are many projective varieties whose rank index is equal to $3$, which is the minimal value for irreducible varieties. The $d$-th Veronese embedding of $\operatorname{\mathbb{P}}^{n}$ satisfies property $\operatorname{QR}(3)$ for all $n\geq 1$ and $d\geq 2$ if $\operatorname{char}(\mathbb{K})\neq 3$ (cf. [8]). Every linearly normal curve of arithmetic genus $g$ and degree $\geq 4g+4$ satisfies property $\operatorname{QR}(3)$ (cf. [16]). Also in [13] the authors found the rank index of various projective varieties, including rational normal scrolls and del Pezzo varieties. For example, if $X$ is a smooth rational normal scroll of dimension $\geq 2$, then $\operatorname{rank-index}(X)=4$.
The main purpose of this paper is to investigate the rank index of projective curves $\mathcal{C}\subset\operatorname{\mathbb{P}}^{r}$ of degree $r+1$. To put things in perspective, we recall some of the previously known results regarding this problem. To give precise statements, we require some notation and definitions. Let $\mathcal{C}\subset\operatorname{\mathbb{P}}^{r}$ be a nondegenerate projective curve of degree $d$. Then $d\geq r$ and equality is attained if and only if $\mathcal{C}$ is a rational normal curve, which is obviously very well-understood. In particular, it holds that $\operatorname{rank-index}(\mathcal{C})=3$ (cf. [8, Corollary 2.4]). If $d=r+1$, then the following two cases may occur:
  1. Case $1.$ $\mathcal{C}$ is a linearly normal curve of arithmetic genus one;
  2. Case $2.$ $\mathcal{C}$ is an isomorphic projection of a rational normal curve $\widetilde{\mathcal{C}}\subset\operatorname{\mathbb{P}}^{r+1}$.
In the former case, it is already known that $\operatorname{rank-index}(\mathcal{C})=3$ (cf. [16, Theorem 1.1]). In the latter case, let
$\tilde{\mathcal{C}}=\mathcal{C}_{r+1}=\left\{[s^{r+1}:s^{r}t:\cdots:t^{r+1}]~|~[s:t]\in\mathbb{P}^{1}\right\}\subset\mathbb{P}^{r+1},d\geq 4$
be the standard rational normal curve of degree $r+1$, and let $\tilde{\mathcal{C}}^{k}$ denote the $k$-th self-join of $\tilde{\mathcal{C}}$. That is, $\tilde{\mathcal{C}}^{k}$ is the closure of the set of points lying in (k1)(k-1)-dimensional linear subspaces spanned by general collections of $k$ points in $\tilde{\mathcal{C}}$. For joins of projective varieties, we refer to [20]. Then our rational curve $\mathcal{C}\subseteq\mathbb{P}^{r}$ of degree $d=r+1$ may be written in the form:
[…]
for some $p\in\operatorname{\mathbb{P}}^{r+1}\setminus\tilde{\mathcal{C}}^{2}$ where $\pi_{p}:\operatorname{\mathbb{P}}^{r+1}\dashrightarrow\operatorname{\mathbb{P}}^{r}$ is the linear projection from $p$. For a point $p\in\operatorname{\mathbb{P}}^{r+1}$, the integer
$\mbox{rank}_{\tilde{\mathcal{C}}}(p)=\min\{k\mid p\in\tilde{\mathcal{C}}^{k}\},$
[…]
Now, let us state our first main result in this paper.
Theorem 1.1.
For the standard rational normal curve $\widetilde{\mathcal{C}}\subset\operatorname{\mathbb{P}}^{r+1}$ and a point $p\in\operatorname{\mathbb{P}}^{r+1}\setminus\widetilde{\mathcal{C}}^{2}$ with $\mbox{rank}_{\tilde{\mathcal{C}}}(p)\geq 4$, let
$\mathcal{C}=\pi_{p}(\tilde{\mathcal{C}})\subseteq\operatorname{\mathbb{P}}^{r}$
be a smooth rational curve of degree $r+1$. Then
  1. (1)
    $\operatorname{rank-index}(\mathcal{C})\leq 4$.
  2. (2)
    Suppose that $p=p_{c}$ is the c-th coordinate point of $\mathbb{P}^{r+1}$ (and hence $3\leq c\leq r-2$). Then $\operatorname{rank-index}(\mathcal{C})=3$.
The proofs of statements (1) and (2) of Theorem˜1.1 are given respectively in Section˜3 and Section˜4.
One of the main reasons why determining the rank index of the curve $\mathcal{C}$ in Theorem˜1.1 is interesting is that, in contrast to all previously studied varieties satisfying property $\operatorname{QR}(3)$, which are linearly normal, the curve $\mathcal{C}$ is not linearly normal. As a consequence, its geometric and algebraic properties depend continuously on the choice of the projection center $p$.
In the proof of Theorem˜1.1(1), we use the fact that there is a unique rational normal surface scroll $\mathcal{S}$ that contains $\mathcal{C}$. Indeed, the homogeneous ideal of $\mathcal{C}$ is generated by that of $\mathcal{S}$ and the quadratic polynomials obtained from the equation of $\mathcal{C}$ as a divisor of $\mathcal{S}$. Note that the existence of such a surface $\mathcal{S}$ can be found in [4, Proposition 1.2] and [3, Theorem 7.3] and the uniqueness of $\mathcal{S}$ is recently proven in [11, Theorem 5.2]. We provide a new constructive proof of the existence of $\mathcal{S}$ by using the apolar ideal of the binary form corresponding to the projection center $p$ in Section˜3.
The proof of Theorem˜1.1(2) is obtained by combining induction on the degree of $\mathcal{C}$ with the investigation of rank $3$ quadratic equations of $\tilde{\mathcal{C}}$ that are also contained in the homogeneous ideal of $\mathcal{C}$.
[…]
Conjecture 1.2.
Let $\mathcal{C}\subset\operatorname{\mathbb{P}}^{r}$ be as in Theorem˜1.1. Then $\operatorname{rank-index}(\mathcal{C})=3$.
Next, we consider the case where $\mbox{rank}_{\tilde{\mathcal{C}}}(p)=3$ and so the homogeneous ideal of $\mathcal{C}$ fails to be generated by quadratic polynomials. Then, $\mathcal{C}$ admits a unique trisecant line, say $\mathbb{L}$. Moreover, the homogeneous ideal of the reducible curve
$X:=\mathcal{C}\cup\mathbb{L}$
is generated by the quadratic polynomials in the homogeneous ideal of $\mathcal{C}$. (See Lemma˜5.1 for details.) Note that $\mathcal{C}$ is $3$-regular (cf. [9, Theorem 3]). Regarding the rank index of $X$, we obtain the following interesting partial result.
Theorem 1.3.
Suppose that $r\geq 4$. Assume that $\operatorname{rank}_{\widetilde{\mathcal{C}}}(p)=3$ and let $\mathcal{C}$, $\mathbb{L}$, and $X=\mathcal{C}\cup\mathbb{L}$ be as above. Then
  1. (1)
    $\operatorname{rank-index}(X)\leq 4$.
  2. (2)
    If $\mathcal{C}\cap\mathbb{L}$ has at most two points, or equivalently, if $\mathcal{C}\cap\mathbb{L}$ contains a point with multiplicity at least two, then $\operatorname{rank-index}(X)=4$.
The proof of Theorem˜1.3 is given in Section˜5.
By Theorem˜1.3, it remains to determine the rank index of $X$ when $\mathcal{C}\cap\mathbb{L}$ consists of three simple points. There are many example such that $X$ has property $\operatorname{QR}(3)$ and hence satisfies $\operatorname{rank-index}(X)=3$. This leads us to the following conjecture.
Conjecture 1.4.
Let $\mathcal{C}$, $\mathbb{L}$, and $X=\mathcal{C}\cup\mathbb{L}$ be as in Theorem˜1.3. Then $\operatorname{rank-index}(X)=3$ if and only if $\mathcal{C}\cap\mathbb{L}$ consists of three distinct reduced points.
Obviously, Theorem˜1.3 proves one direction of ˜1.4.
Organization of the paper.
Section˜2 outlines key definitions and foundational concepts including (smooth) curves of almost minimal degree, apolarity of binary forms, and the rank index of varieties. Section˜3 presents our primary results on the rank index for projected curves. In Section˜4, we focus on the quadratic rank index of projected curves obtained by linear projection from a standard basis in projective spaces. Section˜5 includes explicit calculations for certain projective schemes not covered in the previous sections that provide additional support for conjectures concerning the rank index.
2. Preliminaries
2.1. Apolarity of binary forms
In this subsection we provide some basics related to the notion of apolarity of binary forms, which will play an essential role later. To do so, we fix some notation which will be kept throughout: Let $R:=\mathbb{K}[s,t]$ and $\mathcal{R}:=\mathbb{K}[S,T]$ be two bivariate standard graded polynomial rings over the field $\mathbb{K}$, so that $\deg(s)=\deg(t)=\deg(S)=\deg(T)=1$, $R=\bigoplus_{i\in\mathbb{Z}}R_{i},\mathcal{R}=\bigoplus_{i\in\mathbb{Z}}\mathcal{R}_{i}$ where $R_{i}:=\bigoplus_{j=0}^{i}\mathbb{K}s^{j}t^{i-j}$ and $\mathcal{R}_{i}:=\bigoplus_{j=0}^{i}\mathbb{K}S^{j}T^{i-j}$ are the $i$-th graded components of $R$ respectively of $\mathcal{R}$ for all $i\in\mathbb{Z}$. As usual, we call the polynomials $f\in R_{i}$ and $F\in\mathcal{R}_{i}$ binary forms of degree $i$ in $R$ respectively in $\mathcal{R}$. Correspondingly, we respectively call the sets $R^{\mathrm{hom}}:=\bigcup_{i\in\mathbb{Z}}R_{i}$ and $\mathcal{R}^{\mathrm{hom}}:=\bigcup_{i\in\mathbb{Z}}\mathcal{R}_{i}$ of homogeneous polynomials in $R$ and in $\mathcal{R}$ the sets of binary forms (in $R$ and in $\mathcal{R}$).
Definition 2.1.
For all $i,j\in\mathbb{Z}$ we consider the bilinear map of $\mathbb{K}$-vector spaces
$\circ_{(j,i)}:\mathcal{R}_{j}\times R_{i}\to R_{i-j}\quad\text{and the induced map}\quad\circ:\mathcal{R}^{\mathrm{hom}}\times R^{\mathrm{hom}}\to R^{\mathrm{hom}}$
where $\circ_{(j,i)}$ is defined by
$(F,f)\mapsto F\circ_{(j,i)}f=F\circ f:=\sum_{k=0}^{j}c_{k}\frac{\partial^{j}}{\partial s^{j-k}\partial t^{k}}f$
where $F=\sum_{k=0}^{j}c_{k}S^{j-k}T^{k}$. The apolar ideal of a homogeneous polynomial $f\in R$ is defined to be the ideal generated by all homogeneous polynomials $F\in\mathcal{R}$ for which $F\circ f=0$, and we denote it by $(f)^{\perp}$. In other words,
$(f)^{\perp}:=(F\in\mathcal{R}:F\circ f=0).$
Proposition 2.2.
(cf. [10, Theorem 1.44.]) Suppose $f\in\mathbb{K}[s,t]$ is a binary form of degree $d$. Then, $(f)^{\perp}$ is a complete intersection ideal generated by two binary forms in $\mathcal{R}$. More precisely: If $F_{\ell},F_{h}$ are of minimal respectively maximal degree among all $F\in\mathcal{R}_{\mathrm{hom}}\setminus\{0\}$ with $F\circ f=0$, then
$(f)^{\perp}=(F_{l},F_{h}),$
where $V(F_{l},F_{h})=\emptyset$ and $\deg F_{l}+\deg F_{h}=d+2$ with $\deg F_{l}\leq\deg F_{h}$. If $\deg F_{l}<\deg F_{h}$, then $F_{\ell}$ is essentially uniquely determined by $f$, while $F_{h}$ is essentially unique modulo the principal ideal $\langle F_{l}\rangle$.
For example, if $f=\alpha s+\beta t\in\mathbb{K}[s,t]$ is a linear binary form, then the lower-degree generator of $(f)^{\perp}$ is $F_{l}=\beta S-\alpha T\in\mathcal{R}$ and $F_{h}$ is any quadratic form in $\mathcal{R}$.
The apolarity of forms is a useful tool for studying the decomposition of forms into sums of powers of linear forms. This decomposition is known as the Waring decomposition or the (generalized) additive decomposition of forms. For more details on the generalized additive decomposition and apolarity, see [10].
We remark that for any integer $d\geq 0$, the pairing $\circ:\mathcal{R}_{d}\times R_{d}\to R_{0}=\mathbb{K}$ is perfect and hence forms a non-degenerate inner product of the two vector spaces $\mathcal{R}_{d}$ and $R_{d}$, meaning that for any basis $v_{0},v_{1},\ldots,v_{d}$ of $\mathcal{R}_{d}$ and any basis $w_{0},w_{1},\ldots,w_{d}$ of $R_{d}$ the matrix $v_{\bullet}\circ w_{\bullet}:=(v_{i}\circ w_{j}\mid 0\leq i,j\leq d)\in\mathbb{K}^{(d+1)\times(d+1)}$ is of rank $d+1$. Therefore, we can identify $\mathcal{R}_{d}$ with the dual vector space of $R_{d}$ via the differential operator. In particular, for $f\in R_{d}$ and $g\in\mathcal{R}_{d}$, the inner product $g\circ f$ is given by
$g\circ f=\sum_{i=0}^{d}c_{i}\left(\frac{\partial}{\partial s}\right)^{d-i}\left(\frac{\partial}{\partial t}\right)^{i}f(s,t),$
where $g=g(S,T)=\sum_{i=0}^{d}c_{i}S^{d-i}T^{i}$.
Let $\{S^{d},S^{d-1}T,\cdots,T^{d}\}$ be the monomial basis of $\mathcal{R}_{d}$. Consider the dual basis of this monomial basis for $R_{d}$ with above inner product, which is given by $\{s^{[d]},\cdots,s^{[d-i]}t^{[i]},\cdots,t^{[d]}\}:=\{\frac{1}{d!}s^{d},\cdots,\frac{1}{(d-i)!i!}s^{d-i}t^{i},\cdots,\frac{1}{d!}t^{d}\}$. Indeed,
(2.1) $S^{d-i}T^{i}\circ s^{[d-j]}t^{[j]}=\delta_{ij}$
where $\delta_{ij}$ is the Kronecker delta. Therefore, any form $f(s,t)\in\mathbb{K}[s,t]_{d}$ can be expressed as
(2.2) $f(s,t)=\sum_{i=0}^{d}\left(S^{d-i}T^{i}\circ f(s,t)\right)s^{[d-i]}t^{[i]}.$
We have the following identity:
(2.3) $(\alpha s+\beta t)^{d}=\sum_{i=0}^{d}\binom{d}{i}\alpha^{d-i}\beta^{i}s^{d-i}t^{i}=d!\sum_{i=0}^{d}\alpha^{d-i}\beta^{i}s^{[d-i]}t^{[i]}$
Lemma 2.3.
Let $d$ be a positive integer. For binary forms $F(S,T)\in\mathbb{K}[S,T]_{d}$ and $(\alpha s+\beta t)^{d}\in\mathbb{K}[s,t]_{d}$,
$F(S,T)\circ(\alpha s+\beta t)^{d}=d!F(\alpha,\beta)$
Proof.
Let $F(S,T)=\sum_{j=0}^{d}c_{j}S^{d-j}T^{j}\in\mathbb{K}[S,T]_{d}$. By (2.1) and (2.3),
$F(S,T)\circ(\alpha s+\beta t)^{d}$ $=d!\sum_{i=0}^{d}\sum_{j=0}^{d}\alpha^{d-i}\beta^{i}c_{j}\left(S^{d-i}T^{i}\circ s^{[d-j]}t^{[j]}\right)$
$=d!\sum_{i=0}^{d}c_{i}\alpha^{d-i}\beta^{i}=d!F(\alpha,\beta)$
Thus, the inner product between $F(S,T)$ and $(\alpha s+\beta t)^{d}$ can be interpreted as the evaluation of $F(S,T)$ at $(\alpha,\beta)$.
2.2. Projected curves
We now discuss a correspondence between binary forms of degree $d$ and points in a projective space $\mathbb{P}^{d}$. Let $\nu_{d}:\mathbb{P}^{1}\to\mathbb{P}^{d}$ be the $d$-uple embedding (or the $d$th Veronese map). Let $\mathcal{C}_{d}:=\nu_{d}(\mathbb{P}^{1})\subseteq\mathbb{P}^{d}$ denote the image of the $d$th Veronese map, which is known as the standard rational normal curve.
Given a point $p\in\mathbb{P}^{d}$, consider the vector space of linear forms on $\mathbb{P}^{d}$ that vanish at $p$. Pulling this space back to $\mathbb{P}^{1}$ via $\nu_{d}$ yields a $d$-dimensional vector space of degree $d$ binary forms, which corresponds to a hyperplane in the space of binary forms of degree $d$. We denote by $f_{p}$ the binary form of degree $d$ (unique up to scalar multiple) that is orthogonal to this hyperplane with respect to the inner product. For $p=[p_{0}:p_{1}:\ldots:p_{d}]$, we can write $f_{p}=\sum_{i=0}^{d}p_{i}\binom{d}{i}s^{d-i}t^{i}$, unique up to scalar multiple. In particular, one can see that given a binary form $f_{p}$ in the vector space $\mathbb{K}[s,t]_{d}$, the orthogonal complement of $f_{p}$ with respect to the inner product defined by the differential operator is $((f_{p})^{\perp})_{d}$.
An alternative way to associate a point $p$ with the binary form $f_{p}$ is as follows. Under the $d$-uple embedding $\nu_{d}$, a point $\nu_{d}([\alpha:\beta])$ on the rational normal curve corresponds to the $d$th power $(\alpha s+\beta t)^{d}\in\mathbb{K}[s,t]_{d}$ (using the identification between points of $\mathbb{P}^{1}$ and binary linear forms). Since distinct points on the rational normal curve are in linearly general position, this extends additively to give a correspondence between all points in $\mathbb{P}^{d}$ and binary forms of degree $d$.
Explicitly, for $p\in\mathbb{P}^{d}$, we may express $p$ as a linear combination of $r\leq d+1$ points $p_{i}$ on $\mathcal{C}_{d}$, say $p=\sum_{i=1}^{r}c_{i}p_{i}$. Setting $p_{i}=\nu_{d}([\alpha_{i}:\beta_{i}])$, we have
$f_{p}=\sum_{i=1}^{r}c_{i}(\alpha_{i}s+\beta_{i}t)^{d}\in\mathbb{K}[s,t]_{d}.$
In this correspondence, the torus-fixed point $e_{i}:=[\delta_{i0}:\delta_{i1}:\cdots:\delta_{id}]\in\mathbb{P}^{d}$ corresponds to the monomial $s^{[d-i]}t^{[i]}\in\mathbb{K}[s,t]_{d}$ for $i=0,1,\dots,d$.
Example 2.4.
Let $\mathcal{C}_{6}$ be the standard rational normal curve of degree $6$ in $\mathbb{P}^{6}$, and let $p=e_{2}+e_{4}$ be a point where $e_{2}$ and $e_{4}$ are the second and fourth torus-fixed point in $\mathbb{P}^{6}$, respectively. Then
$(f_{p})^{\perp}=(S^{3}T-ST^{3},S^{4}-S^{2}T^{2}+T^{4})\subset\mathbb{K}[S,T].$
Note that $(f_{p})^{\perp}$ determines a unique degree $6$ form $f_{p}$, up to multiplication by a nonzero scalar in $\mathbb{K}$.
$f_{p}=s^{[4]}t^{[2]}+s^{[2]}t^{[4]}=\frac{1}{4!2!}s^{4}t^{2}+\frac{1}{2!4!}s^{2}t^{4}=\frac{1}{6!}\cdot\left(-t^{6}-s^{6}+\frac{1}{2}(s+t)^{6}+\frac{1}{2}(s-t)^{6}\right).$
The powers of linear forms $\{t^{6},s^{6},(s+t)^{6},(s-t)^{6}\}$ correspond to the four points $\{\nu_{6}([0:1]),\nu_{6}([1:0]),\nu_{6}([1:1]),\nu_{6}([1:-1])\}$ on the standard rational normal curve.
Let $\pi_{p}:\mathbb{P}^{d}\dashrightarrow\mathbb{P}^{d-1}$ be the linear projection map away from a point $p\in\mathbb{P}^{d}$. The projected curve $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})$ can be classified in terms of the $\mathcal{C}_{d}$-rank of the point $p$. For example,
  • If $p$ lies on $\mathcal{C}_{d}$, then $\mathcal{C}$ is a curve of minimal degree.
  • If $p$ lies in $\mathcal{C}_{d}^{2}\setminus\mathcal{C}_{d}$, then $\mathcal{C}$ is an arithmetically Cohen-Macaulay(aCM) curve of almost minimal degree.
  • If $p$ lies in $\mathbb{P}^{d}\setminus\mathcal{C}_{d}^{2}$, then $\mathcal{C}$ is a smooth non-aCM curve of almost minimal degree.
In fact, the graded Betti numbers of $\mathcal{C}$ are determined by the $\mathcal{C}_{d}$-rank of $p$ for $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}^{2}$, see [14].
Through the identification between points in $\mathbb{P}^{d}$ and binary forms of degree $d$, we can regard the projection map $\pi_{p}$ as a $\mathbb{K}$-linear map $\pi_{p}:\mathbb{K}[s,t]_{d}\rightarrow((f_{p})^{\perp})_{d}$ on vector spaces. Namely, for $f_{p}\in\mathbb{K}[s,t]_{d}$ associated with the point $p\in\mathbb{P}^{d}$, $((f_{p})^{\perp})_{d}$ forms a hyperplane orthogonal to $f_{p}$ with respect to the inner product $\circ$. Indeed, given $p$, let $\{p_{1},p_{2},\dots,p_{d}\}$ be a basis for $p^{\perp}=\{p^{\prime}\in\mathbb{P}^{d}:p\circ p^{\prime}=0\}$ for any inner product $\circ$ on the space $\mathbb{P}^{d}$. Since the linear projection $\pi_{p}:\mathbb{P}^{d}\dashrightarrow\mathbb{P}^{d-1}$ is the map sending $q=\alpha_{0}p+\sum_{i=1}^{d}\alpha_{i}p_{i}$ to $\sum_{i=1}^{d}\alpha_{i}p_{i}$, the linear projection $\pi_{p}:\mathbb{K}[s,t]_{d}\rightarrow\mathbb{K}[s,t]_{d}$ sends the binary form $f_{q}=\alpha_{0}f_{p}+\sum_{i=1}^{d}\alpha_{i}f_{i}\in\mathbb{K}[s,t]_{d}$ to $\sum_{i=1}^{d}\alpha_{i}f_{i}$, where $\{f_{1},f_{2},\dots,f_{d}\}$ is the basis for $((f_{p})^{\perp})_{d}$.
Furthermore, by associating a point $p\in\mathbb{P}^{d}$ with the binary form $f_{p}\in\mathbb{K}[s,t]_{d}$ and applying apolarity to this form, we can obtain deeper insights into the projected curve $\mathcal{C}$. In detail, for a point $p\in\mathbb{P}^{d}$, the apolar ideal is given by $(f_{p})^{\perp}=(g_{1},g_{2})\subset\mathbb{K}[S,T]$ with $\deg g_{1}\leq\deg g_{2}$. The $\mathcal{C}_{d}$-rank of the point $p$ equals $\deg(g_{1})$ which also specifies the type of rational normal surface containing the curve $\pi_{p}(\mathcal{C}_{d})$.
2.3. Quadratic rank index of projective schemes
For a nondegenerate projective scheme $X\subset\operatorname{\mathbb{P}}^{r}$ defined by quadratic polynomials, we denote by $I(X)$ the homogeneous vanishing ideal of $X$ in the standard graded polynomial algebra $R=\mathbb{K}[x_{0},x_{1},\ldots,x_{d}]$ and we let $\{Q_{0},Q_{1},\ldots,Q_{t<\binom{d+2}{2}}\}$ be a basis of the $\mathbb{K}$-vector space $I(X)_{2}$. We say $X$ satisfies property $\operatorname{QR}(k)$ for an integer $k$, if $I(X)$ is generated by a collection $\{Q_{0},Q_{1},\ldots,Q_{t}\}$ of quadratic forms of rank at most $k$. We define the quadratic rank-index of a variety $X$ (defined by quadrics) as the minimal number $k$ such that $X$ satisfies property $\operatorname{QR}(k)$. In this section we discuss how to compute the rank index of $X$ given the basis $\{Q_{0},Q_{1},\ldots,Q_{t}\}$.
For each field $\mathbb{F}$ (of characteristic $\neq 2$) and each quadratic polynomial $F\in\mathbb{F}[z_{0},z_{1},\ldots,z_{r}]_{2}$ let $M_{F}\in F^{(r+1)\times(r+1)}$ denote the symmetric matrix associated to $F$, so that $(z_{0},z_{1},\ldots,z_{r})M_{F}(z_{0},z_{1},\ldots,z_{r})^{T}=F$. Each member of $I(X)_{2}$ is of the form
[…]
and satisfies $M_{Q(\underline{a})}=\sum_{i=0}^{t}a_{i}M_{Q_{i}}\in K^{(r+1)\times(r+1)}$. Now, let $A_{0},A_{1},\ldots,A_{t}$ be indeterminates, set $\mathbb{F}:=\mathbb{K}(A_{0},A_{1},\ldots,A_{t})$ and consider the polynomial
$Q:=A_{0}Q_{0}+A_{1}Q_{1}+\cdots+A_{t}Q_{t}\in\sum_{i=0}^{t}A_{i}I(X)_{2}\subseteq\mathbb{F}[z_{0},z_{1},\ldots,z_{r}]_{2}$
which satisfies $M_{Q}=\sum_{i=0}^{t}A_{i}M_{Q_{i}}$. In particular, we have
$M_{Q}\in\mathbb{K}[A_{0},A_{1},\ldots,A_{t}]^{(r+1)\times(r+1)}.$
Moreover, for $1\leq k\leq r$, the form $Q(\underline{a})$ (and hence the matrix $M_{Q(\underline{a})}$) is of rank $\leq r$ if and only if $\underline{a}\in\mathbb{K}^{t+1}$ belongs to the zero-set of the ideal $I(k+1,M_{Q})\subseteq\mathbb{K}[A_{0},A_{1},\ldots,A_{t}]$ generated by all $(k+1)\times(k+1)$ minors of the matrix $M_{Q}$. Now, we identify $\mathbb{P}(I(X)_{2})$ with $\mathbb{P}^{t}$ by sending $Q(\underline{a})]=\mathbb{K}^{*}Q(\underline{a})$ to $a_{0}:a_{1}:\cdots:a_{t}]$, and for $1\leq k\leq r$ we define the locus
$\Phi_{k}(X):=\{[a_{0}:a_{1}:\cdots:a_{t}]\mid\operatorname{rank}(Q(a))\leq k\}\subseteq\mathbb{P}(I(X)_{2})=\mathbb{P}^{t}$
corresponding to all quadratic polynomials of rank at most $k$ in $I(X)$. By the above observation $\Phi_{k}(X)$ is a projective algebraic set in $\mathbb{P}^{t}$. Now, we get a descending filtration of $\mathbb{P}^{t}$ associated to $X$:
$\emptyset=\Phi_{1}(X)=\Phi_{2}(X)\subset\Phi_{3}(X)\subset\Phi_{4}(X)\subset\cdots\subset\Phi_{r}(X)\subset\operatorname{\mathbb{P}}^{t}.$
Note that any change of coordinates $T:\operatorname{\mathbb{P}}^{r}\rightarrow\operatorname{\mathbb{P}}^{r}$ induces a projective equivalence between $\Phi_{k}(X)$ and $\Phi_{k}(T(X))$.
Proposition 2.5.
[13, Proposition 2.1.] Keep the previous notations. Then
$\operatorname{rank-index}(X)=\min\{~s\mid\sqrt{I(s+1,M_{Q})}\quad\mbox{contains no non-zero linear form}\}.$
Definition and Remark 2.6.
For each $k\geq 3$, we define $\delta(X,k)$ as the maximum number of linearly independent quadrics of rank $\leq k$ in $I(X)_{2}$. In particular, $\delta(X,k)$ is one more than the dimension of the linear subspace of $\mathbb{P}^{t}$ spanned by $\Phi_{k}(X)$. Therefore, it holds that
$\delta(X,k)$
$=(t+1)-\left(\text{the maximal}~\sharp~\text{of linearly independent linear forms in}~\sqrt{I(k+1,M_{Q})}\right).$
This invariant is introduced in [11, Definition 6.2] and discussed in the proof of [11, Proposition 6.3].
2.4. Rank $3$ generators
We review the $Q$-map on a given decomposition of the vector bundle over varieties, which produces rank $3$ quadratic generators, introduced in [8]. Let $X$ be as above, let $L=\mathcal{O}_{X}(1)$ and suppose that $X\subset\operatorname{\mathbb{P}}^{r}$ is the linearly normal embedding of $X$ by $L$. Thus there is a natural isomorphism
$\varphi:H^{0}(X,L)\rightarrow R_{1}$
of $\mathbb{K}$-vector spaces. Now, assume that $L$ is decomposed as $L=A^{2}\otimes B$ for some line bundles $A$ and $B$ on $X$ such that $h^{0}(X,A)\geq 2$ and $h^{0}(X,B)\geq 1$. Define the map
$Q_{A,B}:H^{0}(X,A)\times H^{0}(X,A)\times H^{0}(X,B)\rightarrow I(X)_{2}$
by
$Q_{A,B}(s,t,h)=\varphi(s\otimes s\otimes h)\varphi(t\otimes t\otimes h)-\varphi(s\otimes t\otimes h)^{2}.$
This map is well-defined and any rank $3$ quadratic generator of $I(X)$ lies in the image of the $Q_{A,B}$-map.
3. Generators of projected curves
Let $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})$ for $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}$ be the projected curve of degree $d$ in $\mathbb{P}^{d-1}$. In this section, we demonstrate the existence of a parameterized curve that is projectively equivalent to $\mathcal{C}$, and we show that $\mathcal{C}$ satisfies property $\operatorname{QR}(4)$. We note that, since the projected curve is not linearly normal, determining whether this curve satisfies property $\operatorname{QR}(k)$ is not straightforward.
Theorem 3.1.
Let $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}$ and let $f_{p}\in\mathbb{K}[s,t]$ be the binary form corresponding to $p$. Let $g_{1},g_{2}\in\mathbb{K}[S,T]$ be binary forms that generate the ideal $(f_{p})^{\perp}$, and let $d_{i}=\deg g_{i}$ for $i=1,2$ with $2\leq d_{1}\leq d_{2}$ and $d_{1}+d_{2}=d+2$. Then, the projected curve $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})\subset\mathbb{P}^{d-1}$ is projectively equivalent to a curve with the following parametrization:
$\{[g_{2}(\alpha,\beta)\alpha^{d_{1}-2}:\dots:g_{2}(\alpha,\beta)\beta^{d_{1}-2}:g_{1}(\alpha,\beta)\alpha^{d_{2}-2}:\dots:g_{1}(\alpha,\beta)\beta^{d_{2}-2}]\mid[\alpha:\beta]\in\mathbb{P}^{1}\}.$
Proof.
Observe that the vector space $V:=\{g\in\mathbb{K}[S,T]_{d}\mid g\circ f_{p}=0\}\subseteq\mathbb{K}[S,T]_{d}$ is given by
$V$ $=g_{1}\mathbb{K}[S,T]_{d-d_{1}}+g_{2}\mathbb{K}[S,T]_{d-d_{2}}$
$=\langle g_{1}S^{d-d_{1}},\dots,g_{1}T^{d-d_{1}},g_{2}S^{d-d_{2}},\dots,g_{2}T^{d-d_{2}}\rangle.$
Keep in mind that the $d$ vectors
$v_{1}:=g_{2}S^{d-d_{2}},\;v_{2}:=g_{2}S^{d-d_{2}-1}T,\;\ldots,\;v_{d_{1}-1}:=g_{2}T^{d-d_{2}},$
$v_{d_{1}}:=g_{1}S^{d-d_{1}},\;\ldots,\;v_{d-1}:=g_{1}ST^{d-d_{1}-1},\;v_{d}:=g_{1}T^{d-d_{1}}\in\mathbb{K}[S,T]_{d}$
form a basis of the previously introduced subspace $V\subseteq\mathbb{K}[S,T]_{d}$. We consider $\mathbb{K}[s,t]$ as dual to $\mathbb{K}[S,T]_{d}$ by means of the non-degenerate inner product $\circ$ and choose $w_{1},w_{2},\ldots,w_{d}$ as a system dual to $v_{1},v_{2},\ldots,v_{d}$, so that $v_{i}\circ w_{j}=\delta_{ij}$ for $i,j=1,2,\ldots,d$. Observe that $v_{i}\circ f_{p}=0$ for $i=1,2,\ldots,d$ implies that $f_{p}\notin W:=\sum_{j=1}^{d}\mathbb{K}w_{j}$, so that $\mathbb{K}f_{p}+W=\mathbb{K}[s,t]_{d}$. Hence there is a $v_{0}\in\mathbb{K}[S,T]_{d}$ such that $\{v_{0},v_{1},\ldots,v_{d}\}$ is a basis dual to $\{f_{p},w_{1},\ldots,w_{d}\}$.
Since $\{f_{p},w_{1},w_{2},\ldots,w_{d}\}$ is a basis of $\mathbb{K}[s,t]_{d}$, for any $q\in\mathbb{P}^{d}$, we have a presentation
$f_{q}=\alpha_{0}f_{p}+\sum_{i=1}^{d}(v_{i}\circ f_{q})w_{i}=\alpha_{0}f_{p}+(g_{2}S^{d-d_{2}}\circ f_{q})w_{1}+\cdots+(g_{1}T^{d-d_{1}}\circ f_{q})w_{d}.$
If we set $\mathbb{P}^{d-1}:=\mathbb{P}(W)=\mathbb{P}(\mathbb{K}w_{1}+\cdots+\mathbb{K}w_{d})$, the projection map $\pi_{p}:\mathbb{P}^{d}\setminus\{p\}\to\mathbb{P}^{d-1}$ sends
$q$ $=[f_{q}]=[\alpha_{0}f_{p}+(g_{2}S^{d-d_{2}}\circ f_{q})w_{1}+\cdots+(g_{1}T^{d-d_{1}}\circ f_{q})w_{d}]$
$=[\alpha_{0}:g_{2}S^{d-d_{2}}\circ f_{q}:\cdots:g_{1}T^{d-d_{1}}\circ f_{q}]\in\mathbb{P}^{d}\setminus\{p\}$
to
$g_{2}S^{d-d_{2}}\circ f_{q}:\cdots:g_{1}T^{d-d_{1}}\circ f_{q}].$
Now, let $q=[\alpha^{d}:\alpha^{d-1}\beta:\cdots:\alpha\beta^{d-1}:\beta^{d}]\in\mathcal{C}_{d}$ with $\alpha:\beta]\in\mathbb{P}^{1}$. Then we have $f_{q}=(\alpha s+\beta t)^{d}$ and Lemma˜2.3 implies that
$g_{2}S^{d-d_{2}}\circ f_{q}=d!\,g_{2}(\alpha,\beta)\alpha^{d-d_{2}},\;\ldots\;,g_{1}T^{d-d_{1}}\circ f_{q}=d!\,g_{1}(\alpha,\beta)\beta^{d-d_{1}}.$
This proves that the curve $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})$ has the requested parametrization. ∎
By Theorem˜3.1, we can immediately see that $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})$ is contained in a rational surface scroll $\mathcal{S}=S(d-d_{2},d-d_{1})=S(d_{1}-2,d_{2}-2)\subset\mathbb{P}^{d-1}$ as a divisor. This provides a constructive proof of the existence of a surface scroll containing the projected curve $\mathcal{C}$. Moreover we can prove that the ideal of this curve can be generated by quadratic forms of rank at most four.
Corollary 3.2.
For $d\geq 4$, any projected curve $\pi_{p}(\mathcal{C}_{d})\subset\mathbb{P}^{d-1}$ for $p\in\mathbb{P}^{d}$ satisfies property $\operatorname{QR}(4)$.
Proof.
For points $p\in\mathcal{C}_{d}^{2}$, the projected curve $\pi_{p}(\mathcal{C}_{d})$ satisfies property $\operatorname{QR}(3)$, as shown in [16]. Therefore, we assume that $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}^{2}$. Let $f_{p}\in\mathbb{K}[s,t]_{d}$ be the binary form corresponding to $p\in\mathbb{P}^{d}\setminus\mathcal{C}_{d}^{2}$ with $d\geq 4$. By Theorem˜3.1, the projected curve $\pi_{p}(\mathcal{C}_{d})$ is projectively equivalent to the curve
$\mathcal{C}=\left\{[g_{2}(\alpha,\beta)\alpha^{d_{1}-2}:\dots:g_{2}(\alpha,\beta)\beta^{d_{1}-2}:g_{1}(\alpha,\beta)\alpha^{d_{2}-2}:\dots:g_{1}(\alpha,\beta)\beta^{d_{2}-2}]\mid[\alpha:\beta]\in\mathbb{P}^{1}\right\}$
where $(f_{p})^{\perp}=(g_{1},g_{2})$ with $\deg(g_{i})=d_{i}$, $(i=1,2)$, $d_{1}+d_{2}=d+2$, and $2\leq d_{1}\leq d_{2}$.
Therefore, the curve $\mathcal{C}$ is contained in a rational normal surface scroll
$\mathcal{S}$ $=\left\{[\alpha^{d_{1}-2}\gamma:\dots:\beta^{d_{1}-2}\gamma:\alpha^{d_{2}-2}\eta:\dots:\beta^{d_{2}-2}\eta]\mid[\alpha:\beta],[\gamma:\eta]\in\mathbb{P}^{1}\right\}$
$=S(d_{1}-2,d_{2}-2)$
and $\mathcal{C}$ is obtained by the section defined by $g_{1}x-g_{2}y=0$, where $x$ and $y$ are local coordinates on $\mathcal{S}$. In fact, $\mathcal{C}$ is a divisor on $\mathcal{S}$ in the class $H+2F$, where $H$ and $F$ are respectively a hyperplane section and a ruling of $\mathcal{S}$. Note that Greek letters are used for parameters in the parametrization of $\mathcal{S}$, whereas $x$ and $y$ are local coordinates on $\mathcal{S}$.
Therefore, the ideal $I(\pi_{p}(\mathcal{C}_{d}))_{2}$ is generated by the generators of $I(\mathcal{S})$ and quadratic forms corresponding to a basis of $(H-2F)(g_{1}x-g_{2}y)$. Note that the surface scroll $\mathcal{S}$ satisfies property $\operatorname{QR}(4)$ since it is obtained by the $2\times 2$ minors of a $2\times(d-2)$ matrix consisting of linear forms in the coordinate ring of $\mathcal{S}$.
Since H2FH-2F is spanned by $\{S^{d_{1}-j-4}T^{j}x\mid 0\leq j\leq d_{1}-4\}\cup\{S^{d_{2}-j-4}T^{j}y\mid 0\leq j\leq d_{2}-4\}$, $I(\pi_{p}(\mathcal{C}_{d}))_{2}/I(\mathcal{S})_{2}$ is generated by quadratic forms on $\mathcal{S}$ corresponding to
$\{S^{d_{1}-j-4}T^{j}g_{1}(S,T)x^{2}-S^{d_{1}-j-4}T^{j}g_{2}(S,T)xy\mid 0\leq j\leq d_{1}-4\}$
$\cup\;$ $\{S^{d_{2}-j-4}T^{j}g_{1}(S,T)xy-S^{d_{2}-j-4}T^{j}g_{2}(S,T)y^{2}\mid 0\leq j\leq d_{2}-4\}.$
Here, if $d_{1}\leq 3$ (resp. $d_{2}\leq 3$), then the first (resp. second) set is empty. Since we assume $d_{1}+d_{2}\geq 6$ and $d_{1}\leq d_{2}$, the only possible case with $d_{1}\leq 3$ or $d_{2}\leq 3$ is $(d_{1},d_{2})=(3,3)$, in which both sets are empty.
There exist two quadratic forms $q_{i}(S,T),i=1,2$ such that $q_{i}\mid g_{i}$ for $i=1,2$, since $g_{1}$ and $g_{2}$ are binary forms. Let $L_{1,j}$ be the linear form on $\mathcal{S}$ corresponding to $S^{d_{1}-j-4}T^{j}q_{1}x$, $L_{2}$ be the linear form on $\mathcal{S}$ corresponding to $(g_{1}/q_{1})x$, $L_{3,j}$ be the linear form on $\mathcal{S}$ corresponding to $S^{d_{1}-j-4}T^{j}q_{2}x$, and $L_{4}$ be the linear form on $\mathcal{S}$ corresponding to $(g_{2}/q_{2})y$. Then the quadratic generators corresponding to $\{S^{d_{1}-j-4}T^{j}g_{1}x^{2}-S^{d_{1}-j-4}T^{j}g_{2}xy\mid 0\leq j\leq d_{1}-4\}$ are
$\{L_{1,j}L_{2}-L_{3,j}L_{4}\mid 0\leq j\leq d_{1}-4\},$
which is a set of quadratic forms whose ranks are at most four.
Similarly, let $L_{5,j}$ be the linear form on $\mathcal{S}$ corresponding to $S^{d_{2}-j-4}T^{j}q_{1}y$ and $L_{6,j}$ be the linear form on $\mathcal{S}$ corresponding to $S^{d_{2}-j-4}T^{j}q_{2}y$. Then the quadratic generators corresponding to $\{S^{d_{2}-j-4}T^{j}g_{1}xy-S^{d_{2}-j-4}T^{j}g_{2}y^{2}\mid 0\leq j\leq d_{2}-4\}$ are
$\{L_{5,j}L_{2}-L_{6,j}L_{4}\mid 0\leq j\leq d_{2}-4\},$
which is a set of quadratic forms whose ranks are at most four. Thus, any quadratic generator of the ideal of $\mathcal{C}$ is a sum of quadratic forms of rank at most $4$. This implies that the curve $\pi_{p}(\mathcal{C}_{d})$ satisfies property $\operatorname{QR}(4)$. ∎
Example 3.3 (Monomial projections).
Let $e_{c}$ be the c-th standard basis for $\mathbb{P}^{d}$ with $d\geq 4$ and $1\leq c\leq\lfloor\frac{d}{2}\rfloor$. Then the corresponding binary form is given by $f:=f_{e_{c}}=s^{d-c}t^{c}\in\mathbb{K}[s,t]_{d}.$ Therefore, by Theorem˜3.1, the parametrization of the curve $\mathcal{C}=\pi_{e_{c}}(\mathcal{C}_{d})$ is
$\{[\alpha^{d}:\alpha^{d-1}\beta:\cdots:\alpha^{d-c+1}\beta^{c-1}:\alpha^{d-c-1}\beta^{c+1}:\cdots:\beta^{d}]\mid[\alpha:\beta]\in\mathbb{P}^{1}\}.$
Thus, $\mathcal{C}$ is a divisor of
$\mathcal{S}$ $=\{[\alpha^{c-1}\gamma:\alpha^{c-2}\beta\gamma:\cdots:\beta^{c-1}\gamma:\alpha^{d-c-1}\eta:\alpha^{d-c-2}\beta\eta:\cdots:\beta^{d-c-1}\eta]|[\alpha:\beta],[\gamma:\eta]\in\mathbb{P}^{1}\}$
$(=\mathcal{S}(c-1,d-c-1))$
and $\mathcal{C}$ is a section defined by $xT^{c+1}-yS^{d-c+1}$ in $\mathcal{S}$. Since H2FH-2F is spanned by $\{S^{c-j-3}T^{j}x\mid 0\leq j\leq c-3\}\cup\{S^{d-c-j-3}T^{j}y\mid 0\leq j\leq d-c-3\}$, $I(\mathcal{C})_{2}/I(\mathcal{S})_{2}$ is generated by quadratic forms on $\mathcal{S}$ corresponding to
$\{S^{c-j-3}T^{c+j+1}x^{2}-S^{d-j-2}T^{j}xy\mid 0\leq j\leq c-3\}$
$\cup\;$ $\{S^{d-c-j-3}T^{c+j+1}xy-S^{2d-2c-j-2}T^{j}y^{2}\mid 0\leq j\leq d-c-3\}$
Therefore,
[…]
and $I(\mathcal{C})$ is generated by $I(\mathcal{S})$ together with
$\{\,z_{j}z_{c-1}-z_{j}z_{c}\mid 0\leq j\leq c-3\,\}\;\cup\;\{\,z_{c-1}z_{c+j+2}-z_{c}z_{c+j}\mid 0\leq j\leq d-c-3\,\}.$
In fact, since
$I(\mathcal{C})=I(\mathcal{C}_{d})\cap\mathbb{K}[y_{0},\dots,y_{c-1},y_{c+1},\dots,y_{d}],$
where
[…]
this gives a constructive proof of a special case of the result in [5], which states that eliminating a variable from the ideal of a toric variety yields an ideal generated by binomials.
4. Monomial projections
In Corollary˜3.2, we have seen that the projected curves satisfy property $\operatorname{QR}(4)$. However, we will demonstrate that the projected curves also satisfy property $\operatorname{QR}(3)$ in certain cases.
Let $\mathcal{C}_{d,c}=\pi_{e_{c}}(\mathcal{C}_{d})\subset\mathbb{P}^{d-1}$ be the projected curve where $e_{c}$ for $0\leq c\leq d$ is the standard basis of $\mathbb{P}^{d}$. For some $c$, we can immediately see that property $\operatorname{QR}(3)$ is satisfied:
  1. (1)
    If $c=0\text{ or }d$, then $\mathcal{C}_{d,c}$ is a rational normal curve. Therefore, it satisfies property $\operatorname{QR}(3)$ (see [8]).
  2. (2)
    If $c=1\text{ or }d-1$, then $\mathcal{C}_{d,c}$ is a del Pezzo curve. Therefore, it satisfies property $\operatorname{QR}(3)$ (see [16]).
Thus, we assume $2\leq c\leq d-2$ and $d\geq 6$.
We distinguish between the cases $c=2\text{ or }c=d-2$ and $3\leq c\leq d-3$ because, in the case $c=2\text{ or }d-2$, $\mathcal{C}_{d,c}$ is not generated by quadrics, whereas for $3\leq c\leq d-3$, the curve $\mathcal{C}_{d,c}$ is generated by quadrics. For $\mathcal{C}_{d,2}$ (or $\mathcal{C}_{d,d-2}$), we instead consider the scheme $X_{d}=\mathcal{C}_{d,2}\cup\mathbb{L}$ where $\mathbb{L}$ is the trisecant line to $\mathcal{C}_{d,2}$. Note that $X_{d}$ is the scheme generated by quadrics on $\mathcal{C}_{d,2}$.
As we aim to determine whether the curves $\mathcal{C}_{d,c}=\pi_{c}(\mathcal{C}_{d})\subset\mathbb{P}^{d-1}$ satisfy property $\operatorname{QR}(3)$ and to provide additional numerical information beyond simply determining whether the property holds, we introduce the following notation:
Notation and Remark 4.1.
  1. (A)
    There is a combinatorial way to find a generating set of $I(\mathcal{C}_{d,c})$. Suppose I(C_d) = ( 2 ×2 minors of [x0x1⋯xd-1x1x2⋯xd] ) ⊂K[x_0,x_1,…,x_d], and define the index of a quadratic monomial $x_{i}x_{j}$ by $i+j$. Observe that
    1. (1)
      differences of any pairs $(x_{i}x_{j},x_{k}x_{l})$ of quadratic monomials with the same index are generators of $I(\mathcal{C}_{d})$ and
    2. (2)
      the generators of $I(\mathcal{C}_{d})_{2}$ that do not involve $x_{c}$ can be generators of $I(\mathcal{C}_{d,c})_{2}$.
    Therefore, the differences between all pairs of quadratic monomials $(x_{i}x_{j},x_{k}x_{l})$ with same index $i+j=k+l$ for $i,j,k,l\neq c$ are generators of $I(\mathcal{C}_{d,c})_{2}$. Moreover, for each integer $0\leq t\leq 2d$, let $u$ be the number of quadratic monomials $x_{i}x_{j}$ of index $t=i+j$ with $i,j\in\{0,1,\dots,d\}\setminus\{c\}$. Then one obtains u1u-1 linearly independent generators of $I(\mathcal{C}_{d,c})_{2}$.
    Among all polynomials there are $d+1$ quadratic monomials whose support contains $x_{c}$. For $1\leq c\leq d-1$, we lose exactly $d+1$ quadratic generators when eliminating $x_{c}$ from $I(\mathcal{C}_{d})$. Therefore, we have $\binom{d}{2}-(d+1)=\binom{d-1}{2}-2$ linearly independent quadratic generators in $I(\mathcal{C}_{d,c})_{2}$. These generators span $I(\mathcal{C}_{d,c})_{2}$ because $\dim_{\mathbb{K}}(I(\mathcal{C}_{d,c}))=\binom{d-2}{2}+d-4=\binom{d-1}{2}-2=\tfrac{1}{2}(d^{2}-3d-2).$
  2. (B)
    Let $X\subseteq\mathbb{P}^{d}$ be a projective scheme defined by quadratic forms and $I$ be the saturated ideal defining $X$.
    1. (1)
      For quadratic forms $q$ and $q^{\prime}$ in $I$, we write q ∼_I q’ if $q-q^{\prime}$ can be expressed as a sum of rank 3 generators of $I$. In particular,
      1. (i)
        $\sim_{I}$ is an equivalence relation and
      2. (ii)
        If $q\sim_{I}0$, then $q$ is a sum of rank $3$ quadratic forms in $I$.
    2. (2)
      For $0\leq c\leq d$, let $I_{d,c}$ denote the ideal of $\mathcal{C}_{d,c}=\pi_{e_{c}}(\mathcal{C}_{d})$. We write q ∼_(d,c) q’ for quadratic forms $q$ and $q^{\prime}$ in $I_{d,c}$ if $q-q^{\prime}$ can be expressed as a sum of rank 3 generators in $(I_{d,c})_{2}$.
We first claim that $\mathcal{C}_{d,c}$ for $d\geq 6$ and $3\leq c\leq d-3$ satisfies property $\operatorname{QR}(3)$. To show this claim, we begin by proving a series of computational lemmas. We recall from Section˜2.4, given a decomposition of the line bundle $\mathcal{O}_{X}(1)=A^{2}\otimes B$ of $X$,
$Q_{A,B}:H^{0}(X,A)\times H^{0}(X,A)\times H^{0}(X,B)\longrightarrow I(X)_{2}$
is defined by
$Q_{A,B}(s,t,h)=\varphi(s\otimes s\otimes h)\varphi(t\otimes t\otimes h)-\varphi(s\otimes t\otimes h)^{2}.$
Lemma 4.2.
Let $i$, $j$, $k$, and $l$ be integers such that $0\leq i<j\leq d$, i<jki<j-k, $i+j+k\equiv 0\pmod{2}$, and $c\neq i,i+k,j-k,j,\frac{i+j-k}{2},\frac{i+j+k}{2}$. Then,
$x_{i}x_{j}+x_{i+k}x_{j-k}-2x_{\frac{i+j-k}{2}}x_{\frac{i+j+k}{2}}\sim_{(d,c)}0.$
Proof.
We fix the decomposition of the line bundle $\mathcal{O}_{\mathbb{P}^{1}}(d)=A^{2}\otimes B$ where
$A=\mathcal{O}_{\mathbb{P}^{1}}\left(\frac{j-i-k}{2}\right)\text{ and }B=\mathcal{O}_{\mathbb{P}^{1}}(d-j+i+k).$
Then, in the notation introduced at the section 2.4 we have
$Q_{A,B}(S^{\frac{j-i-k}{2}},T^{\frac{j-i-k}{2}},S^{d-j}T^{i}(S^{k}+T^{k}))$
$=$ $(x_{i}+x_{i+k})(x_{j-k}+x_{j})-(x_{\frac{i+j-k}{2}}+x_{\frac{i+j+k}{2}})^{2}$
$=$ $(x_{i}x_{j-k}-x_{\frac{i+j-k}{2}}^{2})+(x_{i+k}x_{j}-x_{i+j+k}^{2})+x_{i}x_{j}+x_{i+k}x_{j-k}-2x_{\frac{i+j-k}{2}}x_{\frac{i+j+k}{2}}.$
Since $(x_{i}x_{j-k}-x_{\frac{i+j-k}{2}}^{2})$ and $(x_{i+k}x_{j}-x_{i+j+k}^{2})$ are rank $3$ generators in $I_{d,c}$, we have $x_{i}x_{j}+x_{i+k}x_{j-k}-2x_{\frac{i+j-k}{2}}x_{\frac{i+j+k}{2}}\sim_{(d,c)}0$. ∎
Lemma 4.3.
For each $0\leq i\leq d-6$, the two quadratic forms $x_{i}x_{i+5}-x_{i+1}x_{i+4}$ and $x_{i+1}x_{i+6}-x_{i+2}x_{i+5}$ satisfy
$(x_{i}x_{i+5}-x_{i+1}x_{i+4})\sim_{(d,i+3)}0\quad\text{and}\quad(x_{i+1}x_{i+6}-x_{i+2}x_{i+5})\sim_{(d,i+3)}0.$
Proof.
For a decomposition $\mathcal{O}_{\mathbb{P}^{1}}(d)=\mathcal{O}_{\mathbb{P}^{1}}(2)^{2}\otimes\mathcal{O}_{\mathbb{P}^{1}}(d-4)$ we have (in the notation introduced at the section 2.4)
$\;Q_{\mathcal{O}_{\mathbb{P}^{1}}(2),\mathcal{O}_{\mathbb{P}^{1}}(d-4)}(S^{2},\;ST+aT^{2},\;S^{d-i-4}T^{i}-2aS^{d-i-5}T^{i+1}+2a^{2}S^{d-i-6}T^{i+2})$
$=$ $\;(x_{i}x_{i+2}-x_{i+1}^{2})+a^{2}(-x_{i}x_{i+4}+x_{i+2}^{2})+2a^{3}(x_{i}x_{i+5}-x_{i+1}x_{i+4})$
$+2a^{4}(x_{i}x_{i+6}-2x_{i+1}x_{i+5}+x_{i+2}x_{i+4})+4a^{5}(-x_{i+1}x_{i+6}+x_{i+2}x_{i+5})$
$+4a^{6}(x_{i+2}x_{i+6}-x_{i+4}^{2})$
for $a\in\mathbb{K}$. Since $x_{i}x_{i+6}+x_{i+2}x_{i+4}-2x_{i+1}x_{i+5}\sim_{(d,i+3)}(x_{i}+x_{i+4})(x_{i+2}+x_{i+6})-(x_{i+1}+x_{i+5})^{2}$ and since the forms $x_{i}x_{i+2}-x_{i+1}^{2}$, $x_{i}x_{i+4}-x_{i+2}^{2}$, $x_{i+2}x_{i+6}-x_{i+4}^{2}$ and $(x_{i}+x_{i+4})(x_{i+2}+x_{i+6})-(x_{i+1}+x_{i+5})^{2}$ are rank three quadratic generators of $I_{d,i+3}$, it follows that
$2a^{3}(x_{i}x_{i+5}-x_{i+1}x_{i+4})-4a^{5}(x_{i+1}x_{i+6}-x_{i+2}x_{i+5})\sim_{(d,i+3)}0$
for any $a\in\mathbb{K}$.
Thus $x_{i}x_{i+5}-x_{i+1}x_{i+4}\sim_{(d,i+3)}0$ and $x_{i+1}x_{i+6}-x_{i+2}x_{i+5}\sim_{(d,i+3)}0$. ∎
Lemma 4.4.
For each $0\leq i\leq d-5$, the two quadratic forms $x_{i}x_{i+5}-x_{i+2}x_{i+3}$ and $x_{i+2}x_{i+5}-x_{i+3}x_{i+4}$ satisfy
$(x_{i}x_{i+5}-x_{i+2}x_{i+3})\sim_{(d,i+1)}0\text{ and }(x_{i+2}x_{i+5}-x_{i+3}x_{i+4})\sim_{(d,i+1)}0.$
Proof.
For a decomposition $\mathcal{O}_{\mathbb{P}^{1}}(d)=\mathcal{O}_{\mathbb{P}^{1}}(2)^{2}\otimes\mathcal{O}_{\mathbb{P}^{1}}(d-4)$ we have (in the notation introduced at the end of Section˜2.4),
$Q_{\mathcal{O}_{\mathbb{P}^{1}}(2),\mathcal{O}_{\mathbb{P}^{1}}(d-4)}(S^{2}+aST,\;T^{2},\;S^{d-i-4}T^{i}-2aS^{d-i-5}T^{i+1})$
$=\quad$ $(x_{i}-3a^{2}x_{i+2}-2a^{3}x_{i+3})(x_{i+4}-2ax_{i+5})-(x_{i+2}-ax_{i+3}-2a^{2}x_{i+4})^{2}$
$=\quad$ $(x_{i}x_{i+4}-x_{i+2}^{2})+2a(-x_{i}x_{i+5}+x_{i+2}x_{i+3})+a^{2}(x_{i+2}x_{i+4}-x_{i+3}^{2})$
$+6a^{3}(x_{i+2}x_{i+5}-x_{i+3}x_{i+4})+4a^{4}(x_{i+3}x_{i+5}-x_{i+4}^{2})$
Since $x_{i}x_{i+4}-x_{i+2}^{2}$, $x_{i+2}x_{i+4}-x_{i+3}^{2}$, and $x_{i+3}x_{i+5}-x_{i+4}^{2}$ are rank 3 quadratic generators in $I_{d,i+3}$,
$2a(-x_{i}x_{i+5}+x_{i+2}x_{i+3})+6a^{3}(x_{i+2}x_{i+5}-x_{i+3}x_{i+4})\sim_{(d,i+1)}0$
for any $a\in\mathbb{K}$.
Hence $x_{i}x_{i+5}-x_{i+2}x_{i+3}\sim_{(d,i+1)}0$ and $x_{i+2}x_{i+5}-x_{i+3}x_{i+4}\sim_{(d,i+1)}0$. ∎
Theorem 4.5.
For $d\geq 6$ and $3\leq c\leq d-3$, the ideal of $\mathcal{C}_{d,c}$ is generated by quadratic forms and satisfies property $\operatorname{QR}(3)$.
Proof.
Since $I_{d,c}\cong I_{d,d-c}$ for any $0\leq c\leq d$, we assume that $3\leq c\leq\lfloor\frac{d}{2}\rfloor$. We recall that the quadratic generators of $I_{d,c}$ are given by binomials of the form $x_{i}x_{j}-x_{k}x_{l}$ where $i+j=k+l$ and $i,j,k,l\neq c$. There are four initial cases for using induction:
$(d,c)=(6,3),(7,3),(8,3),\text{ and }(8,4).$
Initial Case 1 : $d=6,c=3$
index monomials
22 $x_{0}x_{2},x_{1}^{2}$
33 $x_{1}x_{2}$
44 $x_{0}x_{4},x_{2}^{2}$
55 $x_{0}x_{5},x_{1}x_{4}$
66 $x_{0}x_{6},x_{1}x_{5},x_{2}x_{4}$
77 $x_{1}x_{6},x_{2}x_{5}$
88 $x_{2}x_{6},x_{4}^{2}$
99 $x_{4}x_{5}$
$10$ $x_{4}x_{6},x_{5}^{2}$
Table 1. List of monomials with given indices.
The index $2,4,8,10$ generators are $x_{0}x_{2}-x_{1}^{2},x_{0}x_{4}-x_{2}^{2},x_{2}x_{6}-x_{4}^{2},x_{4}x_{6}-x_{5}^{2}$ and they are already rank $3$ quadratic generators. The remaining generators are index $5,6,7$ generators. (See Table˜1.) By Lemma˜4.3 with $i=0$, index $5$ generator $x_{0}x_{5}-x_{1}x_{4}\sim_{(6,3)}0$ and index $7$ generator $x_{1}x_{6}-x_{2}x_{5}\sim_{(6,3)}0$. By Lemma˜4.2, $x_{0}x_{6}+x_{2}x_{4}-2x_{2}x_{4}=x_{0}x_{6}-x_{2}x_{4}\sim_{(6,3)}0$ and $x_{0}x_{6}+x_{2}x_{4}-2x_{1}x_{5}=(x_{0}x_{6}-x_{2}x_{4})+2(x_{2}x_{4}-x_{1}x_{5})\sim_{(6,3)}0$. Hence $I_{6,3}$ satisfies property $\operatorname{QR}(3)$.
Initial Case 2 : $d=7,c=3$.
Since $I_{7,3}$ contains the same generators as $I_{6,3}$, we only need to examine the new generators that include $x_{7}$ as supports. The generators of index $\geq 10$ are given by the $2\times 2$ minors of the matrix
$\begin{pmatrix}x_{4}&x_{5}&x_{6}\\ x_{5}&x_{6}&x_{7}\end{pmatrix}$
These minors correspond to the generators of a rational normal curve and are spanned by rank $3$ quadratic forms.
The remaining monomials are $x_{0}x_{7}$, $x_{1}x_{7}$, and $x_{2}x_{7}$. We need to find quadratic generators containing these monomials that can be expressed as a sum of rank $3$ quadrics without using $x_{3}$. By Lemma˜4.2, we have $x_{0}x_{7}+x_{2}x_{5}-2x_{1}x_{6}\sim_{(7,3)}0$. The generator $x_{1}x_{7}-x_{4}^{2}$ is already a rank $3$ quadratic form. In addition, we obtain $x_{2}x_{7}-x_{4}x_{5}\sim_{(7,3)}0$ by applying Lemma˜4.4 with $i=2$. Thus, we conclude that $I_{7,3}$ satisfies property $\operatorname{QR}(3)$.
Initial Case 3 : $d=8,c=3$
As in the previous case, since $I_{8,3}$ contains the same generators as $I_{7,3}$, we only need to check the new generators that involve $x_{8}$. The generators of index $\geq 11$ are given by the $2\times 2$ minors of the matrix
$\begin{pmatrix}x_{4}&x_{5}&x_{6}&x_{7}\\ x_{5}&x_{6}&x_{7}&x_{8}\end{pmatrix}$
These minors are generators of a rational normal curve, which already satisfies property $\operatorname{QR}(3)$.
The remaining monomials are $x_{0}x_{8}$, $x_{1}x_{8}$, and $x_{2}x_{8}$. The binomials $x_{0}x_{8}-x_{4}^{2}$ and $x_{2}x_{8}-x_{5}^{2}$ are already rank $3$ generators. By Lemma˜4.2, we have $x_{1}x_{8}+x_{2}x_{7}-2x_{4}x_{5}\sim_{(8,3)}0$. Thus, $I_{8,3}$ satisfies property $\operatorname{QR}(3)$.
Initial Case 4 : $d=8,c=4$
Since $I_{7,4}\cong I_{7,3}$, we only need to check the new generators that involve $x_{8}$. The generators of index $\geq 12$ are given by the $2\times 2$ minors of the matrix
$\begin{pmatrix}x_{5}&x_{6}&x_{7}\\ x_{6}&x_{7}&x_{8}\end{pmatrix}$
These minors correspond to generators of a rational normal curve, which already satisfies property $\operatorname{QR}(3)$.
The remaining monomials are $x_{0}x_{8}$, $x_{1}x_{8}$, $x_{2}x_{8}$, and $x_{3}x_{8}$. By Lemma˜4.2, we have $x_{0}x_{8}+x_{2}x_{6}-2x_{3}x_{5}\sim_{(8,4)}0$ and $x_{1}x_{8}+x_{3}x_{6}-2x_{2}x_{7}\sim_{(8,4)}0$. In addition, the binomial $x_{2}x_{8}-x_{5}^{2}$ is already a rank $3$ quadratic form. Moreover, we obtain $x_{3}x_{8}-x_{5}x_{6}\sim_{(8,4)}0$ by Lemma˜4.4 with $i=3$. Hence, $I_{8,4}$ satisfies property $\operatorname{QR}(3)$.
For the inductive hypothesis, assume that the ideal $I_{d-1,c}$ satisfies property $\operatorname{QR}(3)$ for $d\geq 9$ and $c$ such that $3\leq c\leq\lfloor\frac{d}{2}\rfloor$. Since $I_{d,c}$ contains the generators of $I_{d-1,c}$, we only need to find the rank $3$ quadratic forms that involve $x_{d}$. The generators of index $\geq d+c$ are those of a rational normal curve, so the remaining task is to check the quadratic generators with indices between $d$ and d+c1d+c-1. We will consider the cases based on the parity of $d$ and $c$.
Case 1-1 : $d$ is odd and $3\leq c<\frac{d-1}{2}$.
index quadratic forms involving $x_{d}$
$d$ $x_{0}x_{d}+x_{1}x_{d-1}-2x_{\frac{d-1}{2}}x_{\frac{d+1}{2}}$
$d+1$ $x_{1}x_{d}-x_{\frac{d+1}{2}}^{2}$
$d+2$ $x_{2}x_{d}+x_{3}x_{d-1}-2x_{\frac{d+1}{2}}x_{\frac{d+3}{2}}$
$d+2i$ $x_{2i}x_{d}+x_{2i+1}x_{d-1}-2x_{i+\frac{d-1}{2}}x_{i+\frac{d+1}{2}}$
$d+2i+1$ $x_{2i+1}x_{d}-x_{i+\frac{d+1}{2}}^{2}$
[…]
Table 2. Quadratic generators in $I_{d,c}$ for odd $d$ and $3\leq c<\frac{d-1}{2}$.
Since $\frac{d-1}{2}>c$, all binomials in Table˜2 can be expressed as sums of rank 3 quadrics by Lemma˜4.2.
If d+c1d+c-1 is an even number, the quadratic form $x_{c-1}x_{d}-x_{\frac{d+c-1}{2}}^{2}$ is already a rank 3 quadratic form. If d+c1d+c-1 is odd, we use the following binomial:
$x_{c-1}x_{d}+x_{c+2}x_{d-3}-2x_{\frac{d+c-4}{2}}x_{\frac{d+c+2}{2}}\sim_{(d,c)}0.$
The relation holds because c<d4c<d-4, as implied by $c\leq\frac{d-1}{2}$ and $d\geq 9$.
Case 1-2 : $d$ is odd and $c=\frac{d-1}{2}$.
Since $c=\frac{d-1}{2}$, all quadrics in Table˜2 except the index $d$ part does not contain any $x_{c}$. So they can be expressed as sums of rank $3$ quadratic generators by Lemma˜4.2. The only issue arises at index $d$. To resolve this, we replace the form with the following quadratic generator without $x_{c}$ from Lemma˜4.2:
$x_{0}x_{d}+x_{3}x_{d-3}-2x_{\frac{d-3}{2}}x_{\frac{d+3}{2}}\sim_{(d,c)}0.$
Case 2-1 : $d$ is even and $3\leq c<\frac{d}{2}$.
index quadrics containing $x_{d}$
$d$ $x_{0}x_{d}-x_{\frac{d}{2}}^{2}$
$d+1$ $x_{1}x_{d}+x_{2}x_{d-1}-2x_{\frac{d}{2}}x_{\frac{d+2}{2}}$
$d+2i$ $x_{2i}x_{d}-x_{i+\frac{d}{2}}^{2}$
$d+2i+1$ $x_{2i+1}x_{d}+x_{2i+2}x_{d-1}-2x_{i+\frac{d}{2}}x_{i+\frac{d+2}{2}}$
[…]
Table 3. Quadratic generators in $I_{d,c}$ for even $d$ and $3\leq c<\frac{d}{2}$.
Since $\frac{d}{2}>c$, all quadrics listed in Table˜3 are spanned by rank 3 quadratic forms by Lemma˜4.2. For the index d+c1d+c-1, if $c$ is odd, the form $x_{c-1}x_{d}-x_{\frac{d+c-1}{2}}^{2}$ is already a rank $3$ quadratic generator. However, if $c$ is even, then we can use
$x_{c-1}x_{d}+x_{c+2}x_{d-3}-2x_{\frac{d+c-4}{2}}x_{\frac{d+c+2}{2}}\sim_{(d,c)}0.$
as a generator because c<d4c<d-4.
Case 2-2 : $d$ is even and $c=\frac{d}{2}$.
Since $c=\frac{d}{2}$, all quadrics in Table˜3 except index $d$ and $d+1$ does not contain any $x_{c}$. So they can be expressed as sums of rank $3$ quadratic generators by Lemma˜4.2. The issue arises only at indices $d$ and $d+1$. To resolve this, we replace these forms with the following quadratic generators from Lemma˜4.2:
$x_{0}x_{d}+x_{2}x_{d-2}-2x_{\frac{d-2}{2}}x_{\frac{d+2}{2}}\sim_{(d,c)}0\text{ and }x_{1}x_{d}+x_{4}x_{d-3}-2x_{\frac{d-2}{2}}x_{\frac{d+4}{2}}\sim_{(d,c)}0.$
Proof of Theorem˜1.1.
Combine Corollary˜3.2 and Theorem˜4.5. ∎
Now, we consider the case $c=2$. In this case, we have $\operatorname{rank}_{\mathcal{C}_{d}}(p)=3$ and hence $\mathcal{C}$ admits a unique 33–secant line $\mathbb{L}$ and the scheme $X_{d}:=\mathbb{L}\cup\mathcal{C}$ is defined by the quadrics contained in the vanishing ideal $I(\mathcal{C})$ the curve $\mathcal{C}$ (see the comment prior to Theorem 1.3). Our aim is to show, that in this case the scheme $X_{d}$ does not satisfy property $\operatorname{QR}(3)$.
Proposition 4.6.
Let $X_{d}$ be the projective scheme defined by the quadrics on $\mathcal{C}_{d,2}$ for $d\geq 5$. Then we have the following:
  1. (1)
    $X_{d}$ does not satisfy property $\operatorname{QR}(3)$; moreover, in the notation introduced in ˜2.6, we have
  2. (2)
    $\delta(X_{d},4)-\delta(X_{d},3)=2$.
Proof.
We first prove that $\delta(X_{d},4)-\delta(X_{d},3)\geq 2$ for $d\geq 5$. Let $Q=\sum_{i=0}^{\frac{1}{2}(d^{2}-3d-4)}a_{i}Q_{i}$ where $a_{i}\in\mathbb{K}$ for $0\leq i\leq\frac{1}{2}(d^{2}-3d-4)$ and let $M_{Q}$ be the corresponding $d\times d$ symmetric matrix. We claim that there exist at least two independent linear forms in $\sqrt{I(4,M_{Q})}$.
We fix $Q_{0}=x_{0}x_{4}-x_{1}x_{3}$ and $Q_{1}=x_{0}x_{5}-x_{1}x_{4}$. Consider the $4\times 4$ submatrix of $M_{Q}$ obtained by selecting the rows and columns of $M_{Q}$ which correspond to $x_{0},x_{1},x_{3},x_{4}$, respectively:
$\begin{array}[]{c c}&\begin{array}[]{c c c c}x_{0}&x_{1}&x_{3}&x_{4}\\ \end{array}\\ \begin{array}[]{c c c c}x_{0}\\ x_{1}\\ x_{3}\\ x_{4}\end{array}&\begin{bmatrix}0&0&0&a_{0}\\ 0&0&-a_{0}&*\\ 0&-a_{0}&*&*\\ a_{0}&*&*&*\end{bmatrix}\end{array}$.
The determinant of this submatrix is $a_{0}^{4}$ which is contained in $I(4,M_{Q})$. Hence, $a_{0}\in\sqrt{I(4,M_{Q})}$. Now, consider the $4\times 4$ submatrix of $M_{Q}$ obtained by selecting the rows and columns of $M_{Q}$ which correspond to $x_{0},x_{1},x_{4},x_{5}$, respectively:
$\begin{array}[]{c c}&\begin{array}[]{c c c c}x_{0}&x_{1}&x_{4}&x_{5}\\ \end{array}\\ \begin{array}[]{c c c c}x_{0}\\ x_{1}\\ x_{4}\\ x_{5}\end{array}&\begin{bmatrix}0&0&a_{0}&a_{1}\\ 0&0&-a_{1}&*\\ a_{0}&-a_{1}&*&*\\ a_{1}&*&*&*\end{bmatrix}\end{array}$.
The determinant of this submatrix is $g\cdot a_{0}+a_{1}^{4}\in I(4,M_{Q})$ for a cubic polynomial $g$. Since $a_{0}\in\sqrt{I(4,M_{Q})}$, it follows that $a_{1}\in\sqrt{I(4,M_{Q})}$. Therefore, by ˜2.6, $\delta(X_{5},4)-\delta(X_{5},3)\geq 2$.
Now, we show that $\delta(X_{d},4)-\delta(X_{d},3)\leq 2$ by induction on $d\geq 5$. For the initial cases, we consider $d=5$ and $d=6$. When $d=5$, we have
$(I_{5,2})_{2}=\langle x_{4}^{2}-x_{3}x_{5},x_{1}x_{4}-x_{0}x_{5},x_{3}^{2}-x_{1}x_{5},x_{1}x_{3}-x_{0}x_{4}\rangle\subset\mathbb{K}[x_{0},x_{1},x_{3},x_{4},x_{5}].$
Thus, $\delta(X_{5},4)-\delta(X_{5},3)\leq 2$.
Suppose $d=6$. There are $8$ quadratic generators of $I(X_{6})$ as follows:
$Q_{1}=x_{0}x_{4}-x_{1}x_{3},Q_{2}=x_{0}x_{5}-x_{1}x_{4},Q_{3}=x_{0}x_{6}-x_{3}^{2},Q_{4}=x_{1}x_{5}-x_{3}^{2},$
$Q_{5}=x_{1}x_{6}-x_{3}x_{4},Q_{6}=x_{3}x_{5}-x_{4}^{2},Q_{7}=x_{3}x_{6}-x_{4}x_{5},Q_{8}=x_{4}x_{6}-x_{5}^{2}.$
Hence $\delta(X_{6},4)=8$.
We claim that there are $6$ generators can be expressed as a sum of rank $3$ generators without $x_{2}$. We can check that
  1. (1)
    $Q_{3}=x_{0}x_{6}-x_{3}^{2},Q_{4}=x_{1}x_{5}-x_{3}^{2},Q_{6}=x_{3}x_{5}-x_{4}^{2},x_{8}=x_{4}x_{6}-x_{5}^{2}$ are rank 33 itself.
  2. (2)
    $Q_{5}=x_{1}x_{6}-x_{3}x_{4}\sim_{(6,2)}0$ by Lemma˜4.4 with $i=1$.
  3. (3)
    $Q_{6}=x_{3}x_{6}-x_{4}x_{5}=x_{3}x_{6}+x_{4}x_{5}-2x_{4}x_{5}\sim_{(6,2)}0$ by Lemma˜4.2.
Thus, by ˜2.6 we have $\delta(X_{6},4)-\delta(X_{6},3)=2$.
For the inductive hypothesis, let $d\geq 7$ and suppose that $\delta(X_{d-1},4)-\delta(X_{d-1},3)=2$. It is enough to show that the generators in $I_{2,d}\setminus I_{2,d-1}$ can be expressed as sums of rank $3$ quadratic generators.
  1. Case 1 :
    $d$ is odd
For the indices $d$ and $d+1$, the following hold:
$x_{0}x_{d}+x_{1}x_{d-1}-2x_{\frac{d-1}{2}}x_{\frac{d+1}{2}}\sim_{(d,2)}0\text{ and }x_{1}x_{d}-x_{\frac{d+1}{2}}^{2}\sim_{(d,2)}0.$
For indices greater than or equal to $d+2$, all generators of $I_{2,d}$ can be obtained from the $2\times 2$ minors of the matrix
$\begin{pmatrix}x_{3}&x_{4}&\cdots&x_{d-1}\\ x_{4}&x_{5}&\cdots&x_{d}\end{pmatrix}$.
These generators can be expressed as sums of rank $3$ quadratic forms as the rational normal curves satisfy property $\operatorname{QR}(3)$.
  1. Case 2 :
    $d$ is even
For the indices $d$ and $d+1$, we have
$x_{0}x_{d}-x_{\frac{d}{2}}^{2}\sim_{(d,2)}0\text{ and }x_{1}x_{d}+x_{4}x_{d-3}-2x_{\frac{d-2}{2}}x_{\frac{d+4}{2}}\sim_{(d,2)}0.$
For indices greater than or equal to $d+2$, the generators are spanned by rank $3$ quadratic forms by the same reasoning above. ∎
5. $\operatorname{QR}(k)$ properties of $\pi_{p}(\mathcal{C}_{d})$ for a rank $3$ point $p$.
In this section, we investigate whether the space of quadrics on the projected curve $\pi_{p}(\mathcal{C}_{d})$ can be spanned by rank 3 quadratic forms, provided the projection center $p$ is of rank 3. We recall that $\pi_{p}(\mathcal{C}_{d})$ satisfies property $\operatorname{QR}(3)$, for points $p$ of rank at most two.
Let $\operatorname{rank}_{\mathcal{C}_{d}}(p)=l$. Then $\mathcal{C}$ is contained in S(l2,dl)S(l-2,d-l) (cf. [14, Theorem 1.1.(1)]). If $d=4$, then $l=3$ and $\mathcal{C}$ is contained in $S(1,1)$. In this case, $\mathcal{C}$ satisfies only one quadratic equation, namely that of $S(1,1)$, and hence has infinitely many trisecant lines. Therefore, we may assume $d\geq 5$.
Lemma 5.1.
Suppose $p\in\mathcal{C}_{d}^{3}\setminus\mathcal{C}_{d}^{2}$ with $d\geq 5$, and let $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})\subset\mathbb{P}^{d-1}$ be the image of $\mathcal{C}_{d}$ under the linear projection centered at $p$.
  1. (1)
    There exists the unique trisecant line $\mathbb{L}:=S(1)$ of $\mathcal{C}$.
  2. (2)
    $I(\mathcal{C}\cup\mathbb{L})$ is generated by $I(\mathcal{C})_{2}$.
Proof.
  1. (1)
    If $d\geq 5$, then $\mathcal{C}$ is contained in S(1,d3)S(1,d-3) and $\mathcal{C}\equiv H+2F$. Hence $S(1)\equiv H-(d-3)F$ is a trisecant line to $\mathcal{C}$. Conversely, if there exists a trisecant line, then it must be contained in S(1,d3)S(1,d-3), since the ideal of S(1,d3)S(1,d-3) is generated by quadrics. Consequently, $\mathbb{L}=S(1)$ is the unique trisecant line of $\mathcal{C}$.
  2. (2)
    Remark that
    $\mathcal{C}\cup\mathbb{L}\;\equiv\;2H+(5-d)F.$
    1. (a)
      By Theorem 4.3.(1) in [15], $\mathcal{C}\cup\mathbb{L}$ is aCM. Then, by Proposition 3.4 in [15], we know that $I(\mathcal{C}\cup\mathbb{L})$ is generated by quadrics.
    2. (b)
      Clearly, $I(\mathcal{C}\cup\mathbb{L})_{2}\subseteq I(\mathcal{C})_{2}$. On the other hand, $I(\mathcal{C})_{2}\subseteq I(\mathcal{C}\cup\mathbb{L})_{2}$ since $\mathbb{L}$ is the trisecant line to $\mathcal{C}$.
    In consequence, we obtain $I(\mathcal{C})_{2}=I(\mathcal{C}\cup\mathbb{L})_{2}$, and hence $I(\mathcal{C}\cup\mathbb{L})$ is generated by $I(\mathcal{C})_{2}$.
Let $X$ be the scheme defined by the quadrics in $\mathcal{C}$, i.e., $X=\mathcal{C}\cup\mathbb{L}$.
There are three cases depending on how $\mathcal{C}$ and $\mathbb{L}$ intersect, and each case can be described in terms of the multiplicity type of the lower-degree generator $F_{l}$ of the apolar ideal $(f_{p})^{\perp}=(F_{l},F_{h})$:
  1. (1)
    $F_{l}=\ell^{3}$ for a linear binary form $\ell$ $\iff$ $\mathcal{C}$ intersects $\mathbb{L}$ at a triple point.
  2. (2)
    $F_{l}=\ell_{1}^{2}\ell_{2}$ for linearly independent linear binary forms $\ell_{1},\ell_{2}$
    $\iff$ $\mathcal{C}$ intersects $\mathbb{L}$ at a double point and a simple point.
  3. (3)
    $F_{l}=\ell_{1}\ell_{2}\ell_{3}$ for linearly independent linear binary forms $\ell_{1},\ell_{2},\ell_{3}$
    $\iff$ $\mathcal{C}$ intersects $\mathbb{L}$ at three distinct simple points.
By [10, Lemma 1.31 and Proposition 1.36], given the degree of the lower-degree generator of $f_{p}^{\perp}$, we have a unique general additive decomposition of $f_{p}$. Moreover, each term in this general additive decomposition corresponds to a point on $C_{d}$ (with multiplicities). For example, in the case of (2), from the additive decomposition of $f_{p}$, the point $p$ is spanned by a double point and a simple point on $C_{d}$ (corresponding to $\ell_{1}^{2}$ and $\ell_{2}$, respectively). Therefore, the trisecant line $\mathbb{L}$ must intersect $\pi_{p}(C_{d})$ at the image of the double point and the simple point.
Example 5.2.
Let $e_{i}$ for $i=0,\dots,d$ be the torus-fixed points in $\mathbb{P}^{d}$ for $d\geq 5$. Let $\mathbb{L}$ be the trisecant line to $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})$.
  1. (1)
    If $p=e_{2}$, i.e., $(f_{p})^{\perp}=(S^{3},T^{d-1})$, then $\mathcal{C}\cap\mathbb{L}$ consists of a triple point.
  2. (2)
    If $p=e_{1}+e_{d}$, i.e., $(f_{p})^{\perp}=(S^{2}T,dS^{d-1}-T^{d-1})$, then $\mathcal{C}\cap\mathbb{L}$ consists of a double point and a simple point.
  3. (3)
    If $p=\sum_{i=0}^{d}(-1)^{i}e_{i}+(-1)^{d}\sum_{i=0}^{d}e_{i}+(-1)^{d}2e_{d}$, i.e., $(f_{p})^{\perp}=(S^{3}-ST^{2},T^{d-1})$, then $\mathcal{C}\cap\mathbb{L}$ consists of three distinct simple points.
We claim that the scheme $X$ does not satisfy $\operatorname{QR}(3)$ if $\mathcal{C}\cap\mathbb{L}$ contains a non-simple point (i.e., the first two cases). However, in the third case where $\mathcal{C}\cap\mathbb{L}$ consists of three simple points, the scheme $X$ possibly satisfies property $\operatorname{QR}(3)$.
5.1. A triple point.
Let $p$ be a point in $\mathcal{C}_{d}^{3}\setminus\mathcal{C}_{d}^{2}$ for $d\geq 5$ such that $\pi_{p}(\mathcal{C}_{d})\cap\mathbb{L}$ consists of a point with multiplicity $3$. This implies that $(f_{p})^{\perp}=(\ell^{3},g)\subset\mathbb{K}[S,T]$ with $\deg\ell=1$ and $\deg g=d-1$. We claim that the scheme $X:=\pi_{p}(\mathcal{C}_{d})\cup L$ for $d\geq 5$ does not satisfy property $\operatorname{QR}(3)$.
Proposition 5.3.
For $d\geq 5$, let $p$ be a point in $\mathbb{P}^{d}$ such that $(f_{p})^{\perp}=(\ell^{3},g)$ with $\deg\ell=1$ and $\deg g=d-1$.
  1. (1)
    The scheme $X$ does not satisfy property $\operatorname{QR}(3)$. Moreover,
  2. (2)
    $\delta(X,4)-\delta(X,3)\geq 2.$
Proof.
We first note that it suffices to consider the point $p\in\mathbb{P}^{d}$ corresponding to $f_{p}\in\mathbb{K}[s,t]_{d}$ such that $(f_{p})^{\perp}=(T^{3},g(S,T))\subset\mathbb{K}[S,T]$ by using an automorphism of $\mathbb{P}^{1}$ that sends the linear form $\ell$ to $T$ in $\mathbb{K}[S,T]_{1}$. Moreover, we can take any higher degree generator of $(f_{p})^{\perp}$ among any forms in $\{g(S,T)-T^{3}h(S,T)\mid h(S,T)\in\mathbb{K}[S,T]_{d-4}\}$. Therefore, we can assume
$(f_{p})^{\perp}=(T^{3},\alpha_{0}S^{d-1}+\alpha_{1}S^{d-2}T+\alpha_{2}S^{d-3}T^{2})$
for some $\alpha_{0},\alpha_{1},\alpha_{2}\in\mathbb{K}$.
We claim that $\sqrt{I(4,M_{Q})}$ contains at least one linear form. Following the proof of Corollary˜3.2, $I(X)$ is generated by the quadratic generators of
[…]
and
$\left\{z_{1}z_{j+4}-\alpha_{0}z_{2}z_{j+2}-\alpha_{1}z_{2}z_{j+3}-\alpha_{2}z_{2}z_{j+4}\mid 0\leq j\leq d-5\right\}.$
Let $Q=\sum_{i=0}^{N}a_{i}Q_{i}$ be the general element in $\mathbb{K}[a_{0},\dots,a_{N}]$ where $Q_{i}$ are the quadratic generators of $I(X)$ for $0\leq i\leq N=\binom{d-1}{2}-3$. Without loss of generality, suppose that $Q_{0}=-z_{1}z_{2}+z_{0}z_{3}$ and $Q_{1}=-z_{1}z_{3}+z_{0}z_{4}$.
Let $M_{1}$ be the $4\times 4$ submatrix of $M_{Q}$ obtained by selecting the 0th, 1st, 2nd, and 3rd rows and columns of $M_{Q}$:
$M_{1}=\begin{array}[]{c c}&\begin{array}[]{c c c c}z_{0}&z_{1}&z_{2}&z_{3}\\ \end{array}\\ \begin{array}[]{c c c c}z_{0}\\ z_{1}\\ z_{2}\\ z_{3}\end{array}&\begin{bmatrix}0&0&0&a_{0}\\ 0&0&-a_{0}&*\\ 0&-a_{0}&*&*\\ a_{0}&*&*&*\end{bmatrix}\end{array}$.
where * denotes some linear forms in $\mathbb{K}[a_{0},\dots,a_{N}]$. The determinant of this submatrix is $a_{0}^{4}$. Thus $a_{0}\in\sqrt{I(4,M_{Q})}$.
Next, let $M_{2}$ be the $4\times 4$ submatrix of $M_{Q}$ obtained by selecting the 0th, 1st, 3rd, and 4nd rows and columns of $M_{Q}$:
$M_{2}=\begin{array}[]{c c}&\begin{array}[]{c c c c}z_{0}&z_{1}&z_{3}&z_{4}\\ \end{array}\\ \begin{array}[]{c c c c}z_{0}\\ z_{1}\\ z_{3}\\ z_{4}\end{array}&\begin{bmatrix}0&0&a_{0}&a_{1}\\ 0&0&-a_{1}&*\\ a_{0}&-a_{1}&*&*\\ a_{1}&*&*&*\end{bmatrix}\end{array}$.
where * are some linear forms in $\mathbb{K}[a_{0},\dots,a_{N}]$. The determinant of this submatrix is $a_{0}f+a_{1}^{4}\in I(4,M_{Q})$ where $f$ is a cubic form. Since $a_{0}\in\sqrt{I(4,M_{Q})}$, it follows that $a_{1}\in\sqrt{I(4,M_{Q})}$.
Hence, $X$ does not satisfy property $\operatorname{QR}(3)$ and $\delta(X,4)-\delta(X,3)\geq 2$ by ˜2.6. ∎
Considering the analysis of the case $p=e_{2}$ (see Proposition˜4.6), we conjecture that
$\delta(X,4)-\delta(X,3)=2.$
For $d\leq 6$, we have observed that $\delta(X,4)-\delta(X,3)=2$ for randomly generated values of $\alpha_{0},\alpha_{1},\alpha_{2}$ provided by the computer algebra system [6].
5.2. A double point and a simple point.
Let $p$ be a point in $\mathcal{C}_{d}^{3}\setminus\mathcal{C}_{d}^{2}$ for $d\geq 5$ such that $\pi_{p}(\mathcal{C}_{d})\cap\mathbb{L}$ consists of a double point and a simple point. This implies that $(f_{p})^{\perp}=(\ell_{1}^{2}\ell_{2},g)\subset\mathbb{K}[S,T]$ with $\deg\ell_{1}=\deg\ell_{2}=1$ and $\deg g=d-1$. We claim that the scheme $X=\pi_{p}(\mathcal{C}_{d})\cup\mathbb{L}$ for $d\geq 5$ does not satisfy property $\operatorname{QR}(3)$.
Proposition 5.4.
Let $p$ be a point in $\mathbb{P}^{d}$ such that $(f_{p})^{\perp}=(\ell_{1}^{2}\ell_{2},g)\subset\mathbb{K}[S,T]$ with $\deg l_{1}=\deg l_{2}=1$, $l_{1}$ and $l_{2}$ linearly independent, and $\deg g=d-1$. The scheme $X$ does not satisfy property $\operatorname{QR}(3)$ for $d\geq 5$.
Proof.
We first remark that it suffices to consider the point $p\in\mathbb{P}^{d}$ corresponding to $f_{p}\in\mathbb{K}[s,t]_{d}$ such that $(f_{p})^{\perp}=(T^{2}(S+\alpha T),g(S,T))$ for some $\alpha\in\mathbb{K}$, by applying an automorphism on $\mathbb{P}^{1}$ that sends the linear form $\ell_{1}$ to $T$ in $\mathbb{K}[S,T]_{1}$.
Moreover, we can take any higher-degree generator of $(f_{p})^{\perp}$ among all forms in $\{g(S,T)-T^{2}(S+\alpha T)h(S,T)\mid h(S,T)\in\mathbb{K}[S,T]_{d-4}\}$. Therefore, we can assume
$(f_{p})^{\perp}=(T^{2}(S+\alpha T),\beta_{0}S^{d-1}+\beta_{1}S^{d-2}T+\beta_{2}T^{d-1})$
for some $\alpha,\beta_{0},\beta_{1},\beta_{2}\in\mathbb{K}$.
We claim that $\sqrt{I(4,M_{Q})}$ contains at least one linear form. Following the proof of Corollary˜3.2, $I(X)_{2}$ is spanned by the generators of I(S(1,d3))I(S(1,d-3)) and
$\left\{z_{1}z_{j+3}+\alpha z_{1}z_{j+4}-\beta_{0}z_{2}z_{j+2}-\beta_{1}z_{2}z_{j+3}-\beta_{2}z_{j+2}z_{d-1}\mid 0\leq j\leq d-5\right\}.$
Then, $I(4,M_{Q})$ contains $a^{4}$ where $a$ is the coefficient of the quadratic form $z_{0}z_{3}-z_{1}z_{2}$ in $Q$ since $M_{Q}$ contains a $4\times 4$ lower triangular submatrix whose diagonal entries are $\{a,-a,-a,a\}$. Thus, $\sqrt{I(4,M_{Q})}$ contains a linear form $a$, and so $X$ fails to satisfy property $\operatorname{QR}(3)$. ∎
For $d=5$, we observe that $\delta(X,4)-\delta(X,3)=1$ for randomly generated values of $\alpha,\beta_{0},\beta_{1},\beta_{2}$ provided by the computer algebra system [6]. In addition, for a point $p=e_{1}+e_{6}\in\mathbb{P}^{6}$ (where $f_{p}^{\perp}=(S^{2}T,6S^{5}-T^{5})$), we can check that $\delta(X,4)-\delta(X,3)=1$ by [6]. These examples support ˜1.4.
5.3. Three simple points.
Suppose $X=\pi_{p}(\mathcal{C}_{d})\cup\mathbb{L}$ where the intersection $\pi_{p}(\mathcal{C}_{d})\cap\mathbb{L}$ consists of three simple points. We note that it suffices to consider the point $p\in\mathbb{P}^{d}$ such that
$(f_{p})^{\perp}=(S(S+\alpha_{1}T)(S+\alpha_{2}T),\beta_{0}S^{d-1}+\beta_{1}S^{d-2}T+\beta_{2}T^{d-1}),$
for $\alpha_{1},\alpha_{2},\beta_{0},\beta_{1},\beta_{2}\in\mathbb{K}$.
For $d=5$, we check that $X$ satisfies property $\operatorname{QR}(3)$ for randomly generated values of $\alpha_{1},\alpha_{2},\beta_{0},\beta_{1},\beta_{2}$ using the computer algebra system [6].
We claim that $X$ satisfies property $\operatorname{QR}(3)$ if
$p\;$ $=\nu_{d}([1:-1])+(-1)^{d}\nu_{d}([1:1])+(-1)^{d-1}2\nu_{d}([0:1])$
$=\sum_{i=0}^{d}(-1)^{i}e_{i}+(-1)^{d}\sum_{i=0}^{d}e_{i}+(-1)^{d-1}2e_{d}$
[…]
We first verify this claim for low-degree cases, specifically for $d=6$ and $d=7$.
Lemma 5.5.
  1. (1)
    For a point $p\in\mathbb{P}^{6}$ such that $(f_{p})^{\perp}=(S^{3}-ST^{2},T^{5})$, $\pi_{p}(C_{6})\cup\mathbb{L}$ satisfies property $\operatorname{QR}(3)$.
  2. (2)
    For a point $p\in\mathbb{P}^{7}$ such that $(f_{p})^{\perp}=(S^{3}-ST^{2},T^{6})$, $\pi_{p}(C_{7})\cup\mathbb{L}$ satisfies property $\operatorname{QR}(3)$.
Proof of (1).
Let $LT:\mathbb{P}^{6}\to\mathbb{P}^{6}$ be a linear transformation that maps a point $p$ to $e_{0}$, and let $T^{*}$ denote the $\mathbb{K}$-linear homomorphism induced by $LT$. Let $I=(g_{1},\dots,g_{8})$ be the saturated ideal of the rational normal curve $C_{6}$, and let $J=(h_{1},\dots,h_{8})$ where each $h_{i}$ for $i=1,\dots,8$ is the quadratic form obtained by applying $T^{*}$ to the variables in $g_{i}$.
Then,
$I(\pi_{p}(C_{6}))=J\cap\mathbb{K}[y_{1},\dots,y_{6}],$
where $\mathbb{K}[y_{0},y_{1},\dots,y_{6}]$ is the coordinate ring of the image under $LT$. The ideal of $X=\pi_{p}(\mathcal{C}_{6})\cup\mathbb{L}$ is given by $\left(J\cap\mathbb{K}[y_{1},\dots,y_{6}]\right)_{2}$. Considering the isomorphism $f:H^{0}(C_{6},\mathcal{O}_{\mathbb{P}^{1}}(6))\to R_{1}$, the elimination of a variable is equivalent to requiring that the coefficient of the term mapping to $y_{0}$ via the $Q_{A,B}$-map is zero, given a decomposition of the line bundle $\mathcal{O}_{\mathbb{P}^{1}}(6)=A^{2}\otimes B$ on $\mathcal{C}_{6}$.
We fix the decomposition $\mathcal{O}_{\mathbb{P}^{1}}(6)=\mathcal{O}_{\mathbb{P}^{1}}(2)^{2}\otimes\mathcal{O}_{\mathbb{P}^{1}}(2)$ and let $l,m,\text{ and }n$ be general members of $H^{0}(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(2))\cong\mathbb{K}[S,T]_{2}$. The vanishing condition of the coefficient can be decomposed into following three cases:
  1. (i)
    $l=S^{2}-T^{2},~m=\gamma_{0}ST+\gamma_{1}T^{2},~n=2\gamma_{0}\gamma_{1}S^{2}-(\gamma_{0}^{2}+\gamma_{1}^{2})ST$;
  2. (ii)
    $l=S^{2}-ST,~m=\gamma_{0}ST+\gamma_{1}T^{2},~n=(-2\gamma_{0}\gamma_{1}-\gamma_{1}^{2})S^{2}+\gamma_{1}^{2}ST+(\gamma_{0}+\gamma_{1})^{2}T^{2}$;
  3. (iii)
    $l=S^{2}+ST,~m=\gamma_{0}ST,~n=(-2\gamma_{0}\gamma_{1}-\gamma_{1}^{2})S^{2}-\gamma_{1}^{2}ST+(\gamma_{0}+\gamma_{1})^{2}T^{2}$
for $\gamma_{0},\gamma_{1}\in\mathbb{K}\setminus\{0\}$.
Since we obtain degree $6$ binary forms in $\gamma_{0}$ and $\gamma_{1}$ from $Q_{\mathcal{O}_{\mathbb{P}^{1}}(2),\mathcal{O}_{\mathbb{P}^{1}}(2)}$, there are $7$ generators and there are $5$ independent quadratic forms in $I(\pi_{p}(\mathcal{C}_{6}))_{2}$ among them. Therefore, we obtain $5$ independent rank $3$ quadratic generators of $I(\pi_{p}(\mathcal{C}_{6}))$ for each cases and get $8$ independent generators in total, as verified by the computer software system [6].
Proof of (2). As in (1), the ideal of $X$ is given by $\left(J\cap\mathbb{K}[y_{1},\dots,y_{7}]\right)_{2}$ where $J$ is generated by the quadratic forms obtained by applying the $\mathbb{K}$-linear map $T^{*}:\mathbb{K}[x_{0},\dots,x_{7}]\to\mathbb{K}[y_{0},\dots,y_{7}]$ to the variables in the quadratic generators of the defining ideal of $C_{7}\subset\mathbb{P}^{7}$.
We fix the decomposition $\mathcal{O}_{\mathbb{P}^{1}}(7)=\mathcal{O}_{\mathbb{P}^{1}}(2)^{2}\otimes\mathcal{O}_{\mathbb{P}^{1}}(3)$. Let $l$ and $m$ be general elements of $H^{0}(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(2))\cong\mathbb{K}[S,T]_{2}$, and let $n$ be a general element of $H^{0}(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(3))\cong\mathbb{K}[S,T]_{3}$. The cases where $l$, $m$, and $n$ satisfy the vanishing condition for the coefficients are distinguished to be the following three cases:
  1. (i)
    $\begin{cases}l=S^{2}-T^{2},\;m=\gamma_{0}ST+\gamma_{1}T^{2},\\ n=\gamma_{2}(S^{3}-ST^{2})-2\gamma_{0}\gamma_{1}S^{2}T+(\gamma_{0}^{2}+\gamma_{1}^{2})ST^{2}\end{cases}$
  2. (ii)
    $\begin{cases}l=S^{2}-ST,\;m=\gamma_{1}ST+\gamma_{0}T^{2},\\ n=\gamma_{2}(S^{3}-ST^{2})+((\gamma_{1}+\gamma_{0})^{2}-\gamma_{0}^{2})(S^{2}T+ST^{2})-(\gamma_{1}+\gamma_{0})^{2}(T^{3}+ST^{2})\end{cases}$
  3. (iii)
    $\begin{cases}l=S^{2}+ST,\;m=-\gamma_{1}ST+\gamma_{0}T^{2},\\ n=\gamma_{2}(S^{3}-ST^{2})+((\gamma_{1}+\gamma_{0})^{2}-\gamma_{0}^{2})(S^{2}T-ST^{2})-(\gamma_{1}+\gamma_{0})^{2}(T^{3}-ST^{2})\end{cases}$
for $\gamma_{0},\gamma_{1},\gamma_{2}\in\mathbb{K}\setminus\{0\}$.
We obtain $9$ independent quadratic forms in $I(\pi_{p}(\mathcal{C}_{7}))_{2}$ in each case and the 27 quadratic forms span the space $I(\pi_{p}(\mathcal{C}_{d}))_{2}$ which is verified using the computer algebra system [6].
We now prove the claim that $X=\pi_{p}(\mathcal{C}_{d})\cup\mathbb{L}$ satisfies property $\operatorname{QR}(3)$ for the point $p=\sum_{i=0}^{d}(-1)^{i}e_{i}+(-1)^{d}\sum_{i=0}^{d}e_{i}+(-1)^{d-1}2e_{d}$.
Proposition 5.6.
Let $p$ be a point in $\operatorname{\mathbb{P}}^{d}$ (for $d\geq 6$) corresponding to a binary form $f_{p}$ such that $(f_{p})^{\perp}=(S^{3}-ST^{2},T^{d-1})$. Then, the scheme $X=\pi_{p}(\mathcal{C}_{d})\cup\mathbb{L}$ satisfies property $\operatorname{QR}(3)$.
Proof.
Let $LT:\mathbb{P}^{d}\to\mathbb{P}^{d}$ be a linear map on $\mathbb{P}^{d}$ that sends a point $p$ to $e_{0}$. Let $I$ be the ideal generated by the $2\times 2$ minors of the matrix
$\begin{bmatrix}x_{0}&x_{1}&\cdots&x_{d-1}\\ x_{1}&x_{2}&\cdots&x_{d}\end{bmatrix}$
and let $J$ be the ideal generated by the quadratic forms obtained by applying the $\mathbb{K}$-linear map $T^{*}:\mathbb{K}[x_{0},\dots,x_{d}]\to\mathbb{K}[y_{0},\dots,y_{d}]$ to the variables in the generators of $I$. Then
$I(\pi_{p}(\mathcal{C}_{d}))=J\cap\mathbb{K}[y_{1},\dots,y_{d}]$
where $J$ is the ideal generated by the $2\times 2$ minors of
$\begin{bmatrix}y_{0}&y_{1}&y_{2}+y_{0}&y_{3}+y_{1}&\cdots&y_{d-1}+\cdots+y_{1}\\ y_{1}&y_{2}+y_{0}&y_{3}+y_{1}&y_{4}+y_{2}+y_{0}&\cdots&y_{d}\end{bmatrix}$.
Therefore, we obtain the ideal of $X=\pi_{p}(\mathcal{C}_{d})\cup\mathbb{L}$ as follows:
[…]
$+(y_{1}y_{i+2}-y_{1}y_{i}-y_{3}y_{i}:2\leq i\leq d-3)$
The cases $d=6,7$ are covered by Lemma˜5.5.(1) and Lemma˜5.5.(2).
Suppose $d\geq 8$ and $X=\mathbb{L}\cup\pi_{p}(\mathcal{C}_{d^{\prime}})$ satisfies $\operatorname{QR}(3)$ for all integers $d^{\prime}<d$. For brevity, we omit the line bundles $A$ and $B$ from the decomposition $\mathcal{O}_{\mathbb{P}^{1}}(d)=A^{\otimes 2}\otimes B$ in the $Q_{A,B}$-map and simply write $Q$ instead.
Case 1: $d$ is even.
Among the generarors of $I=I(X)$, $2\times 2$ minors of $\begin{bmatrix}y_{2}&y_{3}&\cdots&y_{d-2}\\ y_{3}&y_{4}&\cdots&y_{d-1}\end{bmatrix}$ can be written as a sum of rank $3$ quadratic forms. Therefore, we claim that
$y_{i}y_{d}-y_{i+1}(y_{d-1}+\dots+y_{1})\sim_{I}0\text{ for }2\leq i\leq d-2$
and
$y_{1}y_{i+2}-y_{1}y_{i}-y_{3}y_{i}\sim_{I}0\text{ for }2\leq i\leq d-3.$
By the induction hypothesis, we have
$z_{i}z_{d-1}-z_{i+1}(z_{d-2}+\cdots+z_{0})=\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j})$
for some $Q(l_{j},m_{j},h_{j})\in\mathbb{K}[z_{0},z_{2},z_{3},\dots,z_{d-1}]$. Therefore, we have that
$\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j}T)=y_{i+1}y_{d}-y_{i+2}(y_{d-1}+\cdots+y_{1})$
for $2\leq i\leq d-2$. Note that we didn’t check $y_{2}y_{d}-y_{3}(y_{d-1}+\cdots+y_{1})\sim_{I}0$ yet.
By induction hypothesis,
$w_{1}w_{i+2}-w_{i}(w_{1}+w_{3})=\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j})\in k[w_{0},w_{1},w_{3},\dots,w_{d-2}]$
for $2\leq i\leq d-5$. Then, we have
$\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j}S^{2})=y_{1}(y_{i+2})-(y_{3}+y_{1})y_{i}$
for $2\leq i\leq d-5$.
Moreover, since $\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j}T^{2})=y_{3}y_{i+4}-(y_{5}+y_{3}+y_{1})y_{i+2}$ for $2\leq i\leq d-5$ and $y_{3}y_{i+4}-y_{5}y_{i+2}=(y_{3}y_{i+4}-y_{4}y_{i+3})+(y_{4}y_{i+3}-y_{5}y_{i+2})\sim_{I}0$, we obtain
$y_{1}y_{i+4}-(y_{3}+y_{1})y_{i+2}\sim_{I}0$
for $2\leq i\leq d-5$. This implies that
$y_{1}y_{i+2}-(y_{3}+y_{1})y_{i}\sim_{I}0$
for $2\leq i\leq d-3$.
Now, we claim that
$y_{2}y_{d}-y_{3}(y_{d-1}+\cdots+y_{1})\sim_{I}0.$
Suppose $k$ is an integer such that $d=2k$. Note that
(5.1) $Q(S^{k}-S^{k-2}T^{2},T^{k},1)=(-y_{2}+y_{4})y_{d}-y_{k+2}^{2},$
(5.2) $y_{4}y_{d}-y_{5}(y_{d-1}+\cdots+y_{1}),$
(5.3) $y_{1}y_{5}-(y_{3}+y_{1})y_{3},\text{ and }$
(5.4) $y_{5}y_{d-1}-y_{k+2}^{2}=\sum_{i=0}^{k-4}(y_{i+5}y_{d-1-i}-y_{i+6}y_{d-2-i})$
are spanned by rank $3$ quadratic forms. Therefore,
[…]
$=y_{2}y_{d}-(y_{5}y_{d-3}+y_{5}y_{d-5}+\cdots+y_{5}y_{3})-y_{3}(y_{3}+y_{1})$
$\sim_{I}y_{2}y_{d}-y_{3}(y_{d-1}+y_{d-3}+\cdots+y_{3}+y_{1})$
since $y_{5}y_{d-1-2t}\sim_{I}y_{3}y_{d+1-2t}$ for $1\leq t\leq k-2$.
Case 2: $d$ is odd.
By the inductive hypothesis, we have $z_{i}z_{d-1}-z_{i+1}(z_{d-2}+\cdots+z_{1})=\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j})\in\mathbb{K}[z_{1},z_{2},\dots,z_{d-1}].$ Therefore,
$\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j}T)=y_{i+1}y_{d}-y_{i+2}(y_{d-1}+\cdots+y_{0})\in\mathbb{K}[y_{0},y_{1},\ldots,y_{d}]$
for $2\leq i\leq d-2$.
On the other hand, by the inductive hypothesis, $w_{0}w_{i+2}-(w_{2}+w_{0})w_{i}=\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j})\in\mathbb{K}[w_{0},\dots,w_{d-2}]$ for $2\leq i\leq d-5$. Then,
$\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j}S^{2})=y_{0}y_{i+2}-(y_{2}+y_{0})y_{i}$
for $2\leq i\leq d-5$ and
$\sum_{j=1}^{k}Q(l_{j},m_{j},h_{j}T^{2})=(y_{2}+y_{0})y_{i+4}-(y_{4}+y_{2}+y_{0})y_{i+2}$
for $2\leq i\leq d-5$. Moreover,
$y_{0}y_{i+4}-y_{4}y_{i+2}=(y_{2}y_{i+4}-y_{3}y_{i+3})+(y_{3}y_{i+3}-y_{4}y_{i+2})\sim_{I}0$
for $2\leq i\leq d-5$. Therefore,
$y_{0}y_{i+2}-(y_{2}+y_{0})y_{i}\sim_{I}0$
for $2\leq i\leq d-3$.
Now, we claim that
(5.5) $y_{2}y_{d}-y_{3}(y_{d-1}+\dots+y_{0})\sim_{I}0.$
Let $k$ be the integer with $d=2k+1$. We have
(5.6) $(-y_{2}+y_{4})y_{d}+(-y_{3}+y_{5})(y_{d-1}+\cdots+y_{0})-2(y_{k+2})(y_{k+3})\sim_{I}0$
since (5.6) equals to $Q(S^{k}-S^{k-2}T^{2},\,T^{k},\,S+T)-Q(S^{k}-S^{k-2}T^{2},\,T^{k},\,S)-Q(S^{k}-S^{k-2}T^{2},\,T^{k},\,T)$. Moreover,
(5.7) $y_{4}y_{d}-y_{5}(y_{d-1}+\cdots+y_{0})\sim_{I}0\text{ and }$
(5.8) $y_{0}y_{5}-(y_{2}+y_{0})y_{3}\sim_{I}0.$
Therefore we have
[…]
$=\;$ $-2y_{3}(y_{d-1}+\cdots+y_{4})+2y_{5}(y_{d-1}+\cdots+y_{2})-2y_{k+2}y_{k+3}$
$=\;$ $-2\left((y_{3}y_{d-1}-y_{5}y_{d-3})+(y_{3}y_{d-3}-y_{5}y_{d-5})+\cdots+(y_{3}y_{4}-y_{5}y_{2})\right)$
$+2(y_{5}y_{d-1}-y_{k+2}y_{k+3})$
$\sim_{I}$  $\;0.$
Acknowledgement
We greatly appreciate the referees’ careful reading of the manuscript and their detailed and insightful suggestions. The first named author was supported by the Institute for Basic Science (IBS-R032-D1-2024-a00) and was also supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No. RS-2024-00414849). The second named author was supported by a KIAS Individual Grant(MG083101) at Korea Institute for Advanced Study and also was supported by the Sungshin Women’s University Research Grant of 2025(H20250019). The third named author was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government (MSIT) (No. 2022R1A2C1002784).
Data availability
No data was used for the research described in the article.